### v1

# Semiparametric KSD test: unifying score and distance-based approaches for goodness-of-fit testing

Abstract
Goodness-of-fit (GoF) tests are fundamental tools for assessing model adequacy. Score-based GoF tests are particularly appealing because they require fitting the model only once under the null. However, extending these tests to powerful nonparametric settings is challenging, due to the lack of suitable score functions. Through a class of exponentially tilted models, we show that the resulting score-based GoF tests are equivalent to the tests based on integral probability metrics (IPMs) indexed by a function class. When the class is rich, the test is universally consistent. This simple yet insightful perspective enables reinterpretation of classical distance-based testing procedures—including those based on Kolmogorov–Smirnov distance, Wasserstein-1 distance, and maximum mean discrepancy—as arising from score-based constructions. Building on this insight, we propose a new nonparametric score-based GoF test through a special class of IPM induced by kernelized Stein function class, called semiparametric kernelized Stein discrepancy (SKSD) test. Compared with other nonparametric score-based tests, the SKSD test is computationally efficient and accommodates general nuisance-parameter estimators, supported by a generic parametric bootstrap procedure. The SKSD test is universally consistent and attains Pitman efficiency. Moreover, SKSD test provides simple GoF tests for models with intractable likelihoods but tractable scores with the help of Stein’s identity and we use two popular models, kernel exponential family and conditional Gaussian models, to illustrate the power of our method. Our method achieves power comparable to task-specific normality tests such as Anderson-Darling and Lilliefors, despite being designed for general nonparametric alternatives.
1 Introduction
Consider $n$ observations $X_{1},\ldots,X_{n}$ from an unknown distribution $\mathcal{L}$. A common statistical inference begins by positing a parametric model $\{\mathcal{L}_{\theta}:\theta\in\Theta\}$ for a parameter space $\Theta\subseteq\mathbb{R}^{k}$, which may approximate the data-generating process. Inferential tasks such as confidence interval construction and hypothesis testing often rely on the model being correctly specified. Therefore, before proceeding with inference, it is crucial to assess model adequacy using a goodness-of-fit (GoF) test. In particular, we are interested in the following problem:
$H_{0}:\mathcal{L}\in\{\mathcal{L}_{\theta}\in\mathcal{P}(\mathcal{X}):\theta\in\Theta\}\quad\text{versus}\quad H_{1}:\mathcal{L}\notin\{\mathcal{L}_{\theta}\in\mathcal{P}(\mathcal{X}):\theta\in\Theta\}.$ (1)
A classical approach to goodness-of-fit (GoF) testing is the likelihood ratio test (LRT) [wilks1938large, lehmann2005testing]. However, the LRT can be suboptimal—or even powerless—when alternatives lie outside the assumed parametric family [cox1979theoretical], particularly in the presence of more nuanced nonparametric alternatives. To address this limitation, various extensions have been proposed, most notably the generalized likelihood ratio test [perlman1999emperor, fan2001generalized, fan2005nonparametric, fan2007nonparametric]. Another line of work, which does not directly rely on the likelihood function, constructs nonparametric GoF tests by comparing the estimated distributions under the null and alternative hypotheses using distance measures such as $L_{1}$, $L_{2}$, or $L_{\infty}$ [bickel1973some, azzalini1989use, hardle1993comparing]. Both approaches typically require smoothing techniques or local estimators to capture general alternatives, but as a result, they suffer from the curse of dimensionality [gonzalez2013updated].
In light of these challenges, the celebrated score-based test [rao1948large, rao1973linear] offers an appealing alternative. A key advantage of score-based tests is that they avoid fitting the unrestricted (full) model, thereby sidestepping the need for estimation under a nonparametric specification. Surprisingly, despite their wide use in low-dimensional parametric settings, relatively little attention has been paid to their extension to nonparametric frameworks. A central obstacle is that, in fully nonparametric settings, even defining an appropriate score is nontrivial, making both its construction and the interpretation of the resulting test statistic particularly challenging.
1.1 Our contributions
Our contributions fall into two categories.
Unification of score-based test and IPM-based test.
Under a class of general exponentially tilted models (ETMs), we show that the score under the model is equivalent to a special class of distance measures, known as integral probability metrics (IPMs). IPMs are characterized by a class of test functions $\mathcal{F}$. This powerful connection reveals that GoF tests based on IPMs are actually score-based tests under the ETM when the model is indexed by the same function class $\mathcal{F}$. Therefore, depending on the chosen $\mathcal{F}$, several celebrated nonparametric distance-based tests—such as those based on the Kolmogorov-Smirnov (K-S) [durbin1975kolmogorov, margolin1976tests], Wasserstein-1 (W-1) distances [hallin2021multivariate] and maximum mean discrepancy [key2025composite]—can be reinterpreted from a nonparametric score-based perspective. Despite its simplicity, this connection has not been explicitly recognized in the existing literature. Such duality of score-based and distance-based tests offers a unified interpretation for seemingly different GoF testing procedures and may inspire new test procedures with appealing computational and statistical properties.
A new nonparametric score-based test.
Leveraging such duality, we introduce a nonparametric score-based test grounded in a reproducing kernel Hilbert space (RKHS) induced by kernelized Stein functions [liu2016kernelized]. We refer to this test as the semiparametric test based on kernelized Stein discrepancy, SKSD test and its semiparametric nature arises from the nonparametric kernelized Stein function class within the tilted parametric model. In particular, SKSD test offers several desirable properties:
  1. 1.
    Computational efficiency: Because of the Stein’s identity, the use of kernelized Stein functions endows the SKSD test with a closed-form expression for its test statistic, which can be computed directly from the data without requiring numerical integration. The SKSD statistic can be computed in at most $O(n^{2})$ time, which is dimension-agnostic and can even be reduced to $O(n)$ if certain choice of kernel function is considered, making it far more scalable in practice. In contrast, test statistics based on other IPMs such as the K-S or W-1 distances typically require at least $O(n^{3})$ computation for multivariate data and are thus computationally expensive.
  2. 2.
    Flexibility of nuisance estimation: The semiparametric nature of the SKSD test accommodates general estimator $\hat{\theta}_{n}$ for the “best-fit” parameter $\theta_{0}$ under null. The test remains asymptotically valid and powerful as long as the estimator $\hat{\theta}_{n}$ satisfies a mild asymptotic linear condition. To support this generality, we develop a provably valid parametric bootstrap testing framework that applies to a broad class of estimators, including two most prominent examples: M-estimators and minimum distance estimators.
  3. 3.
    Universal power and asymptotic efficiency: Under mild regularity conditions, the SKSD test is universally powerful against any fixed alternative to the null hypothesis. Moreover, we establish the rate-optimal efficiency of the proposed test by deriving its limiting power under two local contiguous alternatives (i.e., alternatives approaching $H_{0}$ at the $O(n^{-1/2})$ rate). These results establish the Pitman efficiency of the SKSD test.
  4. 4.
    Capability to test models with intractable likelihoods: The proposed test relies only on the score function under the null, making it well-suited for models with intractable likelihoods but tractable score functions. This provides a useful diagnostic tool for complex models such as exponential family graphical models [besag1975statistical, wainwright2008graphical], kernel exponential family models [canu2006kernel], and energy-based models [grathwohl2020learning], which classical GoF test methods find challenging to handle.
Our theoretical framework combines the recent advances in variational representation of kernel-based U-statistic [fernandez2024general] and stable convergence tools developed in [hausler2015stable]. Through extensive simulation results, we demonstrate that the proposed SKSD test attains power comparable to established normality tests, including the Anderson-Darling test [anderson1952asymptotic] and Lilliefors test [lilliefors1967kolmogorov]. We apply the SKSD test to the structural detection of two popular models: kernel exponential family models [canu2006kernel] and conditional Gaussian models [arnold1999conditional], where the likelihood functions are intractable in both cases. The results show that SKSD test is able to effectively detect the true model structure in these complex settings.
1.2 Related work
Broadly speaking, the SKSD test fits into an emerging literature on nonparametric testing with kernel methods. Leveraging the kernel trick, powerful kernel-based procedures have been developed in a variety of settings, including classical two-sample testing [gretton2012kernel, sejdinovic2013equivalence, schrab2023mmd, chatterjee2025boosting], point-null goodness-of-fit testing [chwialkowski2016kernel, liu2016kernelized, balasubramanian2021optimality, hagrass2026minimax], independence testing [gretton2007kernel, gretton2010consistent, deb2020measuring, albert2022adaptive, shekhar2023permutation], and conditional testing [zhang2011kernel, huang2022kernel, chatterjee2024kernel, zhang2025doubly]. In particular, the point-null GoF test with KSD is a special case of our SKSD test when $\Theta$ is only a single point.
Our SKSD test is also closely related to recent works on goodness-of-fit testing under composite null hypotheses. [key2025composite] proposed a composite goodness-of-fit test based on MMD with a dedicated estimation procedure and test statistic, while [bruck2025distribution] studied model selection and specification testing for parametric models via linearized MMD statistics with Gaussian limits. These approaches rely on MMD-based tests and numerical integration, whereas our SKSD test provides a more efficient way to handle intractable models, allowing for flexible nuisance-parameter estimators. We also note that [key2025composite] discussed a special case of a composite KSD-based test in which the nuisance parameter estimate is a particular minimum-KSD estimator, without any theoretical guarantee on the validity. This test can be viewed as a special case of the general SKSD test framework.
1.3 Organization of the paper and notations
We begin by establishing a connection between score-based tests and IPM-based tests in Section 2. Building on this bridge, we introduce the SKSD test in Section 3, together with a generic parametric bootstrap calibration. In the same section, we also derive the asymptotic properties of the SKSD test under the null and show bootstrap consistency. In Section 4, we prove that the SKSD test is universally powerful against general alternative hypothesis and analyze its local power. Section 5 illustrates the practical utility of our method through applications to goodness-of-fit testing of normality, kernel exponential family models, and conditional Gaussian models. We conclude in Section 6 with discussion on limitations and potential future work. All proofs and additional experimental results are deferred to the Appendix. Before proceeding, we summarize the notation used throughout the paper.
Spaces and Measures.
Suppose $\mathcal{X}$ is a Polish space, i.e., a complete and separable metric space, and let $\mathscr{B}(\mathcal{X})$ denote the Borel $\sigma$-algebra generated by the open sets of $\mathcal{X}$. Consider two probability measures $\mathcal{L}_{p},\mathcal{L}_{q}\in\mathcal{P}(\mathcal{X})$, the collection of all probability measures on $(\mathcal{X},\mathscr{B}(\mathcal{X}))$, with differentiable density functions $p$ and $q$. Let $p_{\theta}$ denote the density function of $\mathcal{L}_{\theta}$ with respect to the Lebesgue measure $\lambda$. Define the score function of $\mathcal{L}_{p}$ as $s_{p}(x)=\nabla_{x}\log p(x)$. Note that this score function is different from the score $\bm{s}_{\theta}(x)$ defined in (3) since the differentiation is with respect to $x$ instead of the parameter $\gamma$. We use the bold symbol $\bm{s}_{\theta}(x)$ to denote the latter to avoid confusion.
Vectors, matrices and functions.
For a vector $u,v\in\mathbb{R}^{d}$ in Euclidean space, we use $\|v\|$ to denote its Euclidean norm and $\langle u,v\rangle$ to denote the inner product. For a matrix $A\in\mathbb{R}^{d_{1}\times d_{2}}$, we use $\|A\|_{2}$ to denote its operator norm, i.e., $\|A\|_{2}=\sup_{\|x\|\leq 1}\|Ax\|$. For higher order tensors, the operator norm is defined similarly. For a matrix $A\in\mathbb{R}^{d\times d}$, we use $\mathrm{Tr}(A)$ to denote its trace, i.e., $\mathrm{Tr}(A)=\sum_{i=1}^{d}A_{ii}$. Let $\mathcal{F}_{d,m}$ be the space of functions from $\mathcal{X}\subseteq\mathbb{R}^{d}$ to $\mathbb{R}^{m}$, for integers $d,m\in\mathbb{N}$ and $\mathcal{C}^{\alpha,\beta}(A,B)$ be the space of functions $f:A\times B\to\mathbb{R}$ such that $\partial_{x}^{\alpha}\partial_{y}^{\beta}f(x,y)$ is continuous for all $x\in A$ and $y\in B$.
Kernels.
The kernel in our setting is a measurable function $K(\cdot,\cdot):\mathcal{X}\times\mathcal{X}\to\mathbb{R}$. We say a kernel is symmetric if $K(x,y)=K(y,x)$ for all $x,y\in\mathcal{X}$. We say a kernel is positive definite if for any $x_{1},\ldots,x_{n}\in\mathcal{X}$ and $c_{1},\ldots,c_{n}\in\mathbb{R}$, we have $\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}K(x_{i},x_{j})\geq 0$.
2 Bridging score-based and IPM-based tests
Building on a surprisingly simple tilting model, we establish the connection of score-based tests with integral probability metrics (IPMs). We introduce such equivalence in Section 2.1, harness the insights to define nonparametric score-based tests and discuss the computational complexity of the IPM-based tests (now also score-based tests) in Section 2.2.
2.1 Exponentially tilted model: a revisit on score-based tests
Classical score-based tests, such as the Rao’s score test [rao1948large, rao1973linear], are based on the score of the parametric model. For a parametric model $\{\mathcal{L}_{(\gamma,\theta)}:(\gamma,\theta)\in\Gamma\times\Theta\subseteq\mathbb{R}\times\mathbb{R}^{k}\}$, the score is defined as:
$\bm{s}_{\theta}(X)\equiv\frac{\partial}{\partial\gamma}\frac{1}{n}\sum_{i=1}^{n}\log\mathcal{L}_{(\gamma,\theta)}(X_{i})\Big|_{\gamma=0}.$
Here, $\theta$ should be viewed as a nuisance parameter while $\gamma$ is the main parameter of interest. The GoF testing can be performed with $\bm{s}_{\hat{\theta}_{n}}(X)$ with a consistent estimator $\hat{\theta}_{n}$ of $\theta$ under the null model. We revisit two classical examples of score tests based on $\bm{s}_{\hat{\theta}_{n}}(X)$ to prepare for the generalization to nonparametric models.
Example 1 (Testing mean of Gaussian location shift model, [student1908probable, rao1948large]).
Let $\mathcal{L}_{(\gamma,\theta)}$ be a Gaussian model with mean $\gamma$ and variance $\theta^{2}$. When testing $\gamma=0$, the score can be computed as $\bm{s}_{\hat{\theta}_{n}}(X)=-\frac{1}{\hat{\theta}_{n}^{2}}\frac{1}{n}\sum_{i=1}^{n}X_{i}$. Then if we choose the linear function class, $\mathcal{F}\equiv\{ax+b:a\in[0,1],b\in\mathbb{R}\}$, the following holds:
$\max_{f\in\mathcal{F}}\Big|\frac{1}{n}\sum_{i=1}^{n}f(X_{i})-\mathbb{E}_{\mathcal{L}_{(0,\hat{\theta}_{n})}}[f(X)]\Big|^{2}=(\hat{\theta}_{n})^{2}\max_{a\in[0,1]}|a||\bm{s}_{\hat{\theta}_{n}}(X)|^{2}=(\hat{\theta}_{n})^{2}|\bm{s}_{\hat{\theta}_{n}}(X)|^{2}.$
In other words, the score-based test for testing the mean under Gaussian model can be viewed as a test based on the departure of the first moment of data from the null model.
Example 2 (Testing exponentiality under Gamma family, [moran1951random, haywood2008distribution]).
Let $\mathcal{L}_{(\gamma,\theta)}$ be the Gamma distribution family with shape parameter $\gamma+1$ and scale parameter $1/\theta$. The model is well-defined when when $\gamma>-1$ and $\theta>0$. When $\gamma=0$, the model boils down to the exponential distribution $\{\mathcal{L}_{\theta}\in\mathcal{P}(\mathcal{X}):p_{\theta}=\exp(-\theta x),\theta>0\}$. When testing exponentiality ($\gamma=0$) under the Gamma distribution family, the score can be computed as $\bm{s}_{\hat{\theta}_{n}}(X)=\frac{1}{n}\sum_{i=1}^{n}\log X_{i}-\mathbb{E}_{\mathcal{L}_{(0,\hat{\theta}_{n})}}[\log X]$. Then if we choose the logarithmic function class, $\mathcal{F}\equiv\{a\log(x):a\in[0,1]\}$, the following holds:
$\max_{f\in\mathcal{F}}\Big|\frac{1}{n}\sum_{i=1}^{n}f(X_{i})-\mathbb{E}_{\mathcal{L}_{(0,\hat{\theta}_{n})}}[f(X)]\Big|^{2}=\max_{a\in[0,1]}|a||\bm{s}_{\hat{\theta}_{n}}(X)|^{2}=|\bm{s}_{\hat{\theta}_{n}}(X)|^{2}.$
In other words, the score-based test can be viewed as a test based on the departure of the logarithmic moment of the data from the null model when it comes to exponential model nested in Gamma family.
It is clear that these score-based tests compare the empirical moments of the transformed data with the theoretical moments under the null model. The transformation function class $\mathcal{F}$ is highly problem-specific. Moreover, the null model $p_{\theta}$ is exponentially tilted towards the alternative indexed by the transformation function: $\mathcal{L}_{(\gamma,\theta)}\propto p_{\theta}(x)\cdot\exp(\gamma f(x))$ for $f\in\mathcal{F}$. This can be easily checked for the above two examples. Such observation motivates us to generalize the score-based tests to a broader class of models and function classes under so-called, exponentially tilted model (ETM). Given a class of measurable functions $\mathcal{F}$ and any $f\in\mathcal{F}$, consider the ETM:
$\mathcal{L}_{(\gamma,\theta)}(x;f)=\frac{\exp(\gamma f(x))p_{\theta}(x)}{\int\exp(\gamma f(x))p_{\theta}(x)\,\mathrm{d}x},\quad\text{where}\quad(\gamma,\theta)^{\top}\in\mathbb{R}^{k+1}.$ (2)
In Example 1, the null model $p_{\theta}$ is Gaussian with mean 0 and variance $\theta$. In Example 2, $p_{\theta}$ is an exponential distribution with scale $1/\theta$, assuming $\theta>0$. Let $\hat{\theta}_{n}$ be a consistent estimate of nuisance parameter $\theta$ under the null model. The score with respect to the parameter of interest $\gamma$ is then:
$\bm{s}_{\hat{\theta}_{n}}(X,f)\equiv\frac{\partial}{\partial\gamma}\frac{1}{n}\sum_{i=1}^{n}\log\mathcal{L}_{(\gamma,\hat{\theta}_{n})}(X_{i};f)\Big|_{\gamma=0}=\frac{1}{n}\sum_{i=1}^{n}f(X_{i})-\mathbb{E}_{\mathcal{L}_{(0,\hat{\theta}_{n})}}[f(X)].$ (3)
The score $\bm{s}_{\hat{\theta}_{n}}(X,f)$ simplifies to the empirical average of $f$ evaluated on the data, centered by its expectation under the null. Given a fixed $f\in\mathcal{F}$, the test statistic $|\bm{s}_{\hat{\theta}_{n}}(X,f)|^{2}$ can be used to detect deviations from the null model (2). The choice of $f$ determines the type of alternative the test is sensitive to, as exemplified in Examples 1-2. By aggregating over all $f\in\mathcal{F}$, we arrive at the following test statistic:
$\max_{f\in\mathcal{F}}|\bm{s}_{\hat{\theta}_{n}}(X,f)|=\max_{f\in\mathcal{F}}\Big|\frac{1}{n}\sum_{i=1}^{n}f(X_{i})-\mathbb{E}_{\mathcal{L}_{(0,\hat{\theta}_{n})}}[f(X)]\Big|.$ (4)
The maximum of difference-in-mean in (4) is known as the integral probability metric (IPM, [muller1997integral]). Intuitively speaking, the test based on IPMs compare any moment difference of the data with the null model over all the functions in a prespecified function class $\mathcal{F}$, and different choices of $\mathcal{F}$ yield different IPMs. In this way, the well-studied score-based testing procedures can be interpreted as tests based on IPMs under specific choices of $\mathcal{F}$. It is thus natural to ask whether the reverse is also true: can we interpret any IPM-based tests as score-based tests with different choices of $\mathcal{F}$? We address this question in the next section.
2.2 Nonparametric score-based tests via characteristic IPMs
Tests based on IPMs can be powerful against nonparametric alternatives when sufficiently rich function classes $\mathcal{F}$ are used. The form and expressiveness of $\mathcal{F}$ fundamentally determine the capacity of the aggregated score statistic to capture discrepancies between the null and alternative. This insight has motivated a rich body of literature on hypothesis testing using IPMs [gretton2012kernel, sriperumbudur2012empirical, sejdinovic2013equivalence, liu2016kernelized, paik2023maximum]. In particular, we define an IPM is characteristic if
$\mathrm{IPM}(\mathcal{L}_{\theta},\mathcal{L};\mathcal{F})\equiv\max_{f\in\mathcal{F}}\left|\mathbb{E}_{\mathcal{L}}[f(X)]-\mathbb{E}_{\mathcal{L}_{\theta}}[f(X)]\right|=0\quad\text{if and only if}\quad\mathcal{L}=\mathcal{L}_{\theta}.$ (5)
Examples of popular characteristic IPMs and related divergences include $\{f:f=\mathbbm{1}(A),A\subseteq\mathcal{X}\}$, Kolmogorov-Smirnov (K-S) distance; $\{f:|f(x)-f(y)|/\|x-y\|_{2}\leq 1\}$, Wasserstein-1 (W-1) distance; a reproducing kernel Hilbert space (RKHS): MMD. Through the equivalence in (4), tests based on these IPMs can be interpreted as nonparametric score-based tests under the exponentially tilted models (2), provided the IPM is characteristic. We formally define such tests below:
Definition 1 (Nonparametric score-based test).
A score-based test statistic of the form $\sup_{f\in\mathcal{F}}|\bm{s}_{\theta}(X,f)|^{2}$, with $X\sim\prod_{i=1}^{n}\mathcal{L}(X_{i})$, is called a nonparametric score-based test if $\mathcal{F}$ induces a characteristic IPM, $\mathrm{IPM}(\mathcal{L}_{\theta},\mathcal{L};\mathcal{F})$.
According to the definition, the simplest way to verify that a score-based test is nonparametric is to confirm that the corresponding IPM is characteristic, as per definition (5). It is known that the K-S and W-1 distances are characteristic under mild assumptions. Similarly, MMD is characteristic when appropriate kernels and regularity conditions are satisfied [gretton2012kernel, muandet2017kernel]. Despite being characteristic, however, critical computation challenges arise in the case of a nonparametric score-based test, when $\mathcal{F}$ is chosen to be sufficiently rich and complex. To illustrate this, we discuss the computational complexity of several test statistics induced by different IPMs below.
Kolmogorov-Smirnov (K-S) Distance.
In the univariate case ($d=1$), the classical Kolmogorov-Smirnov (K-S) distance between an empirical distribution and a known distribution can be computed efficiently using sorting algorithms, yielding a computational complexity of $O(n\log n)$. However, in the bivariate case ($d=2$), the worst-case computational complexity increases to $O(n^{2})$ [peacock1983two, fasano1987multidimensional]. When $d>2$, computing the K-S distance becomes an NP-hard to compute [gnewuch2009finding].
Wasserstein-1 (W-1) Distance.
Similarly, when $d=1$, the W-1 distance can be computed efficiently in $O(n\log n)$ time via sorting. For $d>1$, however, the problem can be formulated as an assignment problem through discretization, and the computational complexity generally grows to $O(n^{3})$ [kuhn1955hungarian], although approximation algorithms exist to improve efficiency [altschuler2017near].
Maximum Mean Discrepancy (MMD).
Among nonparametric score-based tests, MMD provides computationally efficient testing procedures. Computing the MMD test statistic requires Monte Carlo (MC) integration. If $m$ denotes the number of MC samples, the overall computational cost becomes $O(n^{2}+2nm+m^{2})$ [gretton2012kernel]. When $m\ll n$, the cost is effectively quadratic; however, to control the MC error adequately, one often needs $m\gg n$, in which case the total complexity increases to $O(m^{2})$.
Distance Metric Computational Complexity
$d=1$ $d=2$ $d>2$
K-S distance $n\log n$ $n^{2}$ NP-hard
W-1 distance $n\log n$ $n^{3}$ $n^{3}$
MMD (U-/V- statistic) $n^{2}+m^{2}$ $n^{2}+m^{2}$ $n^{2}+m^{2}$
KSD (U-/V- statistic) $n^{2}$ $n^{2}$ $n^{2}$
Table 1: Computational complexity of various distance-based tests. The displayed values represent the asymptotic order of time complexity.
Kernelized Stein Discrepancy (KSD).
A special class of IPM is KSD, which is also a special case of MMD [liu2016kernelized]. It can be shown to be characteristic under certain mild conditions [barp2024targeted]. In contrast to MMD, KSD does not require MC integration for any expectation. Leveraging the Stein’s identity, KSD test statistic can be computed with $O(n^{2})$ time complexity, although further acceleration is possible [jitkrittum2017linear] (see Appendix A.2 for details). Table 1 summarizes the computational complexity of the distance-based tests discussed above.
From a GoF testing perspective, compared to K-S, W-1 distances and MMD, KSD is significantly more efficient especially when it comes to multivariate settings. Therefore, KSD emerges as a natural choice for nonparametric score-based testing, especially in the multivariate setting. Nevertheless, for univariate cases, i.e. $d=1$, the higher computational cost of kernel-based methods may be justified by gains in test power, a topic we revisit in Appendix L.
3 Semiparametric KSD test
In the following sections, we will first review the preliminaries of KSD in Section 3.1 and then propose the main methodology in Section 3.2 with the investigation of its asymptotic behaviors in Section 3.3. We establish a calibration procedure based on parametric bootstrap and study its weak convergence in Section 3.4.
3.1 Preliminaries: kernelized Stein class and KSD
First introduced in the machine learning community, kernelized Stein discrepancy (KSD, [liu2016kernelized]) was proposed as a measure of model discrepancy in settings where the model likelihood is intractable. Its popularity is largely due to its computational efficiency (see Table 1), its desirable characteristic property that distinguishes between probability distributions, and, crucially, its ability to handle intractable models. We refer the reader to [anastasiou2023stein] for a comprehensive review of these developments. KSD belongs to the general class of Stein discrepancy (SD), which is defined via Stein operator $\mathcal{A}_{p}:\mathcal{F}_{d,d}\to\mathcal{F}_{d,1}$. Recalling $s_{p}=\nabla_{x}\log p(x)$ is the score function and Stein operator is defined as $\mathcal{A}_{p}g(x)=\sum_{j=1}^{d}\partial_{j}g_{j}(x)+g^{\top}(x)s_{p}(x)$, where $g\in\mathcal{F}_{d,d}$, and $g_{j}(x)$ denotes the $j$-th coordinate of $g(x)$. Under suitable regularity conditions, the following identity holds: $\mathbb{E}_{X\sim\mathcal{L}_{p}}[\mathcal{A}_{p}g(X)]=0$, as established, for example, in Proposition 1 of [gorham2015measuring]. This is the so-called Stein’s identity [stein1972bound, gorham2015measuring]. Using this identity, the SD between $\mathcal{L}_{p}$ and $\mathcal{L}_{q}$ can be defined as $\mathrm{SD}(\mathcal{L}_{p},\mathcal{L}_{q})=\sup_{g\in\overline{\mathcal{F}}_{d,d}}\left|\mathbb{E}_{p}[\mathcal{A}_{p}g(X)]-\mathbb{E}_{q}[\mathcal{A}_{p}g(X)]\right|$ for a compact set $\overline{\mathcal{F}}_{d,d}\subseteq\mathcal{F}_{d,d}$ with $\mathbb{E}_{p}[\mathcal{A}_{p}g(X)]=0$ by Stein’s identity. If we take $\mathcal{F}_{d,d}=\mathcal{H}^{d}$, the product space of a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, and $\overline{\mathcal{F}}_{d,d}$ to be the unit ball in $\mathcal{H}^{d}$, then the resulting SD is called the kernelized Stein discrepancy (KSD):
$\mathrm{KSD}(\mathcal{L}_{p},\mathcal{L}_{q})\equiv\sup_{\begin{subarray}{c}\|g\|_{\mathcal{H}^{d}}\leq 1\end{subarray}}\left|\mathbb{E}_{p}[\mathcal{A}_{p}g(X)]-\mathbb{E}_{q}[\mathcal{A}_{p}g(X)]\right|=\sup_{\begin{subarray}{c}\|g\|_{\mathcal{H}^{d}}\leq 1\end{subarray}}\left|\mathbb{E}_{q}[\mathcal{A}_{p}g(X)]\right|.$ (6)
We refer readers who are unfamiliar with RKHS to [kanagawa2018gaussian] for a detailed introduction and review. If we further define the function class to be the kernelized Stein class, $\mathcal{F}_{\mathrm{KSD}}=\{\mathcal{A}_{p}g:\|g\|_{\mathcal{H}^{d}}\leq 1\}$, then definition (6) fits the framework of the IPM in (4):
$\mathrm{KSD}(\mathcal{L}_{p},\mathcal{L}_{q})=\sup_{f\in\mathcal{F}_{\mathrm{KSD}}}\left|\mathbb{E}_{p}[f(X)]-\mathbb{E}_{q}[f(X)]\right|$
with $\mathbb{E}_{p}[f(X)]=0$ for any $f\in\mathcal{F}_{\mathrm{KSD}}$. Furthermore, one can show under mild conditions on the kernel function, the KSD is characteristic and we refer interested readers to Appendix A.4 for more details.
3.2 SKSD test with general nuisance estimators
Recalling $\hat{\mathcal{L}}_{n}$ is the empirical measure $\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$, our proposed semiparametric KSD (SKSD) test statistic is defined as:
$T_{n}(X,\hat{\theta}_{n})\equiv\mathrm{KSD}^{2}(\mathcal{L}_{\hat{\theta}_{n}},\hat{\mathcal{L}}_{n})=\sup_{g\in\mathcal{H}^{d}}(\mathbb{E}_{\hat{\mathcal{L}}_{n}}[\mathcal{A}_{p_{\hat{\theta}_{n}}}g(X)])^{2}.$ (SKSD)
For the optimization problem (SKSD), a closed-form solution can be obtained by applying the reproducing property (the well-known “kernel trick”), yielding:
$T_{n}(X,\hat{\theta}_{n})=\mathbb{E}_{X,X^{\prime}\sim\mathcal{L}_{p_{\hat{\theta}_{n}}}}[h_{p_{\hat{\theta}_{n}}}(X,X^{\prime})]=\frac{1}{n^{2}}\sum_{i,j=1}^{n}h_{p_{\hat{\theta}_{n}}}(X_{i},X_{j}),$ (7)
where the explicit form of $h$ is provided in Appendix A.1. The alternative formulation (7) enables a more computationally efficient estimate of KSD, as it only requires evaluating the expectation of a bivariate function $h$. A nice property of the SKSD test statistic is that it does not rely on the likelihood function of the model $\mathcal{L}_{\hat{\theta}_{n}}$. In fact, the function $h$ only involves the score function $s_{p_{\hat{\theta}_{n}}}$ under the model $\mathcal{L}_{\hat{\theta}_{n}}$ and kernel function $K$. Moreover, evaluating function $h$ does not require any Monte Carlo simulation and the V-statistic (7) can be computed in quadratic time (recall Table 1). Moreover, there are certain choices of kernel function that can reduce the complexity to even $O(n)$, such as linear kernel (See Appendix A.2 for more details).
Another appealing property of SKSD is the compatibility with general estimation procedure for $\hat{\theta}_{n}$. When $\hat{\theta}_{n}=\theta_{0}\in\Theta$, test statistic (SKSD) boils down to the KSD statistic [liu2016kernelized, chwialkowski2016kernel]. In most of statistical practice, however, the parameter $\theta_{0}$ is unknown and needs to be estimated from the data. The SKSD test statistic is designed to accommodate the general estimation procedure for $\theta_{0}$. We allow the estimation of $\hat{\theta}_{n}$ to be highly problem-specific and impose no restrictions on the estimation procedure, provided the resulting estimator satisfies a so-called uniform asymptotic linear estimate, formalized in the following definition.
Definition 2 (Uniform asymptotic linear estimate).
Consider the estimation procedure $\mathcal{E}:\mathcal{X}^{n}\rightarrow\Theta\subseteq\mathbb{R}^{k}$ and an open set $S\subseteq\Theta$. If there exists a bivariate function $I(\cdot,\cdot):\mathcal{X}\times\Theta\rightarrow\mathbb{R}^{k}$ such that for any $X=(X_{1},\ldots,X_{N})\overset{\mathrm{i.i.d.}}{\sim}\mathcal{L}_{\theta}\in\mathcal{L}_{S}$, the following holds: $\mathbb{E}[I(X_{i},\theta)]=0,\ \sup_{\theta\in\Theta}\mathbb{E}[\left\|I(X_{i},\theta)\right\|^{2}]<\infty$ and
$\sup_{\theta\in S}\Big|\sqrt{n}(\hat{\theta}_{n}-\theta)-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}I(X_{i},\theta)\Big|\overset{p}{\rightarrow}0\quad\text{where}\quad\hat{\theta}_{n}=\mathcal{E}(X).$ (8)
Then we say $I$ is an influence function and estimator $\hat{\theta}_{n}$ is an $(S,\mathcal{E})$-uniformly asymptotic linear estimate (UALE) with influence function $I$.
Throughout this paper, we will assume that $\hat{\theta}_{n}$ is a UALE with an estimation procedure $\mathcal{A}$, formalized in the following assumption.
Assumption 1 (Regularity of the estimate).
Suppose there exists $\theta_{0}\in\Theta$ such that $\hat{\theta}_{n}-\theta_{0}=o_{\mathbb{P}}(1)$. Furthermore, suppose there exists $\delta>0$ such that $B_{\theta_{0}}(\delta)$ is an open ball centered at point $\theta_{0}$ with radius $\delta$. Then $\hat{\theta}_{n}$ is a $(B_{\theta_{0}}(\delta),\mathcal{E})$-UALE with influence function $I$ and a prespecified estimation algorithm $\mathcal{E}$.
Definition 2 is a slightly stronger requirement than the classical asymptotic linear estimate (ALE) assumption [van2000asymptotic, Chapter 5,] because of the uniformity requirement, but still considerably general. Assumption 1 requires the estimator $\hat{\theta}_{n}$ to admit a limit $\theta_{0}$ in probability with asymptotic linear expansion uniformly over a neighborhood of the limit. In particular, classical parametric estimators such as M-estimator and minimum distance estimator [wolfowitz1957minimum] can be shown to be UALE under very mild regularity conditions. The formulation of these estimators can be found in Appendix A.3.
3.3 Consistency and asymptotic null distribution
In this section, we study the large-sample property of the SKSD test statistic, and the following theorem shows that the SKSD test is consistent.
Theorem 1 (Consistency of $\mathrm{SKSD}$ test statistic).
Under Assumption 1 and the regularity conditions in Appendix D, we have $T_{n}(X,\hat{\theta}_{n})\overset{p}{\rightarrow}\mathrm{KSD}^{2}(\mathcal{L}_{\theta_{0}},\mathcal{L})$ as $n\rightarrow\infty$.
The proof of Theorem 1 is provided in Appendix F.1. Theorem 1 shows that the SKSD test statistic converges to the KSD statistic under both null and alternative hypothesis, which is a natural extension of the KSD test statistic to the semiparametric setting. Further, we make the following remark on the limit parameter $\theta_{0}$.
Remark 1 (On the limit parameter $\theta_{0}$).
It is worth noting that the limit of the test statistic $T_{n}(X,\hat{\theta}_{n})$ depends on the realization of the parameter $\theta_{0}$ in Assumption 1, which in turn can be related to the choice of estimation algorithm $\mathcal{E}$. In particular, under the null hypothesis any consistent estimator satisfies $\mathcal{L}=\mathcal{L}_{\theta_{0}}$, so the test statistic converges to zero as desired. However, under the alternative hypothesis, different estimation algorithms may lead to different limits $\theta_{0}$ even when the data are generated from the same distribution. For example, when $\mathcal{E}$ is empirical risk minimization (ERM) with respect to different loss functions, the corresponding $\mathcal{L}_{\theta_{0}}$ is the projection of the true distribution $\mathcal{L}$ onto the model class $\{\mathcal{L}_{\theta}:\theta\in\Theta\}$ under the discrepancy metrics induced by those losses. This freedom of choice can, however, substantially affect the power of the resulting test procedure. We investigate the impact of different estimation algorithms on power in Sections 5.1 and 5.3.
We now establish the asymptotic distribution of the observed test statistic $T_{n}(X,\hat{\theta}_{n})$, which we refer to as $T_{n}$ when there is no ambiguity. To formally state the following results, we define the following functional $S^{*}:\mathcal{H}^{d}\times\mathcal{X}\times\Theta\to\mathbb{R}$:
$S^{*}(f,X,\theta)=\mathcal{A}_{\theta}f(X)+\big\langle\mathbb{E}_{X^{\prime}\sim P_{\theta}}\left[[\nabla_{\theta}s_{\theta}(X^{\prime})]^{\top}f(X^{\prime})\right],I(X,\theta)\big\rangle.$ (9)
To formally state the theorem, Let $K_{x}$ be feature map of $x$ in $\mathcal{H}$ (see Appendix A.1 for a rigorous definition) and $\mathbf{K}_{x}=K_{x}\bm{1}_{d}$, where $\bm{1}_{d}$ is the $d$-dimensional all-ones vector. The following theorem characterizes the asymptotic null distribution of the SKSD test statistic.
Theorem 2 (Asymptotic null distribution).
Suppose $X_{1},\ldots,X_{n}$ are i.i.d. samples from $\mathcal{L}_{\theta_{0}}$ with $\theta_{0}\in\Theta$. Under Assumption 1 and regularity conditions in Appendix D, we have $nT_{n}$ converges in distribution to $W$ as $n\rightarrow\infty$, i.e.,
$\limsup_{n\rightarrow\infty}\left|\mathbb{P}_{H_{0}}[nT_{n}\leq t]-\mathbb{P}[W\leq t]\right|=0\quad\forall t\in\mathbb{R},$
where $W\overset{d}{=}\sum_{u=1}^{\infty}\lambda_{u}Z_{u}^{2}$ with $Z_{u}$ to be independent standard normal variables and $\lambda_{u}$’s to be the eigenvalues of the operator $T_{\sigma}:\mathcal{H}^{d}\to\mathcal{F}_{d,1}$ defined as $T_{\sigma}(f)(x)=\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}[S^{*}(f,X,\theta_{0})S^{*}(\mathbf{K}_{x},X,\theta_{0})]$ for any $f\in\mathcal{H}^{d}$ and $x\in\mathcal{X}$.
The proof of Theorem 2 is provided in Appendix F.2. We refer readers to Appendix B for additional results on the properties of this test statistic. The proof of the theorem is sketched in the following remark.
Remark 2 (Proof sketch of Theorem 2).
First, we show that under Assumption 1, the test statistic $T_{n}(X,\hat{\theta}_{n})$ is asymptotically equivalent to the supremum of a stochastic process indexed by the function class $\mathcal{H}^{d}$, whose structure depends on the estimation algorithm $\mathcal{E}$. This requires the generalization of common analysis of degenerate V-statistics to accommodate the nuisance parameter estimation. Second, we adopt functional central limit theorem to show the stochastic process above converges to a Gaussian process weakly. The target result is then obtained by applying the continuous mapping theorem to the supremum of the Gaussian process.
3.4 Bootstrap test: agnostic consistency guarantee
Informed by Theorem 2, the SKSD test statistic converges to a complicated distribution (infinite weighted sum of $\chi^{2}$) under $H_{0}$ (see Theorem 2). In light of this challenge, we propose a practical level-$\alpha$ testing procedure based on parametric bootstrap. The bootstrap test is summarized in Algorithm 1.
Input: Distribution class $\{\mathcal{L}_{\theta}:\theta\in\Theta\}$, data $X=(X_{1},\ldots,X_{n})$, estimation algorithm $\mathcal{E}:\mathcal{X}^{n}\rightarrow\Theta\subseteq\mathbb{R}^{k}$, bootstrap size $B$.
1 Obtain estimate $\hat{\theta}_{n}=\mathcal{E}(X)$;
2 Compute $T_{n}\equiv T_{n}(X,\hat{\theta}_{n})$ as in (SKSD);
3 for $b=1,2,\dots,B$ do
4   Generate bootstrap sample $\tilde{X}^{(b)}=(\tilde{X}_{1}^{(b)},\ldots,\tilde{X}_{n}^{(b)})\overset{\text{i.i.d.}}{\sim}\mathcal{L}_{\hat{\theta}_{n}}$;
5   Obtain bootstrap estimate $\tilde{\theta}_{n}^{(b)}=\mathcal{E}(\tilde{X}^{(b)})$;
6   Compute the resampled statistic $\widetilde{T}_{n}^{(b)}\equiv T_{n}(\tilde{X}^{(b)},\tilde{\theta}_{n}^{(b)})$.
7 end for
Output: $p$-value $\sum_{b=1}^{B}\mathbbm{1}(\widetilde{T}_{n}^{(b)}\geq T_{n})/B$.
Algorithm 1 Semiparametric KSD GoF testing procedure
We want to point out that using parametric bootstrap is certainly not new in the literature. It has been a standard approach for goodness-of-fit testing [freedman1981bootstrapping, genest2008validity]. In the context of SKSD test, in addition to resolving the intractable distribution, there are two other critical reasons to favor the parametric bootstrap procedure. First, in the general context of GoF testing, it has been pointed out that other forms of bootstrap methods can fail including nonparametric bootstrap [bollen1992bootstrapping] and wild bootstrap [key2025composite, brueck2025composite]. Second, the proposed bootstrap procedure (Algorithm 1) is agnostic to the estimation procedure $\mathcal{E}$ and there is no need to estimate any nuisance parameter involved. We make the following remark on the computation complexity of the proposed bootstrap procedure.
Remark 3 (Computation complexity of Algorithm 1).
The computational complexity of Algorithm 1 is driven by three components: (1) the cost of the estimation procedure $\mathcal{E}$, denoted by $E(n)$ per fit with $n$ samples; (2) the cost of computing the test statistics $T_{n}$ and $\widetilde{T}_{n}$, which is $O(n^{2})$ per evaluation; and (3) the cost of generating bootstrap samples, denoted by $S(n)$ to obtain $n$ samples. Consequently, the total complexity is on the order of $O(B(E(n)+n^{2}+S(n)))$. In regimes where estimation or resampling is substantially more expensive (e.g., $E(n)\gg n^{2}$ or $S(n)\gg n^{2}$), the total cost can greatly exceed $O(Bn^{2})$. One possible remedy that avoids repeated resampling and refitting is to construct a Neyman-orthogonalized kernel [chernozhukov2022locally] so that the influence function no longer appears in (9) and wild bootstrap can be applied to achieve the validity [escanciano2024gaussian]. However, the resulting procedure has complexity $O(n^{3}+Bn^{2})$ (see Appendix C for details). Hence, there is a trade-off between the cost of computing the test statistic and the costs associated with estimation and resampling.
The following result shows the asymptotic distribution of the bootstrap test statistic $\widetilde{T}_{n}$ matches the asymptotic distribution of $T_{n}(X,\hat{\theta}_{n})$ with arbitrary UALE estimators under general hypothesis.
Theorem 3 (Bootstrap consistency).
Suppose Assumption 1 and regularity conditions in Appendix D hold. Write $\widetilde{T}_{n}\equiv T_{n}(\tilde{X},\tilde{\theta}_{n})$ and $\mathcal{F}_{n}$ as the $\sigma$-algebra induced by data $X_{1},\ldots,X_{n}$. Then for any given $\delta>0$, we have
$\lim_{n\rightarrow\infty}\mathbb{P}\left[\big|\mathbb{P}[n\widetilde{T}_{n}\leq t|\mathcal{F}_{n}]-\mathbb{P}[W\leq t]\big|>\delta\right]=0\quad\text{for any}\quad t\in\mathbb{R},$ (10)
where random variable $W$ is defined as in Theorem 2.
The proof of Theorem 3 is provided in Appendix F.3. We make two remarks illustrating the significance of Theorem 3.
Remark 4 (Agnosticity of Theorem 3 on hypothesis).
Theorem 3 is a strong result in the context of bootstrap consistency. Unlike most of the existing bootstrap consistency results in kernel-based testing literature, Theorem 3 provides a universal weak limit regardless of how the data is generated, i.e., from the null or alternative hypothesis, despite the complexity of the asymptotic null distribution. The closest result to Theorem 3 is Theorem 5 in [key2025composite]. However, they only provide bootstrap validity under the null hypothesis when the MMD is used as the test statistic at the presence of a nuisance parameter. Therefore, the results cannot be used to analyze the power of the test. In contrast, our result allows us to understand the asymptotic behavior of the proposed test under both null and alternative hypotheses, providing particularly strong theoretical support for power analysis.
Remark 5 (Machinery to prove Theorem 3).
To prove Theorem 3, we regard the result as a conditional version of weak convergence result in Theorem 2. The extra critical techniques used here are results on stable convergence [hausler2015stable], which provides a precise depiction of the weak convergence of the conditional distribution in the sense of Markov kernel convergence. We then show the proof strategy in Theorem 2 can be adapted to the conditional case, and be proved to hold stably. The property of martingale stable convergence leads to the target result.
Consider the bootstrap test (letting $B\rightarrow\infty$ in Algorithm 1):
$\phi_{n,\alpha}\equiv\mathbbm{1}(T_{n}\geq\mathbb{Q}_{1-\alpha}(\widetilde{T}_{n}^{(b)})),$ (SKSD-test)
where $\mathbb{Q}_{1-\alpha}(\mu)$ denotes the $(1-\alpha)$-th quantile of the measure $\mu$. A direct corollary of Theorem 3 is that the proposed bootstrap test $\phi_{n,\alpha}$ is asymptotically valid under the null hypothesis.
Corollary 1.
Suppose the assumptions in Theorems 2 and 3 hold. Then, we have $\lim_{n\rightarrow\infty}\mathbb{P}_{H_{0}}[\phi_{n,\alpha}=1]=\alpha$.
Corollary 1 guarantees the asymptotic validity of the SKSD test under a generic nuisance estimation procedure. In practice, when applying the proposed method with a particular class of estimators, analysts can readily verify Assumption 1, along with other standard regularity conditions.
4 Power analysis of SKSD test
With the asymptotic distribution of $T_{n}(X,\hat{\theta}_{n})$ and the bootstrap consistency result, we can now establish the asymptotic validity and consistency of the proposed bootstrap test $\phi_{n,\alpha}$.
Theorem 4 (Universal consistency against fixed alternative).
Suppose Assumptions 1 and regularity conditions in Appendix D hold. If $\inf_{\theta\in\Theta}\mathrm{KSD}(\mathcal{L}_{\theta},\mathcal{L})>0$ under $H_{1}$, then we have $\lim_{n\rightarrow\infty}\mathbb{P}_{H_{1}}[\phi_{n,\alpha}=1]=1$.
Theorem 4 shows that the proposed bootstrap test is consistent against any fixed alternative distribution that is separated from the null model in terms of KSD. This result is quite general as it holds for any UALE estimator $\hat{\theta}_{n}$ and does not require any further specification on the alternative distribution. We emphasize that the agnostic bootstrap consistency result under general hypotheses (Theorem 3) is what makes such power analyses possible. In practice, what is more interesting is the power behavior of the proposed test under local alternative models, which we will study in the following section. Throughout the rest of the section, we allow the law generating data, $\mathcal{L}$, to depend on the sample size $n$ and adopt the triangular array setup. We will use $\mathcal{L}_{n}$ to emphasize the dependence. Consider the following assumption.
Assumption 2 (Local altervatives with regularity).
Suppoe either of the two conditions holds:
  1. a
    (Multiplicative local alternative, [jankova2020goodness]) Consider $H_{1,n}:\mathcal{L}_{n}\propto p_{\theta_{0}}(x)(1+h(x)/\sqrt{n})$ for some function $h(x):\mathbb{R}^{d}\to\mathbb{R}$. Further, suppose $h(X)$ has finite second moment under the null hypothesis, i.e., $\mathbb{E}_{\theta_{0}}[h^{2}(X)]<\infty.$
  2. b
    (Additive local alternative, [huber1965robust, niu2022distribution, chatterjee2025boosting]) Consider $H_{1,n}:\mathcal{L}_{n}=(1-\gamma/\sqrt{n})p_{\theta_{0}}(x)+(\gamma/\sqrt{n})g(x)$ for some density function $g(x):\mathbb{R}^{d}\to\mathbb{R}$ and constant $\gamma>0$. Further, suppose $\mathbb{E}_{\theta_{0}}\big[\big(\frac{g(X)}{p_{\theta_{0}}(X)}\big)^{2}\big]<\infty$ and denote $\mathcal{L}_{g}$ as the law with density $g$.
The regularity conditions in each case essentially imply the existence of the local alternatives as proper distributions. We are now ready to state the main result on the power analysis under multiplicative local alternatives. The proof of Theorem 5 is provided in Appendix G.
Theorem 5 (Power under local alternative models).
Suppose Assumptions 1, 2 and regularity conditions in Appendix D hold. Then for data generated from local alternative model , we have
$\lim_{n\rightarrow\infty}\mathbb{P}_{H_{1,n}}[\phi_{n,\alpha}=1]=\beta\quad\text{where}\quad\beta=\mathbb{P}[W_{1}\geq w_{1-\alpha}]\geq\alpha.$
The random variable $W_{1}\overset{d}{=}\sum_{u=1}^{\infty}\lambda_{u}(Z_{u}+\mu_{u})^{2}$ and $w_{1-\alpha}$ is the $(1-\alpha)$-quantile of $W$, where $\lambda_{u}$’s, $Z_{u}$’s and $W$ are defined in Theorem 2. If we write the eigenfunction corresponding to eigenvalue $\lambda_{u}$ of the operator $T_{\sigma}$ (defined in Theorem 2) as $\phi_{u}$, then we have $\mu_{u}=\mathbb{E}_{\theta_{0}}[S^{*}(\phi_{u},X,\theta_{0})h(X)]$ under Assumption 2a and $\mu_{u}=\gamma\mathbb{E}_{\mathcal{L}_{g}}[S^{*}(\phi_{u},X,\theta_{0})]$ under Assumption 2b. In particular, when there exists $u\in\mathbb{N}_{+}$ such that $\mu_{u}\neq 0$, we have $\beta>\alpha$.
Theorem 5 characterizes the asymptotic power function of our test under both multiplicative and additive local alternatives. The asymptotic power is fully characterized by the mean shift parameters $\mu_{u}$ and SKSD test $\phi_{n,\alpha}$ will have non-trivial power for local alternatives scaling with $1/\sqrt{n}$ as long as some $\mu_{u}\neq 0$. For multiplicative local alternatives, equivalently, this means the existence of some $f\in\mathcal{H}^{d}$ such that $\mathbb{E}_{\theta_{0}}[S^{*}(f,X,\theta_{0})h(X)]\neq 0$. This requires that under measure $\mathbb{P}_{\theta_{0}}$, tilting function $h(X)$ is correlated with at least one functional in the set $\{S^{*}(f,X,\theta_{0}):f\in\mathcal{H}^{d}\}$. For additive local alternatives, correspondingly, this is equivalent to the existence of some $f\in\mathcal{H}^{d}$ such that $\mathbb{E}_{\mathcal{L}_{g}}[S^{*}(f,X,\theta_{0})]\neq 0$. Overall, Theorem 5 shows the Pitman efficiency of our test. We demonstrate the finite-sample power of the SKSD test under these alternatives in Section K.
5 Application with SKSD test
In this section, we demonstrate the practical utility of our proposed SKSD test through three representative applications: testing normality in Section 5.1, determining order of kernel exponential family in Section 5.2, and detecting graphical structure in conditional Gaussian models in Section 5.3.
5.1 Testing normality
Testing normality is arguably the most classical GoF testing problem. A variety of specialized goodness-of-fit tests for normality have also been developed. In this section, we conduct simulations to assess the Type-I error and power of SKSD test for normality under different nuisance estimation procedures.
Experiment setup.
In this experiment, we consider four different data generating processes (DGPs) with Gaussian models the null model: (1) Gaussian distribution with different means $\mu$; (2) non-centered Student-$t$ distribution with varying degrees of freedom $\nu$; (3) mixture of two Gaussians with the same mixture weights and different mean shift $\delta$; and (4) non-centered generalized $\chi^{2}$ distribution with varying power parameter $\alpha$. Specifically, the DGPs are defined as follows: We refer to Appendix J.1 for the explicit forms of the distributions.
Methods compared.
To provide a comprehensive evaluation, we carry out simulations to examine finite-sample performance under various nuisance estimation strategies. The estimation and testing procedures used in this experiment are as follows. For estimation, we use minimum-KSD estimator [barp2019minimum] and MLE are used to estimate the parameters $(\mu,\sigma)$ in the null model $N(\mu,\sigma^{2})$. For testing, we consider the SKSD test, likelihood-ratio test (LRT), specialized normality tests, the Anderson-Darling test [anderson1952asymptotic] and the Lilliefors test [lilliefors1967kolmogorov], which are widely used in practice. The LRT in this context conincides with the well-known Vuong’s test [vuong1989likelihood]. Since we do not find a general method to calibrate the test invovled the nuisance parameter estimation, we use parametric bootstrap to approximate the null distribution of all tests.
Simulation parameter and kernel choices.
The sample size is set to $n=100$. The kernel used in the SKSD test is the Gaussian kernel with bandwidth selected based on the median heuristic [gretton2012kernel]. The nominal level $\alpha$ is set to be $0.05$. We keep the kernel and level choice fixed across all following experiments in this section.
Figure 1: Power curves of SKSD goodness-of-fit tests for Gaussian models under various null and alternative distributions. The data generating processes across the subfigures: top left: Gaussian distribution; top right: non-centered Student-$t$ distribution; bottom left: mixture of Gaussians; bottom right: non-centered generalized $\chi^{2}$ distribution. The black dashed line indicates the test level.
Simulation results.
The results are summarized in Figure 1, where “min-KSD + KSD” and “MLE + KSD” curves correspond to our proposed test. First, when data are generated under the null hypothesis, the SKSD test and the specialized normality tests (Anderson-Darling and Lilliefors) achieve the nominal level without noticeable inflation. In contrast, the LRT-based with either the minimum-KSD or MLE estimator is substantially conservative. Second, under the three alternatives, the SKSD tests exhibit power comparable to that of the specialized normality tests. In some settings, such as the Gaussian mixture model, the SKSD tests even outperform both the Anderson-Darling and Lilliefors tests. The power of the LRT-based tests, however, are quite unstable across different setups. Third, the choice of estimation algorithm $\mathcal{A}$ on the nuisance parameter does lead to differences in the power of the SKSD tests although the power gap is not substantial.
5.2 Order detection of kernel exponential model
Let $q$ be a reference density on $\mathbb{R}^{d}$, and let $\kappa:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ be a reproducing kernel associated with RKHS $\mathcal{H}_{\kappa}$. The kernel exponential family model [canu2006kernel] is indexed by a function $f\in\mathcal{H}_{\kappa}^{d}$ and is given by
$p_{f}(x)\propto q(x)\exp\left(\left\langle f,\kappa(x,\cdot)\mathbf{1}_{d}\right\rangle_{\mathcal{H}_{\kappa}^{d}}\right),\quad x\in\mathbb{R}^{d},$
where we set $\kappa$ to be the standard Gaussian kernel and $q(x)$ is set to be the density of $N(0,9)$. In this case, $p_{f}$ is parameterized by the functional element $f$, and the normalizing constant is typically intractable. In practice, because $\mathcal{H}_{\kappa}^{d}$ is infinite dimensional, the element $f$ cannot be estimated directly. A common strategy is to approximate $f$ by a finite-rank representation with sufficient complexity, which yields accurate density approximations [strathmann2015gradient, sutherland2018efficient]. In this experiment, we therefore focus on goodness-of-fit testing for the finite-rank approximation of the kernel exponential family.
Experiment setup.
For the null model, we assume $f(\cdot)=\sum_{\ell=1}^{p}\theta_{\ell}\phi_{\ell}(\cdot)$, with rank $p$, coefficients $\theta_{\ell}\in\mathbb{R}$ and basis function $\phi_{\ell}$. The following two settings are considered: (1) fix $p=1$ for the null model and generate data from the model with various parameter $\theta_{2}$; (2) consider the null models with different rank $p$ and generate data from the model with fixed rank $p=5$ and various choices of $\theta_{2}$. For both settings, we generate the sample with metroplis-adjusted Langevin algorithm (MALA, [roberts2001optimal]) and set the sample size to be $n=200$. We use the minimum-KSD estimator to estimate the parameters in the density model. This has become a standard approach in the literature when dealing with the nuisance parameter estimation problem [matsubara2022robust, liu2024robustness]. We refer the reader to Appendix J.2 for the complete simulation setup.
Results.
Figure 2: Power curve of SKSD goodness-of-fit tests for kernel exponential family. The red dashed line indicates the test level.
$\theta_{2}$ $p=1$ $p=2$ $p=3$ $p=4$
$0.0$ 0.900 0.354 0.250 0.080
$0.5$ 0.916 0.360 0.322 0.138
$1.0$ 0.784 0.450 0.280 0.202
$1.5$ 0.862 0.404 0.354 0.236
$2.0$ 0.938 0.382 0.348 0.148
$2.5$ 0.952 0.436 0.378 0.222
$3.0$ 0.932 0.334 0.446 0.232
Table 2: SKSD test power for various values of $\theta_{2}$ and null model rank pp.
In the first setting, data are generated from the null model when $\theta_{2}=0$ and from the alternative otherwise. Hence, the rejection rate should be close to the nominal level $\alpha=0.05$ at $\theta_{2}=0$ and increase as $|\theta_{2}|$ grows. Figure 2 confirms this: the Type-I error is close to $\alpha$ at $\theta_{2}=0$, and the power increases with $|\theta_{2}|$, reaching nearly 1 once $|\theta_{2}|\geq 3$, illustrating that the SKSD test is consistent and powerful in this setting. In the second setting, as $p$ increases, the null model becomes a higher-rank approximation of the kernel exponential family, which is supposed to capture more information on $f$. Table 2 shows that the power is highest at $p=1$ and generally decreases with $p$. This suggests that SKSD test can be a powerful stepwise model selection strategy for the kernel exponential family model.
5.3 Graph structure detection in conditional Gaussian model
In this application, we consider quadratic interaction model in Gaussian conditional family [gelman1991note, arnold1999conditional], which is defined as follows:
$p_{\gamma}(x)\propto\exp\left(\sum_{1\leq i\neq j\leq d}\Sigma_{ij}(x^{(i)})^{2}(x^{(j)})^{2}+\sum_{k=1}^{d}\gamma_{k}^{(2)}(x^{(k)})^{2}+\sum_{l=1}^{d}\gamma_{\ell}^{(1)}x^{(\ell)}\right)$ (11)
where in $x\in\mathbb{R}^{d}$ with parameters $\theta=(\Sigma,\gamma^{(1)},\gamma^{(2)})$ and $\Sigma\in\mathbb{R}^{d\times d}$ and $\gamma^{(1)},\gamma^{(2)}\in\mathbb{R}^{d}$. The key property of this model is that the conditional distributions are in Gaussian family. Note that the conditional Gaussian model is substantially different from the Gaussian graphical model and indeed, using graphical Lasso will lead to poor performance in terms of the graph recovery task [lin2016estimation]. Note the normalizing constant in (11) is infeasible to compute as it involves high-dimensional intractable integrals. However, the score of the model can be easily computed and we refer to Appendix J.3 for the explicit form of the score function.
Experiment setup.
Without loss of generality, we consider $\Sigma$ is a symmetric matrix, i.e., $\Sigma_{ij}=\Sigma_{ji}$ for all $i,j$, and make the following constraints on the parameter space: $\Sigma_{ij}\leq 0$, $\gamma_{i}^{(2)}<0$, which ensures the existence of the joint and conditional probability distributions. We want to know if a quadratic interaction model with a ring structure, i.e., $\Sigma_{ij}=0$ for all $|i-j|\neq 1\text{ or }d-1$, can fit the model well enough. When generating the data, we consider alternatives only introducing second-order interaction terms between non-adjacent nodes in the graph: $\Sigma_{ij}=0$ for all |ij|>2|i-j|>2 or |ij|<d2|i-j|<d-2. We are interested in two settings: (1) fix dimension $d=8$ and vary the magnitude of the second-order interaction $\varepsilon$. We are under the null model when $\varepsilon=0$ and under the alternative otherwise; (2) fix the magnitude of the second-order interaction terms $\varepsilon=0.5$ and vary the dimension $d$ to generate data from different distributions, which are all under the alternatives. The parameter $\varepsilon\geq 0$ (see Appendix J.3 for the definition of $\varepsilon$) controls the magnitude of the second-order interaction terms in both settings. Sampling from model (11) can be easily implemented by Gibbs sampler [arnold2001conditionally]. We use sample size $n=500$ throughout the experiments. We refer to Appendix J.3 for additional details on the experiments.
Methods compared.
For estimation, we use minimum-KSD estimator [barp2019minimum] and score matching estimator [hyvarinen2005estimation, lin2016estimation] to estimate the parameters of the interaction terms in the null model with 1D grid structure. Both estimators admit closed-form solutions, which are given in Appendix J.3. For testing, the SKSD test with parametric bootstrap cutoff is considered (Algorithm 1).
Results.
The experimental results for the two settings are summarized in Figure 3. In the first setting with fixed $d$, the power of the test increases with $\varepsilon$ for both the score-matching estimator and the minimum-KSD estimator. Under the null, the Type I error of both estimators is well controlled at the nominal level. In the second setting, the power increases with $d$. Intuitively, in higher dimensions there are more node pairs, so the cumulative signal for detecting second-order interactions is stronger as the number of interacting nodes grows. In both setups, when the signal is weak (small $\varepsilon$ or small $d$), the score-matching estimator yields slightly higher power. As the signal strengthens, however, the minimum-KSD estimator becomes more powerful. Since score matching has $O(n)$ time complexity whereas minimum-KSD has $O(n^{2})$, these results highlight a trade-off between statistical power and computational cost. They also suggest that the choice of estimation algorithm can substantially affect the power of the SKSD test. We investigate the mechanism behind this pattern in the next section.
Figure 3: Top: power curves of SKSD goodness-of-fit tests for quadratic interaction graphical model. The left panel demonstrates the power of our proposed test with varying $\varepsilon$. The right panel shows the power of the same test with fixed $\varepsilon$ and varying $d$. Bottom: the orange curves are the ratios between RMSEs of the estimators and the green curves are the ratios of projection distances between estimators onto the null model, both defined as in (12).
Unveiling the mystery between different estimators.
It turns out that the different estimators trade off estimation error and separation from the null in SKSD tests. To understand these patterns, we compare the root mean squared error (RMSE) of the estimators and the distance of their population projections onto the null model. Let $\Sigma^{\mathrm{proj}}_{\mathrm{KSD}}$ (resp. $\Sigma^{\mathrm{proj}}_{\mathrm{SM}}$) be the projection of the true interaction matrix $\Sigma$ onto the null model parameter space under the loss functions of minimum-KSD (resp. score-matching) estimation procedure. For each estimator $\mathcal{A}\in\{\mathrm{KSD},\mathrm{SM}\}$, we define
$\mathrm{RMSE}(\mathcal{A})=\Bigl(\mathbb{E}\bigl[\|\hat{\Sigma}_{\mathcal{A}}-\Sigma^{\mathrm{proj}}_{\mathcal{A}}\|_{\mathrm{F}}^{2}\bigr]\Bigr)^{1/2},\quad\mathrm{Proj\text{-}dist}(\mathcal{A})=\|\Sigma^{\mathrm{proj}}_{\mathcal{A}}-\Sigma\|_{\mathrm{F}},$ (12)
where $\|\cdot\|_{\mathrm{F}}$ is the Frobenius norm. The RMSE reflects estimation accuracy of different estimators, while the projection distance measures how far the projected model is from the null under alternative. Lower RMSE and larger projection distance are expected to yield higher power. To deconfound the finite-sample error, we compute the population projections using $10^{4}$ Monte Carlo samples and the RMSE using $10^{3}$ independent repetitions under the simulation setup.
The resulting ratios $\mathrm{RMSE}(\mathrm{KSD})/\mathrm{RMSE}(\mathrm{SM})$ and $\mathrm{Proj\text{-}dist}(\mathrm{KSD})/\mathrm{Proj\text{-}dist}(\mathrm{SM})$ are also shown in Figure 3. In the first setting, as $\varepsilon$ increases, the RMSE ratio decreases while the projection-distance ratio increases, matching the trend of power curves: score matching is favored when $\varepsilon$ is small (RMSE dominates), and minimum-KSD gains an advantage for larger $\varepsilon$ (projection distance dominates). In the second setting, as $d$ increases, the RMSE ratio remains nearly constant while the projection-distance ratio increases, consistent with the faster growth in power for minimum-KSD relative to score matching. In other words, it is the projection distance, i.e., the distance between the projected model and the null model, that matters most in determining the relative performance of the two estimators.
6 Discussion
In this paper, we establish a simple yet powerful connection between score-based and distance-based GoF tests and IPMs through the lens of general exponentially tilted models. This framework generalizes the classical parametric score-based test to a nonparametric setting and allows us to reinterpret many existing IPM-based methods as nonparametric score-based tests. Motivated by the computational and statistical limitations of existing tests, we propose a new nonparametric score-based test, the SKSD test paired with parametric bootstrap calibration, which can be efficiently computed and is broadly applicable to a wide range of parametric models. We prove its asymptotic properties under both the null and (local) alternative hypotheses, and we demonstrate the flexibility and effectiveness of the proposed test through extensive experiments on various models, including settings in which only the score function is available.
There are many promising directions for future work. First, it would be interesting to extend our framework to discrete distributions, where the score function can be defined through the discrete Stein operator [yang2018goodness]. Second, the choice of bandwidth in the kernel function of the SKSD test is a challenging problem and can have a substantial impact on the power of the resulting test. It would be valuable to draw inspiration from recent work on improving the power of kernel-based tests by aggregating multiple kernels [schrab2022ksd, schrab2023mmd, chatterjee2025boosting]. Lastly, the null model can be more complex, going beyond the fixed-dimensional setting to high- or even infinite-dimensional cases [zhu2020distance, zhu2021interpoint, yan2023kernel, huang2023high, gao2023two], and the composite nature of the GoF problem further complicates the analysis. An important direction is to investigate how estimation error in high-dimensional settings manifests in the nonparametric GoF testing problem and how this, in turn, suggests principled solutions.
Appendix A Additional details on KSD
A.1 Derivation of closed-form solution for KSD
We first explicitly define the function $h_{p}$ for some density $p$ introduced in Eqn. (7):
$h_{p}(x,y)$ $=s_{p}(x)^{\top}s_{p}(y)K(x,y)+s_{p}(x)^{\top}\nabla_{y}K(x,y)$
$\qquad\qquad\qquad\qquad+s_{p}(y)^{\top}\nabla_{x}K(x,y)+\mathrm{Tr}(\nabla_{x}\nabla_{y}K(x,y)).$
Then, we show how to derive the closed-form solution of KSD in Eqn. (6). To see this, we recall the Riesz representation theorem [reed1972methods, Theorem II.4,] so that for an RKHS $\mathcal{H}$, there exists a positive definite kernel $K:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ such that the feature map $\psi_{x}\in\mathcal{H}$ for all $x\in\mathcal{X}$ satisfies $K(x,\cdot)=\psi_{x}(\cdot)$ and $K(x,y)=\langle\psi_{x},\psi_{y}\rangle_{\mathcal{H}}$. We additionally define the operation of Stein operator $\mathcal{A}_{p}$ on a real-value function (e.g., feature map) as $\xi_{p,x}(\cdot)\in\mathcal{H}^{d}$:
$\xi_{p,x}(\cdot)\equiv s(x)\psi_{x}(\cdot)+\nabla\psi_{x}(\cdot).$
The reproduciable property of $\mathcal{H}$ leads to
$\sup_{g\in B_{\mathcal{H}^{d}}(1)}\left|\mathbb{E}_{q}[\mathcal{A}_{p}g(X)]\right|^{2}$ $=\sup_{g\in B_{\mathcal{H}^{d}}(1)}\left\langle g(\cdot),\mathbb{E}_{X\sim q}\left[\xi_{p,X}(\cdot)\right]\right\rangle_{\mathcal{H}^{d}}^{2}$
$=\left\|\mathbb{E}_{X\sim q}\left[\xi_{p,X}(\cdot)\right]\right\|_{\mathcal{H}^{d}}^{2}$
$=\left\langle\mathbb{E}_{X\sim q}\left[\xi_{p,X}(\cdot)\right],\mathbb{E}_{Y\sim q}\left[\xi_{p,Y}(\cdot)\right]\right\rangle_{\mathcal{H}^{d}}$
$=\mathbb{E}_{X,Y\sim q}\left[\langle\xi_{p,X}(\cdot),\xi_{p,Y}(\cdot)\rangle_{\mathcal{H}^{d}}\right]$
$=\mathbb{E}_{X,Y\sim q}\left[\mathcal{A}_{p}^{1}\mathcal{A}_{p}^{2}[K(X,Y)I_{d}]\right],$ (13)
where the superscript $1$ (resp. $2$) means operating on the first (resp. second) argument of the function. By direct calculation, we have
$\mathcal{A}_{p}^{1}\mathcal{A}_{p}^{2}K(X,Y)I_{d}=h_{p}(X,Y),$ (14)
where the function $h_{p}$ is defined above. Plugging Eqn. (14) back into Eqn. (13) gives
$\mathrm{KSD}^{2}(\mathcal{L}_{p},\mathcal{L}_{q})=\sup_{g\in B_{\mathcal{H}^{d}}(1)}\left|\mathbb{E}_{q}[\mathcal{A}_{p}g(X)]\right|^{2}=\mathbb{E}_{q}[h_{p}(X,Y)],$
which proves Eqn. (6).
A.2 Computation of V-statistic for KSD
Recall that the formula of V-statistic for KSD is given by
$T_{n}(X,\theta)=\mathrm{KSD}^{2}(\mathcal{L}_{\theta},\hat{\mathcal{L}}_{n})=\frac{1}{n^{2}}\sum_{i,j=1}^{n}h_{p_{\theta}}(X_{i},X_{j}).$
For general kernel function, $h_{p_{\theta}}(X_{i},X_{j})$ can be computed through the closed-form expression above, so the computation cost of the statistic is $O(n^{2})$. When linear kernel $K(x,y)=x^{\top}y$ is applied in the computation, we have
$h_{p_{\theta}}(x,y)$ $=s_{p_{\theta}}(x)^{\top}s_{p_{\theta}}(y)(x^{\top}y)+s_{p_{\theta}}(x)^{\top}\nabla_{y}(x^{\top}y)$
$\qquad\qquad\qquad\qquad+s_{p_{\theta}}(y)^{\top}\nabla_{x}(x^{\top}y)+\mathrm{Tr}(\nabla_{x}\nabla_{y}x^{\top}y)$
$=s_{p_{\theta}}(x)^{\top}s_{p_{\theta}}(y)(x^{\top}y)+s_{p_{\theta}}(x)^{\top}x+s_{p_{\theta}}(y)^{\top}y+d$
$=\mathrm{Tr}(xs_{p_{\theta}}(x)^{\top}ys_{p_{\theta}}(y)^{\top}+xs_{p_{\theta}}(x)^{\top}+ys_{p_{\theta}}(y)^{\top}+I_{d})$
$=\mathrm{Tr}\left(\left(xs_{p_{\theta}}(x)^{\top}+I_{d}\right)(ys_{p_{\theta}}(y)^{\top}+I_{d})\right).$
By plugging the above equation into the V-statistic, we have
$T_{n}(X,\theta)$ $=\frac{1}{n^{2}}\sum_{i,j=1}^{n}h_{p_{\theta}}(X_{i},X_{j})$
$=\frac{1}{n^{2}}\sum_{i,j=1}^{n}\mathrm{Tr}\left(\left(X_{i}s_{p_{\theta}}(X_{i})^{\top}+I_{d}\right)(X_{j}s_{p_{\theta}}(X_{j})^{\top}+I_{d})\right)$
$=\mathrm{Tr}\left(\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}s_{p_{\theta}}(X_{i})^{\top}+I_{d}\right)^{2}\right),$
which requires only $O(n)$ time to compute.
A.3 Two general classes of estimators
We present two most popular estimation frameworks to estimate the nuisance parameter $\theta_{0}$.
  • M-estimator: For some general loss function $\ell(\cdot,\cdot)$, we define the true yet unknown estimand $\theta_{0}$ to be the population risk minimizer and empirical risk minimizer:
    $\theta_{n}=\arg\min_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\ell(\theta,X_{i}).$ (15)
    Special instances in this class include maximum likelihood estimator (MLE), maximum pseudo-likelihood estimate (MPLE) for models with intractable likelihood [besag1975statistical].
  • Minimum distance estimator: For some statistical distance $D(\cdot,\cdot)$, we define the estimand and estimator as, respectively:
    $\theta_{n}=\arg\min_{\theta\in\Theta}D(\mathcal{L}_{\theta},\mathcal{L}_{n}).$ (16)
    Special instances in this class include minimum Wasserstein estimator [bernton2019parameter], minimum MMD estimator[briol2019statistical, niu2023discrepancy, alquier2024universal] and minimum-KSD estimator [barp2019minimum, matsubara2022robust].
A.4 Nonparametric characteristicity
Moreover, as an IPM, KSD is characteristic (see Definition (5)) by the following proposition.
Proposition 1 (Theorem 2.2 in [chwialkowski2016kernel]).
Suppose $\mathbb{E}_{\mathcal{L}_{q}}h_{P}(X,X)<\infty$, and $\mathbb{E}_{\mathcal{L}_{q}}\|\nabla\log p(X)-\nabla\log q(X)\|^{2}<\infty$. If kernel $K$ is $C_{0}$-universal. Then $\mathrm{KSD}$ is characteristic, i.e., $\mathrm{KSD}(\mathcal{L}_{p},\mathcal{L}_{q})=0$ if and only if $\mathcal{L}_{p}=\mathcal{L}_{q}$.
Proposition 1 ensures that SKSD test statistic, $T_{N}(X,\hat{\theta}_{N})$, can be used to distinguish the true data generating distribution $\mathcal{L}$ from the fitted model $\mathcal{L}_{\hat{\theta}_{N}}$ as long as $\hat{\theta}_{N}\rightarrow\theta_{0}$ for some $\theta_{0}\in\Theta$. The characteristic property is especially important for the diagnostic purpose of the nonparametric GoF test. This results in the power against any alternative distribution $\mathcal{L}$ that is not in the model $\mathcal{L}_{\Theta}$ (see Theorem 4).
A.5 Connection to moment test
The score statistic can be interpreted from the moment equation guaranteed by Stein identity, $\mathbb{E}_{\hat{\mathcal{L}}_{n}}[\mathcal{A}_{p_{\hat{\theta}_{n}}}g(X)]=0$ for any test function $g$ in the RKHS $\mathcal{H}^{d}$. In other words, (SKSD) is a nonparametric moment test. It turns out that the moment test has been widely studied in the goodness-of-fit (GoF) testing literature, especially in the econometrics community [white1996estimation].
Appendix B Auxiliary results in Section 3.4
Theorem 1 shows that the limit distribution of the test statistic is a mixture chi-squared random variables. In this case, we can show that the limit distribution is almost surely finite, which ensures the existence of the quantile we use in the follow-up test and prevents the degenerations of the test statistic. We summarize this result in the following proposition, the proof of which is provided in Appendix F.5.
Proposition 2.
The limit distribution of the test statistic $nT_{n}(X,\hat{\theta}_{n})$ is almost surely finite, i.e., for $W$ and $\{\lambda_{u}\}_{u=1}^{\infty}$ defined in Theorem 2,
$\mathbb{P}\left(W<\infty\right)=\mathbb{P}\left(\sum_{u=1}^{\infty}\lambda_{u}Z_{u}^{2}<\infty\right)=1.$ (17)
Proposition 2 also implies other limit distributions used in the follow-up sections are also almost surely finite. For example, the limit of test statistics under multiplicative local alternatives, i.e. $W_{1}$ defined in Section 4, is almost surely finite. In fact, notice that
$W_{1}\leq 2\sum_{u=1}^{\infty}\lambda_{u}Z_{u}^{2}+2\sum_{u=1}^{\infty}\lambda_{u}\mu_{u}^{2}\leq 2W+2\sum_{u=1}^{\infty}\lambda_{u}\mu_{u}^{2}.$
Combining with $\sum_{u=1}^{\infty}\mu_{u}^{2}<\infty$ (shown in the proof of Theorem 5), we have $W_{1}<\infty$ a.s. The similar argument can be applied to $W_{1}^{\prime}$ in Section 4, which ensures it is also almost surely finite.
Although Theorem 3 is enough for the asymptotic validity of the bootstrap tests considered in this paper, we mention here that the stronger results automatically hold in this setting. Indeed, the convergence holds uniformly in the cutoff $t$, and we summarize it as an corollary for further use. The quantities in Corollary 2 follow the definitions in Theorem 3. The proof of Corollary 2 is provided in Appendix F.6.
Corollary 2.
Under the same assumptions as in Theorem 3, we have
$\lim_{n\rightarrow\infty}\mathbb{P}\left[\sup_{t\in\mathbb{R}}\left|\mathbb{P}[n\tilde{T}_{n}\leq t|\mathcal{F}_{n}]-\mathbb{P}[W\leq t]\right|>\delta\right]=0.$ (18)
With $\lim_{n\to\infty}\hat{\theta}_{n}=\theta_{0}$ almost surely, we additionally have
$\lim_{n\rightarrow\infty}\sup_{t\in\mathbb{R}}\left|\mathbb{P}[n\tilde{T}_{n}\leq t|\mathcal{F}_{n}]-\mathbb{P}[W\leq t]\right|=0\quad\text{almost surely}.$ (19)
Eqn. (18) is a stronger version of Theorem 3 in the sense that the quantile convergence holds uniformly in $t$. This is particularly useful when we want to do testing with multiple levels. Eqn. (19) reveals that when stronger assumption on the estimator is imposed, stronger convergence guarantees on the parametric bootstrap can be obtained. With the almost sure convergence, the bootstrap test is asymptotic valid for almost all sequences of data generated from $\mathcal{L}$, while in Theorem 3, only a probabilistic guarantee is provided.
Appendix C Neyman orthogonal SKSD test
In this section, we discuss how to construct a Neyman orthogonal SKSD test and its computational cost discussed in Section 3.4. For SKSD test, we prove in Eqn. (27) that
$T_{n}(X,\hat{\theta}_{n})=\mathrm{KSD}^{2}(\mathcal{L}_{\hat{\theta}_{n}},\mathcal{L}_{n})=\sup_{f\in\mathcal{B}_{\mathcal{H}^{d}}(1)}S_{n}(f)^{2}+o_{p}(1),$
where $\mathcal{B}_{\mathcal{H}^{d}}(1)$ is the unit ball in $\mathcal{H}^{d}$, and the functional $S_{n}(f)$ is defined as
$S_{n}(f)$ $=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left[\mathcal{A}_{\theta_{0}}f(X_{i})+\Big\langle\mathbb{E}_{X^{\prime}\sim P_{\theta_{0}}}\left[[\nabla_{\theta}s_{\theta_{0}}(X^{\prime})]^{\top}f(X^{\prime})\right],I(X_{i},\theta_{0})\Big\rangle\right],$
where the second term in the summation is due to the effect of parameter estimation, and leads to the incorrectness of the wild bootstrap procedure discussed in Section 3.4. To fix, it is possible to consider the following transformation of $f$ such that
$\widetilde{f}\coloneqq f-\nabla_{\theta}s_{\theta_{0}}(X)\left(\mathbb{E}_{\theta_{0}}[\nabla_{\theta}s_{\theta_{0}}(X)^{\top}\nabla_{\theta}s_{\theta_{0}}(X)]\right)^{-1}\mathbb{E}_{\theta_{0}}[\nabla_{\theta}s_{\theta_{0}}(X)^{\top}f(X)],$
which leads to the Neyman orthogonality condition:
$\text{and}\quad\mathbb{E}_{\theta_{0}}[[\nabla_{\theta}s_{\theta_{0}}(X)]^{\top}\widetilde{f}(X)]=0\text{ for any }f\in\mathcal{H}^{d}.$
Let $r_{\theta}(X)=\nabla_{\theta}s_{\theta}(X)$ and $\mathbf{K}(x,y)\in\mathbb{R}^{d\times d}$ be the matrix-value kernel. The transform function $\widetilde{f}$ corresponds to a transformed kernel, which we refer as the Neyman orthogonal kernel [escanciano2024gaussian] defined as
$\widetilde{\mathbf{K}}(x_{i},x_{j})$ $=\mathbf{K}(x_{i},x_{j})-r_{\theta_{0}}(x_{i})\left(\mathbb{E}_{\theta_{0}}[r_{\theta_{0}}(X)^{\top}r_{\theta_{0}}(X)]\right)^{-1}\mathbb{E}_{\theta_{0}}[r_{\theta_{0}}(X)^{\top}\mathbf{K}(X,x_{j})]$
$\quad-\mathbb{E}_{\theta_{0}}[\mathbf{K}(x_{i},X)r_{\theta_{0}}(X)]\left(\mathbb{E}_{\theta_{0}}[r_{\theta_{0}}(X)^{\top}r_{\theta_{0}}(X)]\right)^{-1}r_{\theta_{0}}(x_{j})^{\top}$
$\quad+r_{\theta_{0}}(x_{i})\left(\mathbb{E}_{\theta_{0}}[r_{\theta_{0}}(X)^{\top}r_{\theta_{0}}(X)]\right)^{-1}\mathbb{E}_{\theta_{0}}[r_{\theta_{0}}(X)^{\top}\mathbf{K}(X,X^{\prime})r_{\theta_{0}}(X^{\prime})]$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left(\mathbb{E}_{\theta_{0}}[r_{\theta_{0}}(X)^{\top}r_{\theta_{0}}(X)]\right)^{-1}r_{\theta_{0}}(x_{j})^{\top}.$
Let $\mathbf{\Pi}(x_{i},x_{j})=r_{\theta_{0}}(x_{i})\left(\mathbb{E}_{\theta_{0}}[r_{\theta_{0}}(X)^{\top}r_{\theta_{0}}(X)]\right)^{-1}r_{\theta_{0}}(x_{j})^{\top}$. For simplicity, we then omit the dependence on $\theta_{0}$ in all expectation calculation in the following. Then the Neyman orthogonal kernel can be written as
$\widetilde{\mathbf{K}}(x_{i},x_{j})=\mathbf{K}(x_{i},x_{j})-\mathbb{E}[\mathbf{\Pi}(x_{i},X)\mathbf{K}(X,x_{j})]$ $-\mathbb{E}[\mathbf{K}(x_{i},X)\mathbf{\Pi}(X,x_{j})^{\top}]$
$+\mathbb{E}[\mathbf{\Pi}(x_{i},X)\mathbf{K}(X,X^{\prime})\mathbf{\Pi}(X^{\prime},x_{j})^{\top}].$
[…]
$=s_{\theta}(x)^{\top}\widetilde{\mathbf{K}}(x,y)s_{\theta}(y)+s_{\theta}(x)^{\top}\nabla_{2}\cdot\widetilde{\mathbf{K}}(x,y)+\nabla_{1}\cdot\widetilde{\mathbf{K}}(x,y)s_{\theta}(y)+\nabla_{1}\cdot\nabla_{2}\cdot\widetilde{\mathbf{K}}(x,y),$ (21)
where $\nabla\cdot$ is the divergence, and we adopt the convention that $\nabla_{1}\cdot\mathbf{K}\in\mathbb{R}^{1\times d}$ and $\nabla_{2}\cdot\mathbf{K}\in\mathbb{R}^{d\times 1}$ for matrix-value kernel $\mathbf{K}\in\mathbb{R}^{d\times d}$. For the derivative of the kernel, we have
$\nabla_{1}\cdot\widetilde{\mathbf{K}}(x_{i},x_{j})=$ $\nabla_{1}\cdot\mathbf{K}(x_{i},x_{j})-\mathbb{E}[\nabla_{1}\cdot\mathbf{\Pi}(x_{i},X)\mathbf{K}(X,x_{j})]-\mathbb{E}[\nabla_{1}\cdot\mathbf{K}(x_{i},X)\mathbf{\Pi}(X,x_{j})^{\top}]$
$+\mathbb{E}[\nabla_{1}\cdot\mathbf{\Pi}(x_{i},X)\mathbf{K}(X,X^{\prime})\mathbf{\Pi}(X^{\prime},x_{j})^{\top}],$
$\nabla_{2}\cdot\widetilde{\mathbf{K}}(x_{i},x_{j})=$ $\nabla_{2}\cdot\mathbf{K}(x_{i},x_{j})-\mathbb{E}[\mathbf{\Pi}(x_{i},X)\nabla_{2}\cdot\mathbf{K}(X,x_{j})]-\mathbb{E}[\nabla_{2}\cdot\mathbf{K}(x_{i},X)\mathbf{\Pi}(X,x_{j})^{\top}]$
$+\mathbb{E}[\mathbf{\Pi}(x_{i},X)\nabla_{2}\cdot\mathbf{K}(X,X^{\prime})\mathbf{\Pi}(X^{\prime},x_{j})^{\top}],$
$\nabla_{1}\cdot\nabla_{2}\cdot\widetilde{\mathbf{K}}(x_{i},x_{j})=$ $\nabla_{1}\cdot\nabla_{2}\cdot\mathbf{K}(x_{i},x_{j})-\mathbb{E}[\nabla_{1}\cdot\mathbf{\Pi}(x_{i},X)\nabla_{2}\cdot\mathbf{K}(X,x_{j})]$
$-\mathbb{E}[\nabla_{1}\cdot\mathbf{K}(x_{i},X)\nabla_{2}\cdot\mathbf{\Pi}(X,x_{j})^{\top}]$
$+\mathbb{E}[\nabla_{1}\cdot\mathbf{\Pi}(x_{i},X)\mathbf{K}(X,X^{\prime})\nabla_{2}\cdot\mathbf{\Pi}(X^{\prime},x_{j})^{\top}].$
For the Neyman orthogonalized SKSD statistic, we can show that wild bootstrap is valid to approximate the null distribution of the test statistic. The test procedure is summarized in Algorithm 2.
Input: Distribution class $\{\mathcal{L}_{\theta}:\theta\in\Theta\}$, data $X=(X_{1},\ldots,X_{n})$, Neyman orthogonal kernel $\widetilde{\mathbf{K}}$, estimation algorithm $\mathcal{E}:\mathcal{X}^{n}\rightarrow\Theta\subseteq\mathbb{R}^{k}$, bootstrap size $B$.
1 Obtain estimate $\hat{\theta}_{n}=\mathcal{E}(X)$;
2 Compute $T_{n,\widetilde{\mathbf{K}}}\equiv T_{n,\widetilde{\mathbf{K}}}(X,\hat{\theta}_{n})$ as in Eqns. (20) and (C);
3 for $b=1,2,\dots,B$ do
4   Generate wild bootstrap weight $\{w_{i}^{(b)}\}_{i\in[n]}\overset{\mathrm{i.i.d.}}{\sim}\text{Unif}\{-1,+1\}$;
[…]
6 end for
Output: $p$-value $\sum_{b=1}^{B}\bm{1}(\widetilde{T}_{n,\widetilde{\mathbf{K}}}^{(b)}\geq T_{n,\widetilde{\mathbf{K}}})/B$.
Algorithm 2 Neyman orthogonalized SKSD GoF testing procedure
To compute the test statistic $T_{n,\widetilde{\mathbf{K}}}$, we need to apply $n\times n$ matrix multiplications, which has computational complexity $O(n^{3})$. For each resampled statistic based on wild bootstrap, we need additional $O(n^{2})$ computation. Thus, the total computational complexity of the Neyman orthogonalized SKSD test is $O(n^{3}+n^{2}B)$. When $n\gg B$, the computational complexity is dominated by $O(n^{3})$, which can be more expensive than the standard SKSD test. Additionally, notice that the construction of the Neyman orthogonal kernel $\widetilde{\mathbf{K}}$ involves several expectation terms under the null model, which may not have closed-form expressions. We may need to approximate these expectation terms using Monte Carlo methods, which introduces additional computational overhead in sampling and potential approximation errors on the kernel function.
Appendix D Regularity conditions
Assumption 3 (Regularity on the support set $\Theta$).
Suppose $\Theta\subseteq\mathbb{R}^{k}$ is a bounded open convex set.
Assumption 4 (Regularity on the distribution).
$\{\mathcal{L}_{\theta}\}_{\theta\in\Theta}$ is a family of distribution in $\mathbb{R}^{d}$ with support $\mathcal{X}$ and density function $p_{\theta}(x)$. For $P_{\theta}$-almost every $x$, $p_{\theta}(x)\in\mathcal{C}^{1,4}(\mathcal{X},\Theta)$, and for all $\theta\in\Theta$, there exists constants $C_{1},\ldots C_{4}$ such that for $r=1,2$, $m=2,3$, the score function $s_{\theta}(x)=\nabla_{x}\log p_{\theta}(x)$ satisfies
$\mathbb{E}_{X\sim\mathcal{L}}\left[\|s_{\theta}(X)\|^{r}\right]\leq C_{1}<\infty.$
$\mathbb{E}_{X\sim\mathcal{L}}\left[\|\nabla_{\theta}s_{\theta}(x)\|_{2}^{r}\right]\leq C_{2}<\infty.$
$\mathbb{E}_{X\sim\mathcal{L}}\left[\|\nabla_{\theta}^{(m)}s_{\theta}(x)\|_{2}^{r}\right]\leq C_{3}<\infty,$
where $\nabla^{(m)}$ denotes the $m$-th gradient operator.
Remark 6.
The last line in Assumption 4 can be weaken to
$\mathbb{E}_{X\sim\mathcal{L}}\left[\|\nabla_{\theta}^{(m)}s_{\theta}(x)\|_{2}^{4/3+\kappa}\right]\leq C_{3}<\infty,$
for some $\kappa>0$, while all the results keep the same.
Assumption 5 (Regularity on the kernel).
The kernel $K(x,x^{\prime})$ is a measurable function in $\mathcal{X}\times\mathcal{X}$, and is symmetric and positive definite. There exists some universal constant $C_{K}$ such that for all $x,x^{\prime}\in\mathcal{X}$, we have
$K(x,x^{\prime}),\|\nabla_{x}K(x,x^{\prime})\|,\mathrm{Tr}(\nabla_{x}\nabla_{x^{\prime}}K(x,x^{\prime}))\leq C_{K}.$
Appendix E Preliminaries on stable convergence
For our testing procedure presented in Algorithm 1, we calibrate the distribution of the test statistic under the null hypothesis using parametric bootstrap, i.e., resample from the null model conditional on the observed data. Thus, to establish the asymptotic validity of the bootstrap test (SKSD-test), the limit behaviors of the test statistic and the bootstrap statistic need to be analyzed.
In Theorem 2, we establish the weak convergence of the test statistic. However, similar results on the bootstrap statistic $\widetilde{T}_{n}$ do not imply the validity of the test, since $(T_{n},\widetilde{T}_{n}^{(1)},\ldots,\widetilde{T}_{n}^{(B)})$ need to be considered jointly. Therefore, it suffices to study the asymptotic behavior of the conditional quantile function $\mathbb{P}[n\widetilde{T}_{n}\leq t|\mathcal{F}_{n}]$, which is a random variable.
Classical weak convergence results are established for the convergence of the quantile function, which is not sufficient for our purpose in this conditional convergence scenario, and a stronger notion of weak convergence is needed. To this end, we apply the notion of stable convergence [hausler2015stable], which is a stronger form of weak convergence and is suitable for studying the asymptotic behavior of conditional distributions.
Definition 3 (Stable convergence).
Suppose there is a sequence of random variables $(Y_{n})_{n\geq 1}$ defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ with values in a Polish space $(\mathcal{Y},\mathcal{B})$. We say that $Y_{n}$ converges $\mathcal{G}$-stably to a Markov kernel $Q(\omega,\cdot)$ for a sub-$\sigma$-algebra $\mathcal{G}\subseteq\mathcal{F}$, and write as
$Y_{n}\overset{d}{\rightarrow}Q(\omega,\cdot)\quad\mathcal{G}-\text{stably},$
if for every $\mathcal{G}$-measurable random variable $Z\in L_{1}(\mathbb{P})$ and every bounded continuous function $f:\mathcal{Y}\to\mathbb{R}$, we have
$\lim_{n\to\infty}\mathbb{E}[Zf(Y_{n})]=\int_{\Omega}\int_{\mathcal{Y}}Z(\omega)f(y)Q(\omega,\mathrm{d}y)\mathbb{P}(\mathrm{d}\omega).$ (22)
When the limit Markov kernel $Q(\omega,\cdot)$ does not depend on $\omega$, we may omit the dependence on $\omega$ and write the limit as a probability measure. For simplicity, in the manuscript, we also write
$Y_{n}\overset{d}{\to}Y\quad\mathcal{G}-\text{stably}$
for a random variable $Y$ defined on the same probability space, if $Y_{n}$ converges $\mathcal{G}$-stably to the Markov kernel induced by the conditional distribution of $Y|\mathcal{G}$.
Generally speaking, stable convergence requires the weak limit of the sequence $Y_{n}$ keeps “stable” when conditioning on different sub-$\sigma$-algebra $\mathcal{G}$. The following lemma, adapted from Theorem 3.2 in [hausler2015stable], provides an equivalent characterization of stable convergence from this perspective.
Lemma 1.
Suppose the setting in Definition 3 holds. Then $Y_{n}$ converges $\mathcal{G}$-stably to $Y$ is equivalent to
$Y_{n}\overset{d}{\rightarrow}Y\kern 5.0pt\text{under }\mathbb{P}_{F}\text{ for all }F\in\mathcal{G}\text{ with }\mathbb{P}(F)>0,$
where $\mathbb{P}_{F}(\cdot)=\mathbb{P}(\cdot|F)\equiv\mathbb{P}(\cdot\cap F)/\mathbb{P}(F)$.
With Lemma 1, it is clear that if we take $\mathcal{G}=\{\emptyset,\Omega\}$, then stable convergence reduces to the usual weak convergence. In the context of our bootstrap test, we aim to show that the bootstrap statistic converges stably to some limit distribution under different realizations of the data generating process. Thus, by Eqn. (22), the joint distribution of $(T_{n},\widetilde{T}_{n}^{(1)},\ldots,\widetilde{T}_{n}^{(B)})$ can be characterized, which leads to the asymptotic validity of the bootstrap test.
A key technical result applied in the proofs is the stable version of the central limit theorem. The following lemma, adapted from Theorem 6.1 in [hausler2015stable], states the conditions under which a martingale difference array converges stably to a normal distribution:
Lemma 2 (Stable Central Limit Theorem).
Let $(Y_{nk})_{1\leq k\leq k_{n},n\in\mathbb{N}}$ be a square integrable martingale difference array adapted to the array $(\mathcal{F}_{nk})_{1\leq k\leq k_{n},n\in\mathbb{N}}$. Let $\mathcal{G}_{nk}=\cap_{m\geq n}\mathcal{F}_{mk}$ for $n\in\mathbb{N}$ and $1\leq k\leq k_{n}$, and $\mathcal{G}=\sigma(\cup_{n\in\mathbb{N}}\mathcal{G}_{nk_{n}})$. Assume that for some $\mathcal{G}$-measurable real, non-negative random variable $\eta$, as $n\to\infty$,
$\sum_{k=1}^{k_{n}}\mathbb{E}[Y_{nk}^{2}|\mathcal{F}_{n,k-1}]\overset{p}{\to}\eta^{2},$ (N)
and for every $\varepsilon>0$,
$\sum_{k=1}^{k_{n}}\mathbb{E}[Y_{nk}^{2}\bm{1}(|Y_{nk}|\geq\varepsilon)|\mathcal{F}_{n,k-1}]\overset{p}{\to}0.$ (CLB)
The condition (N) and (CLB) can be regarded as the conditional form of moment condition and Lindeberg’s condition. Then
$\sum_{k=1}^{k_{n}}Y_{nk}\overset{d}{\to}\mathcal{N}(0,\eta^{2})\quad\mathcal{G}-\text{stably}.$ (23)
With Lemma 2, we further establish the stable convergence of random processes in Lemma 5 and finally prove the main parametric bootstrap result in Theorem 3.
Appendix F Proofs of results in Section 3
F.1 Proof of Theorem 1
Proof of Theorem 1.
By the definition of the SKSD test statistic, we have
$T_{n}(X,\hat{\theta}_{n})=\sup_{g\in B_{\mathcal{H}^{d}}(1)}\left|\mathbb{E}_{\hat{\mathcal{L}}_{n}}[\mathcal{A}_{p_{\hat{\theta}_{n}}}g(X)]\right|^{2}=\frac{1}{n^{2}}\sum_{i,j=1}^{n}h_{p_{\hat{\theta}_{n}}}(X_{i},X_{j}).$
By Assumption 1, we have $\hat{\theta}_{n}\overset{p}{\rightarrow}\theta_{0}$ as $n\rightarrow\infty$. The term of interest $T_{n}(X,\hat{\theta}_{n})$ can be decomposed as:
$T_{n}(X,\hat{\theta}_{n})=\frac{1}{n^{2}}\sum_{i,j=1}^{n}h_{p_{\theta_{0}}}(X_{i},X_{j})+\frac{1}{n^{2}}\sum_{i,j=1}^{n}\left(h_{p_{\hat{\theta}_{n}}}(X_{i},X_{j})-h_{p_{\theta_{0}}}(X_{i},X_{j})\right).$ (24)
Under the regularity conditions, the law of large numbers for U-statistics [van2000asymptotic, Theorem 12.3] implies that
$\frac{1}{n^{2}}\sum_{i,j=1}^{n}h_{p_{\theta_{0}}}(X_{i},X_{j})\overset{p}{\rightarrow}\mathbb{E}_{\mathcal{L}_{\theta_{0}}}[h_{p_{\theta_{0}}}(X,X^{\prime})]=\mathrm{KSD}^{2}(\mathcal{L}_{\theta_{0}},\mathcal{L}).$ (25)
For the second term, we define $\theta_{t}=t\hat{\theta}_{N}+(1-t)\theta_{0}$ for $t\in[0,1]$. By the fundamental theorem of calculus, we can write
$\frac{1}{n^{2}}\sum_{i,j=1}^{n}\left(h_{p_{\hat{\theta}_{n}}}(X_{i},X_{j})-h_{p_{\theta_{0}}}(X_{i},X_{j})\right)$
$=\frac{1}{n^{2}}\sum_{i,j=1}^{n}\int_{0}^{1}\left\langle\nabla_{\theta}h_{p_{\theta_{t}}}(X_{i},X_{j}),\hat{\theta}_{n}-\theta_{0}\right\rangle\mathrm{d}t$
$\leq\left(\int_{0}^{1}\frac{1}{n^{2}}\sum_{i,j=1}^{n}\left\|\nabla_{\theta}h_{p_{\theta_{t}}}(X_{i},X_{j})\right\|\mathrm{d}t\right)\cdot\|\hat{\theta}_{n}-\theta_{0}\|.$
By the regularity conditions, we have for any $\theta\in\Theta$,
$\mathbb{E}\left[\frac{1}{n^{2}}\sum_{i,j=1}^{n}\left\|\nabla_{\theta}h_{p_{\theta}}(X_{i},X_{j})\right\|\right]$ $=\frac{1}{n}\mathbb{E}\left[\|\nabla_{\theta}h_{p_{\theta}}(X,X)\|\right]+\frac{n-1}{n}\mathbb{E}\left[\|\nabla_{\theta}h_{p_{\theta}}(X,X^{\prime})\|\right]$
$\leq 2(C_{1}C_{2}+C_{1})C_{K}<\infty,$
where $C_{1}$, $C_{2}$ and $C_{K}$ are the constants in the regularity conditions. By Chebyshev’s inequality, the integral term is bounded in probability. Additionally, by Assumption 1, we obtain
$\left\|\hat{\theta}_{n}-\theta_{0}\right\|=\left\|\frac{1}{n}\sum_{i=1}^{n}I(X_{i},\theta_{0})\right\|+O_{P}(1/\sqrt{n})=o_{p}(1).$
Combining the above results, we conclude that
$\frac{1}{n^{2}}\sum_{i,j=1}^{n}\left(h_{p_{\hat{\theta}_{n}}}(X_{i},X_{j})-h_{p_{\theta_{0}}}(X_{i},X_{j})\right)=O_{p}(1)\cdot o_{p}(1)=o_{p}(1).$ (26)
Finally, Plugging Eqn. (25) and (26) into Eqn. (24) implies that
$T_{n}(X,\hat{\theta}_{n})\overset{p}{\rightarrow}\mathrm{KSD}^{2}(\mathcal{L}_{\theta_{0}},\mathcal{L}).$
F.2 Proof of Theorem 2
Proof of Theorem 2.
Theorem 2 can be proved by following the strategy mentioned in the main text. We refer the audience to the proofs of Theorem 3 or Corollary 2 as natural extensions of this proof strategy.
Here, we provide a specific viewpoint to regard the test statistic under the null is as a special realization of the bootstrap (resampling) setting in Algorithm 1 in the following sense:
[…]
where We choose $\mathcal{F}_{n}=\{\phi,\Omega\}$, which $\Omega$ is defined to be the whole sample space. Then, the proof of Corollary 2 can be directly adapted to the current setting. The result from Step 1 in the proof of Corollary 2 tells us that the asymptotic variational form of the test statistic can be written as
$nT_{n}=\sup_{f\in\mathcal{B}_{\mathcal{H}^{d}}(1)}S_{n}(f)^{2}+o_{p}(1),$
where the functional $S_{n}(f)$ is defined as
$S_{n}(f)$ $:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}S^{*}(f,X_{i},\theta_{0})$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left[\mathcal{A}_{\theta_{0}}f(X_{i})+\Big\langle\mathbb{E}_{X^{\prime}\sim P_{\theta_{0}}}\left[[\nabla_{\theta}s_{\theta_{0}}(X^{\prime})]^{\top}f(X^{\prime})\right],I(X_{i},\theta_{0})\Big\rangle\right],$ (27)
and $\mathcal{B}_{\mathcal{H}^{d}}(1)$ is the unit ball in $\mathcal{H}^{d}$ and $S^{*}(f,X,\theta)$ is defined in Eqn (9).
Continuing with the analysis in the Step 2 in the proof of Corollary 2, we again verify that
$\{S_{n}(f)\}_{f\in\mathcal{H}}\overset{\mathbb{P}_{\theta_{0}}}{\to}\{S(f)\}_{f\in\mathcal{H}}\overset{d}{=}W\quad\text{ as }n\to\infty.$
Thus, for any $\omega\in\Omega$, we have
$\limsup_{n\rightarrow\infty}\left|\mathbb{P}_{H_{0}}[nT_{n}\leq t|\{\phi,\Omega\}](\omega)-\mathbb{P}[W\leq t]\right|=0\quad\forall t\in\mathbb{R}.$
Dropping the redundant terms, we have
$\limsup_{n\rightarrow\infty}\left|\mathbb{P}_{H_{0}}[nT_{n}\leq t]-\mathbb{P}[W\leq t]\right|=0\quad\forall t\in\mathbb{R},$
which is exactly the conclusion of Theorem 2. ∎
F.3 Proof of Theorem 3
Proof of Theorem 3.
Before going into the proof details, we first recall the definition of $S^{*}$ in Eqn (9):
$S^{*}(f,X,\theta)=\mathcal{A}_{\theta}f(X)+\Big\langle\mathbb{E}_{X^{\prime}\sim P_{\theta}}\left[[\nabla_{\theta}s_{\theta}(X^{\prime})]^{\top}f(X^{\prime})\right],I(X,\theta)\Big\rangle.$
The moment properties of $S^{*}(f,X,\theta)$ we used in what follows are summarized as Lemma 3 with proof provided in Appendix H.1.
Lemma 3.
$S^{*}(f,X,\theta)$ satisfies the moment conditions below. For any $\theta\in\Theta$, $X\sim\mathbb{P}_{\theta}$ and $f\in\mathcal{H}^{d}$, $S^{*}(f,X,\theta)$ satisfies
  1. (1)
    $\mathbb{E}_{X\sim\mathbb{P}_{\theta}}S^{*}(f,X,\theta)=0$.
  2. (2)
    $\mathbb{E}_{X\sim\mathbb{P}_{\theta}}[S^{*}(f,X,\theta)^{2}]\lesssim\mathbb{E}_{X\sim\mathbb{P}_{\theta}}[\|f(X)\|^{2}]+\left\|\mathbb{E}_{X\sim\mathbb{P}_{\theta}}\left[[\nabla_{\theta}s_{\theta}(X)]^{\top}f(X)\right]\right\|^{2},$ which is uniformly bounded in $\theta\in\Theta$.
Now we are ready to prove Theorem 3. We split the proof into three steps.
Step 1: Asymptotic expansion.
We first derive asymptotic variational form of the test statistic $nT_{n}$, which is essential for the following proof on the weak convergence of the bootstrap distribution. The result is summarized in the following lemma.
Lemma 4.
Introduce the following random functionals:
$\widetilde{S}_{n}(f)$ $\equiv\frac{1}{\sqrt{n}}\sum_{i=1}^{n}S^{*}(f,\widetilde{X}_{i},\hat{\theta}_{n}).$
The asymptotic variational form of the bootstrap statistic can be written as
$n\widetilde{T}_{n}=\sup_{f\in\mathcal{B}_{\mathcal{H}^{d}}(1)}\widetilde{S}_{n}(f)^{2}+o_{p}(1),$
The proof of Lemma 4 can be found in Appendix H.2. Lemma 4 implies that two statistics: $n\widetilde{T}_{n}$ and $\sup_{f\in\mathcal{B}_{\mathcal{H}^{d}}(1)}\widetilde{S}_{n}(f)^{2}$ are asymptotically equivalent. Notice that the statement of Theorem 3 is equivalent to as $n\to\infty$,
$\mathbb{P}\left[n\tilde{T}_{n}\leq t\big|\mathcal{F}_{n}\right]\overset{p}{\to}\mathbb{P}[W\leq t]\quad\forall t\in\mathbb{R},$
which is the goal of the proof. By Lemma 20, the conditional tail probability of these two statistics are also asymptotic equivalent, i.e.,
$\mathbb{P}\left[n\tilde{T}_{n}\leq t\big|\mathcal{F}_{n}\right]=\mathbb{P}\left[\sup_{f\in\mathcal{B}_{\mathcal{H}^{d}}(1)}\widetilde{S}_{n}(f)^{2}\leq t\Big|\mathcal{F}_{n}\right]+o_{p}(1)\quad\forall t\in\mathbb{R}.$
Thus, it suffices to prove
$\mathbb{P}\left[\sup_{f\in\mathcal{B}_{\mathcal{H}^{d}}(1)}\widetilde{S}_{n}(f)^{2}\leq t\Big|\mathcal{F}_{n}\right]\overset{p}{\to}\mathbb{P}[W\leq t]\quad\forall t\in\mathbb{R}.$ (28)
In the following proof, we focus on proving Eqn. (28).
Step 2: Weak convergence in the functional space.
Our proof relies on the concept of regular conditional distribution and convergence of Markov kernel, especially stable convergence. We provide introduction and basic results on the concept of stable convergence in Appendix E. See [hausler2015stable] for a detailed introduction to this framework.
We prove the desired stable convergence result through the following lemma, which is a generalization of [fernandez2024general, Theorem 1 ]. We first introduce some notation. Define $\mathcal{F}=\sigma(\cup_{n=1}^{\infty}\mathcal{F}_{n})$ to be the smallest $\sigma$-algebra generated by the whole data sequence. Notice that $\widetilde{T}_{n}$ and $\widetilde{S}_{n}$ both depend on the data through $\hat{\theta}_{n}$, which only depends $\{X_{i}\}_{i=1}^{n}$. Based on this observation, deriving the convergence results conditioning on $\mathcal{F}$ is equivalent to conditioning on $\mathcal{F}_{n}$. Thus, for simplicity, we work on the $\mathcal{F}$-stable convergence of the test statistic in the proof in this part.
Lemma 5 (Functional stable convergence).
Suppose we have a sequence of bounded linear random functionals $\{\widetilde{S}_{n}\}_{n=1}^{\infty}$, which are maps from $\mathcal{H}^{d}\to\mathbb{R}$. Suppose the following three conditions hold:
  1. (1)
    There exists a functional $S(f)$ and a covariance operator $\sigma(f_{1},f_{2})$ such that for any $f\in\mathcal{H}$, we have
    $\widetilde{S}_{n}(f)\overset{d}{\to}S(f)\sim\mathcal{N}(0,\sigma(f,f))\quad\mathcal{F}-\text{stably},$ (C1.1)
  2. (2)
    For some orthonormal basis $(\phi_{i})_{i=1}^{\infty}$ of RKHS $\mathcal{H}^{d}$, we have
    $\sum_{i=1}^{\infty}\sigma(\phi_{i},\phi_{i})<\infty.$ (C1.2)
  3. (3)
    For some orthonormal basis $(\phi_{i})_{i=1}^{\infty}$ of RKHS $\mathcal{H}^{d}$ and any $\varepsilon>0$, we have
    $\lim_{k\to\infty}\limsup_{n\to\infty}\mathbb{P}\left(\sum_{i=k}^{\infty}\widetilde{S}_{n}(\phi_{i})^{2}\geq\varepsilon\,\bigg|\,F\right)=0,$ (C1.3)
    for any $F\subseteq\mathcal{F}$ with $\mathcal{L}(F)>0$, where the measure $\mathcal{L}(\cdot)$ is taken over data generation process.
Then we have
$\{\widetilde{S}_{n}(f)^{2}\}_{f\in\mathcal{H}^{d}}\overset{d}{\to}\{S(f)^{2}\}_{f\in\mathcal{H}^{d}}\quad\mathcal{F}-\text{stably}.$
The proof of Lemma 5 is provided in Appendix H.3. We postpone the verification of conditions (C1.1)-(C1.3) to Appendix F.7. With Lemma 5 in hand, we conclude, as $n\to\infty$,
$\{\widetilde{S}_{n}(f)^{2}\}_{f\in\mathcal{H}^{d}}\overset{d}{\to}\{S(f)^{2}\}_{f\in\mathcal{H}^{d}}\quad\mathcal{F}-\text{stably}.$ (29)
Define the operator norm from $\mathcal{H}^{d}$ to $\mathbb{R}$ as, for any operator $T:\mathcal{H}^{d}\to\mathbb{R}$,
$\|T\|_{\mathcal{H}^{d}\to\mathbb{R}}^{2}=\sup_{\|f\|_{\mathcal{H}^{d}}\leq 1}T(f)^{2}.$ (30)
By continuous mapping theorem, we have
$\sup_{f\in\mathcal{B}_{\mathcal{H}^{d}}(1)}\widetilde{S}_{n}(f)^{2}=\|\widetilde{S}_{n}\|_{\mathcal{H}^{d}\to\mathbb{R}}^{2}\overset{d}{\to}\|S\|_{\mathcal{H}^{d}\to\mathbb{R}}^{2}=\sup_{f\in\mathcal{B}_{\mathcal{H}^{d}}(1)}S(f)^{2}\quad\mathcal{F}-\text{stably}.$
Since the closed-form expression for $S(f)$ is given, the above limit can be written as an infinite mixture of independent $\chi^{2}$ random variables. Define random variable $W$ as:
$W=\sum_{u=1}^{\infty}\lambda_{u}Z_{u}^{2},$
where $\lambda_{u}$’s are eigenvalues of operator $T_{\sigma}$ (See Eqn. (38) for the definition) and $Z_{u}$’s are independent standard Gaussian random variables. See Proposition 2 for the existence of $W$. By direct calculation, we have
$\sup_{f\in\mathcal{B}_{\mathcal{H}^{d}}(1)}S(f)^{2}=\sup_{\|f\|_{\mathcal{H}^{d}}\leq 1}\left\langle\sum_{i=1}^{\infty}\sqrt{\lambda_{i}}Z_{i}\phi_{i},f\right\rangle_{\mathcal{H}^{d}}=\sum_{i=1}^{\infty}\lambda_{i}Z_{i}^{2}\overset{d}{=}W.$
Step 3: Convergence of the tail probability.
Finally, we show that the stable convergence of the test statistic implies the convergence of quantiles in probability, which is exactly Eqn. (28).
We use the definition of stable convergence in [hausler2015stable, Page 6 ] to prove it. Recall Definition of the operator norm in Eqn. (30). Let us choose
$g_{1}\in L_{1}(\mathcal{L}^{\infty})\text{ and measurable w.r.t. }\mathcal{F}\kern 5.0pt\text{ and }\kern 5.0ptg_{2}(\cdot)=\bm{1}(\,\cdot\leq t).$
where $f$ is specified later. The definition of stable convergence together with Eqn. (29) tells that
$\lim_{n\to\infty}\mathbb{E}\left[g_{1}(\mathcal{D})\mathbb{E}\left[g_{2}\left(\|\widetilde{S}_{n}\|_{\mathcal{H}^{d}\to\mathbb{R}}^{2}\right)\big|\mathcal{F}_{n}\right]\right]=\mathbb{E}[g_{1}(\mathcal{D})]\mathbb{E}\left[g_{2}(\|S\|_{\mathcal{H}^{d}\to\mathbb{R}}^{2})\right],$
which is equivalent to
$\lim_{n\to\infty}\mathbb{E}\Big[g_{1}(\mathcal{D})\mathbb{P}\left[\|\widetilde{S}_{n}\|_{\mathcal{H}^{d}\to\mathbb{R}}^{2}\leq t|\mathcal{F}_{n}\right]\Big]=\mathbb{E}[g_{1}(\mathcal{D})]\mathbb{P}[W\leq t],$ (31)
where $\mathcal{D}$ is the data sequence in $\mathcal{X}^{\infty}$. To get the final result, we need a useful technical lemma. The proof is provided in Appendix H.4.
Lemma 6.
Suppose for all bounded continuous function $f$, we have
$\lim_{n\to\infty}\mathbb{E}[f(X_{n})\cdot X_{n}]=C\cdot\lim_{n\to\infty}\mathbb{E}[f(X_{n})],$ (32)
where $C$ is a numerical constant. Then we have $X_{n}\overset{p}{\to}C$.
Since $\mathbb{P}(\cdot|\mathcal{F}_{n})$ is measurable with respect to $\mathcal{F}_{n}$ thus $\mathcal{F}$, we can choose $g_{1}(\mathcal{D})$ to be any measurable functions of the regular conditional probability $\mathbb{P}\left[\|\widetilde{S}_{n}\|_{\mathcal{H}^{d}\to\mathbb{R}}^{2}\leq t|\mathcal{F}_{n}\right]$ in Eqn. (31). Since by our definition, $g_{1}$ covers all $\mathcal{F}$-measurable functions in $L_{1}(\mathcal{L}^{\infty})$, it also covers all bounded continuous function. Applying Lemma 6, we have
$\mathbb{P}\left[\|\widetilde{S}_{n}\|_{\mathcal{H}^{d}\to\mathbb{R}}^{2}\leq t|\mathcal{F}_{n}\right]=\mathbb{P}\left[\sup_{\|f\|_{\mathcal{H}^{d}}\leq 1}\widetilde{S}_{n}(f)^{2}\leq t\bigg|\mathcal{F}_{n}\right]\overset{p}{\rightarrow}\mathbb{P}[W\leq t].$
Noticing that we have no constraint on $t$ in the proof, the above equation proves Eqn. (28), thus proves Theorem 3. ∎
F.4 Proof of Theorem 4
Proof of Theorem 4.
Theorem 4 consists of two parts: the asymptotic validity under the null hypothesis and the consistency under the alternative hypothesis. We provide the proofs of two parts separately.
Proof of asymptotic validity under the null hypothesis.
We use the results in Theorem 2 and 3 combining with [niu2022reconciling, Lemma 3 ] to prove the asymptotic validity of the test. We state their lemma here for completeness.
Lemma 7.
Consider two hypothesis tests based on the same test statistic $T_{1,n}$ but different critical values:
$\phi_{n}^{1}\equiv\mathbbm{1}(T_{1,n}>T_{2,n});\quad\phi_{n}^{2}\equiv\mathbbm{1}(T_{1,n}>w_{1-\alpha}).$
If the critical value of the first converges in probability to that of the second:
$T_{2,n}\overset{p}{\to}w_{1-\alpha},$ (33)
and the test statistic does not accumulate near the limiting critical value:
$\lim_{\delta\to 0}\lim_{n\to\infty}\mathbb{P}(|T_{1,n}-w_{1-\alpha}|<\delta)=0,$ (34)
then the two tests are asymptotically equivalent:
$\lim_{n\to\infty}\mathbb{P}[\phi_{n}^{1}=\phi_{n}^{2}]=1.$
To apply this lemma, we denote the $(1-\alpha)$-quantile of the distribution of $W$ as $w_{1-\alpha}$. Consider the case that $T_{1,n}=nT_{n}$ defined in Eqn. (SKSD) and $T_{2,n}=n\mathbb{Q}_{1-\alpha}(\widetilde{T}_{n})$. Now we can check two conditions in the lemma.
To check Eqn. (33), we first directly apply the result from Theorem 3 to obtain
$\mathbb{P}[n\tilde{T}_{n}\leq t|\mathcal{F}_{n}]\overset{p}{\to}\mathbb{P}[W\leq t]\quad\forall t\in\mathbb{R},$
which is equivalent to $n\widetilde{T}_{n}|\mathcal{F}_{n}\overset{d,p}{\to}W$ with notation used in [niu2022reconciling]. Applying Lemma 19, we have
$\mathbb{Q}_{1-\alpha}(n\widetilde{T}_{n})\overset{p}{\to}\mathbb{Q}_{1-\alpha}(W)=w_{1-\alpha}.$
To check Eqn. (34), we notice that the limit distribution of $T_{1,n}=nT_{n}$ is a mixture of $\chi$-squared distributions, which is a continuous distribution. Define the density of $W$ as $p_{W}(w)$ and $C_{W}=\sup_{w\in[w_{1-\alpha}-1,w_{1-\alpha}+1]}p_{W}(w)$. Then we have
$\lim_{\delta\to 0}\lim_{n\to\infty}\mathbb{P}(|nT_{n}-w_{1-\alpha}|<\delta)\leq\lim_{\delta\to 0}2C_{W}\delta=0.$
In conclusion, Lemma 7 gives us the asymptotic equivalence between the proposed test (SKSD-test) and the test $\phi_{n,\alpha}$ which is asymptotically valid by Theorem 3. Thus, we conclude the asymptotic validity of the test $\phi_{n,\alpha}$, i.e.,
$\lim_{n\rightarrow\infty}\mathbb{P}_{H_{0}}[\phi_{n,\alpha}=1]=\alpha.$
Proof of consistency under the alternative hypothesis.
We first use the following statement to characterize the limit behavior of statistic $T_{n}$ under the alternative hypothesis $H_{1}$. For any sequence of estimator $\hat{\theta}_{n}\in\Theta$, there exists $\varepsilon>0$ such that
$\lim_{n\rightarrow\infty}\mathbb{P}_{H_{1}}\left(T_{n}>\varepsilon\right)=1.$ (35)
Proof of Statement (35).
In the proof, we explicitly choose
$\varepsilon=\frac{1}{2}\inf_{\theta\in\Theta}\mathrm{KSD}(\mathcal{L}_{\theta},\mathcal{L}).$
By the condition $\inf_{\theta\in\Theta}\mathrm{KSD}(\mathcal{L}_{\theta},\mathcal{L})>0$, we confirm that $\varepsilon>0$. We decompose the target probability in the following way:
$\mathbb{P}_{H_{1}}\left(T_{n}\geq\varepsilon\right)$ $=\mathbb{P}_{H_{1}}\left(\mathrm{KSD}(\mathcal{L}_{\hat{\theta}_{n}},\mathcal{L})+(T_{n}-\mathrm{KSD}(\mathcal{L}_{\hat{\theta}_{n}},\mathcal{L}))>\varepsilon\right)$
$\geq\mathbb{P}_{H_{1}}\left(2\varepsilon+(T_{n}-\mathrm{KSD}(\mathcal{L}_{\hat{\theta}_{n}},\mathcal{L}))>\varepsilon\right)$
$=\mathbb{P}_{H_{1}}\left(T_{n}-\mathrm{KSD}(\mathcal{L}_{\hat{\theta}_{n}},\mathcal{L})>-\varepsilon\right)$
$\geq\mathbb{P}_{H_{1}}\left(\left|T_{n}-\mathrm{KSD}(\mathcal{L}_{\hat{\theta}_{n}},\mathcal{L})\right|<\varepsilon\right),$
where in the first inequality, we use the fact that $\mathrm{KSD}(\mathcal{L}_{\hat{\theta}_{n}},\mathcal{L})\geq\inf_{\theta\in\Theta}\mathrm{KSD}(\mathcal{L}_{\theta},\mathcal{L})=2\varepsilon$. The last expression can be bounded by the following lemma, which is indeed a stronger result. The proof of this result can be found in [matsubara2022robust, Lemma 2 ].
Lemma 8.
With Assumption 3 and 4, we have
$\lim_{n\to\infty}\sup_{\theta\in\Theta}|\mathrm{KSD}(\mathcal{L}_{\theta},\mathcal{L}^{n})-\mathrm{KSD}(\mathcal{L}_{\theta},\mathcal{L})|=0\quad\text{almost surely},$
where $\mathcal{L}^{n}$ is the empirical distribution of $X_{1},\ldots,X_{n}$.
Noticing that $\mathrm{KSD}(\mathcal{L}_{\hat{\theta}_{n}},\mathcal{L}^{n})=T_{n}$ and that the almost sure convergence implies convergence in probability, we have
$\lim_{n\rightarrow\infty}\mathbb{P}_{H_{1}}\left(T_{n}\geq\varepsilon\right)=\lim_{n\rightarrow\infty}\mathbb{P}_{H_{1}}\left(\left|T_{n}-\mathrm{KSD}(\mathcal{L}_{\hat{\theta}_{n}},\mathcal{L})\right|<\varepsilon\right)=1.$
With the help of Lemma 7 and the same analysis in the proof of validity above, we can get an asymptotically equiavlent test $\phi^{\prime}_{n}=\mathbbm{1}(nT_{n}>w_{1-\alpha})$. By Eqn. (35), we have
$\lim_{n\to\infty}\mathbb{P}_{H_{1}}\left(nT_{n}>w_{1-\alpha}\right)\geq\lim_{n\to\infty}\mathbb{P}_{H_{1}}\left(nT_{n}>n\varepsilon\right)=1.$
Thus, the asymptotic power can be calculated as
$\lim_{n\rightarrow\infty}\mathbb{P}_{H_{1}}[\phi_{n,\alpha}=1]=\lim_{n\rightarrow\infty}\mathbb{P}_{H_{1}}[\phi^{\prime}_{n}=1]=1.$
F.5 Proof of Proposition 2
Proof of Proposition 2.
We will verify the conditions in the Kolmogorov’s three series theorem [durrett2019probability, Page 150]. By choosing the truncation threshold $A$ in the theorem to be $\infty$, the first condition in the three series holds automatically. Thus, it suffices to show that
$\sum_{i=1}^{\infty}\mathbb{E}[\lambda_{i}Z_{i}^{2}]<\infty,\quad\sum_{i=1}^{\infty}\mathrm{Var}[\lambda_{i}Z_{i}^{2}]<\infty.$
In the proof of Theorem 3, we have shown that $\lambda_{i}\geq 0$ and $\sum_{i=1}^{\infty}\lambda_{i}=\text{Trace}(T_{\sigma})=\sum_{i=1}^{\infty}\sigma(\phi_{i},\phi_{i})<\infty$. Since $Z_{i}^{2}$ is a chi-squared random variable with degree 1, we have $\mathbb{E}[\lambda_{i}Z_{i}^{2}]=\lambda_{i}$ and $\mathrm{Var}[\lambda_{i}Z_{i}^{2}]=2\lambda_{i}^{2}$, which implies the later two conditions. ∎
F.6 Proof of Corollary 2
Proof of Corollary 2.
Eqn. (18) is a direct consequence of Theorem 3 and Lemma 18. Thus, we omit the proof here, and focus on proving Eqn. (19) as follows.
Define $\Omega$ to be data sample space, i.e., $\Omega=\mathcal{X}^{\mathbb{N}}=\{(x_{1},x_{2},\cdots):x_{i}\in\mathcal{X}\}$. Define an event $\mathcal{E}=\{\omega\in\Omega:\lim_{n\rightarrow\infty}\hat{\theta}_{n}(\omega)\rightarrow\theta_{0}\}$. By our additional assumption, we have $\mathbb{P}(\mathcal{E})=1$. Thus we can work under this event.
For a fixed $\omega\in\mathcal{E}$, $\hat{\theta}_{n}$ is a fixed sequence in $\Theta$ converging to $\theta_{0}$. We aim to prove that
$\lim_{n\rightarrow\infty}\mathbb{P}[n\tilde{T}_{n}\leq t|\mathcal{F}_{n}](\omega)=\mathbb{P}[W\leq t].$
Under this fixed $\omega$, the data generation processes of $\widetilde{X}_{i}$ is a fixed sequence of distributions, denoted as $\mathbb{P}_{\hat{\theta}_{n}}$. With this notation, we can equivalently write the target in a clearer form:
[…]
Step 1: Asymptotic expansion.
The step 1 in this proof is a correspondence to the step 1 of the proof of Theorem 3. Instead of a decomposition with respect to probability $\mathcal{L}^{n}\times\mathbb{P}_{\hat{\theta}_{n}}^{n}$, we need to work with decomposition with respect to conditional probability.
Lemma 9.
Conditional on data sequence $\mathcal{D}_{n}$ (i.e., for a fixed sequence of estimator $\hat{\theta}_{n}$), the bootstrap statistic admits the following asymptotic variational form:
$n\widetilde{T}_{n}=\sup_{f\in\mathcal{B}_{\mathcal{H}}(1)}\widetilde{S}_{n}(f)^{2}+o_{\mathbb{P}_{\hat{\theta}_{n}}}(1),$
where $\widetilde{S}_{n}(f)$ is defined in Lemma 4.
The proof of Lemma 9 is similar to the proof of Lemma 4, which we will not repeat here. Lemma 9 implies that $n\widetilde{T}_{n}$ and $\sup_{f\in\mathcal{B}_{\mathcal{H}}(1)}\widetilde{S}_{n}(f)^{2}$ are asymptotic equivalent under conditional measure $\mathbb{P}_{\hat{\theta}_{n}}$. By Slutsky’s Theorem and Eqn. (36), it suffices to prove
$\sup_{f\in\mathcal{B}_{\mathcal{H}}(1)}\widetilde{S}_{n}(f)^{2}\overset{\mathbb{P}_{\hat{\theta}_{n}}}{\rightarrow}W\quad\text{ as }n\to\infty.$ (37)
Step 2: Weak convergence in the functional space.
We directly prove Eqn. (37) in this part. [fernandez2024general, Theorem 1 ] is used to prove this result. We state it here for convenience.
Lemma 10.
With the same setup in Lemma 5, we introduce the new conditions:
  1. (1)’
    There exists a covariance functional $\sigma(f_{1},f_{2})$ such that for any $f\in\mathcal{H}$, we have
    $\widetilde{S}_{n}(f)\overset{d}{\to}S(f)\sim\mathcal{N}(0,\sigma(f,f)),$ (C2.1)
  2. (3)’
    For some orthonormal basis $(\phi_{i})_{i=1}^{\infty}$ of RKHS $\mathcal{H}$ and any $\varepsilon>0$, we have
    $\lim_{k\to\infty}\limsup_{n\to\infty}\mathbb{P}_{\hat{\theta}_{n}}\left(\sum_{i=k}^{\infty}\widetilde{S}_{n}(\phi_{i})^{2}\geq\varepsilon\right)=0,$ (C2.3)
Suppose the conditions (C2.1), (C1.2) and (C2.3) hold, then we have
$\{\widetilde{S}_{n}(f)^{2}\}_{f\in\mathcal{H}}\overset{d}{\to}\{S(f)^{2}\}_{f\in\mathcal{H}}.$
The proof of Lemma 10 is similar to the proof of Lemma 5, which we will not repeat here. Notice that (C1.2) is not a probabilistic statement, and it is verified in the proof of Theorem 3 and still holds in this case. The verification of (C2.1) is very similar the verification of (C1.1) in the proof of Theorem 3. We can check that, as $n\to\infty$,
$\mathbb{E}[S^{*}(f,\widetilde{X},\hat{\theta}_{n})^{2}]\to\sigma(f,f)\qquad\text{and}$
$\mathbb{E}\left[S^{*}(f,\widetilde{X},\hat{\theta}_{n})^{2}\bm{1}\left(S^{*}(f,\widetilde{X},\hat{\theta}_{n})\geq\varepsilon\sqrt{n}\right)\right]\to 0$
still hold for any $\varepsilon>0$. Then by standard Lindeberg-Feller Central Limit Theorem [durrett2019probability, Theorem 3.4.10 ], we verify Condition (C2.1).
For Condition (C2.3), we can verify it by the Markov inequality: For the third condition, we use the Markov’s inequality to obtain as $n\to\infty$,
$\mathbb{P}_{\hat{\theta}_{n}}\left(\sum_{k\geq i}\widetilde{S}_{n}(\phi_{k})^{2}\geq\varepsilon\right)\leq\frac{1}{\varepsilon}\sum_{k=i}^{\infty}\mathbb{E}_{\hat{\theta}_{n}}\left[\widetilde{S}_{n}(\phi_{k})^{2}\right]\to\frac{1}{\varepsilon}\sum_{k=i}^{\infty}\sigma(\phi_{k},\phi_{k}).$
Since $\sigma(\phi_{k},\phi_{k})>0$ is summable by Condition (C1.2), we obtain
$\lim_{i\to\infty}\sum_{k=i}^{\infty}\sigma(\phi_{k},\phi_{k})=0,$
which leads to Condition (C2.3):
$\lim_{i\to\infty}\limsup_{n\to\infty}\mathbb{P}_{\theta_{0}}\left(\sum_{k\geq i}S_{n}(\phi_{k})^{2}\geq\varepsilon\right)=\lim_{i\to\infty}\sum_{k=i}^{\infty}\sigma(\phi_{k},\phi_{k})=0.$
After verifying the three conditions, we can directly apply Lemma 10 to obtain
$\{\widetilde{S}_{n}(f)^{2}\}_{f\in\mathcal{H}}\overset{d}{\to}\{S(f)^{2}\}_{f\in\mathcal{H}}.$
As it is shown in the Step 3 in the proof of Theorem 3, the limit distribution of the variational form can be expressed as
$\sup_{\|f\|_{\mathcal{H}}\leq 1}S(f)^{2}=\sum_{i=1}^{\infty}\lambda_{i}Z_{i}^{2}\overset{d}{=}W.$
Then by the continuous mapping theorem, we obtain
$\sup_{\|f\|_{\mathcal{H}}\leq 1}\widetilde{S}_{n}(f)^{2}\overset{\mathbb{P}_{\hat{\theta}_{n}}}{\rightarrow}\sup_{\|f\|_{\mathcal{H}}\leq 1}S(f)^{2}\overset{d}{=}W\quad\text{ as }n\to\infty,$
which proves Eqn. (37), thus completes the proof of Corollary 2.
F.7 Verification of the conditions in Lemma 5
F.7.1 Proof of Condition (C1.1)
We mainly use Lemma 2 to prove this result, which is stated in Appendix E with background of stable convergence for reader’s convenience.
We first explicitly define the $S(f)$ and recall the definition of $\sigma(f_{1},f_{2})$ on the right hand side of Eqn. (C1.1):
$S(f)=\sum_{i=1}^{\infty}\sqrt{\lambda_{i}}\langle\phi_{i},f\rangle_{\mathcal{H}^{d}}\,Z_{i}\quad\text{and}\quad\sigma(f_{1},f_{2})=\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}[S^{*}(f_{1},X,\theta_{0})S^{*}(f_{2},X,\theta_{0})],$
where $\lambda_{i}$’s are the eigenvalues of the operator $T_{\sigma}:\mathcal{H}^{d}\to\mathcal{F}_{d,1}$ defined as
$T_{\sigma}(f)(x)=\sigma(f,\mathbf{K}_{x})\quad\text{for any }f\in\mathcal{H}^{d},x\in\mathcal{X},$ (38)
where $K_{x}$ is the feature map, $\mathbf{K}_{x}=K_{x}\bm{1}_{d}$ and $Z_{i}$’s are independent standard normal random variables. In the following proof, we stick to the notation $\widetilde{X}$ generated from $\mathbb{P}_{\hat{\theta}_{n}}$ without additional explanation.
First notice a direct consequence of Lemma 3 is that for any $f\in\mathcal{H}^{d}$,
$\mathbb{E}\left[S^{*}(f,\widetilde{X},\hat{\theta}_{n})|\mathcal{F}_{n}\right]=0,$ (39)
which ensures that we can fit our setting into the martingale framework in Lemma 2 and apply its result. Thus, we only need to verify the conditional moment condition (N) and conditional Lindeberg’s condition (CLB).
Verification of Condition (N)
To verify the conditional moment condition (N), we need to show for any $f\in\mathcal{H}$, as $n\to\infty$,
$\mathbb{E}[S^{*}(f,\widetilde{X},\hat{\theta}_{n})^{2}|\mathcal{F}_{n}]\overset{p}{\to}\sigma(f,f)=\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}[S^{*}(f,X,\theta_{0})^{2}]<\infty,$ (40)
by Lemma 3. Actually, we have
$\mathbb{E}[S^{*}(f,\widetilde{X},\hat{\theta}_{n})^{2}|\mathcal{F}_{n}]$
$=\int_{\mathcal{X}}S^{*}(f,x,\hat{\theta}_{n})S^{*}(f,x,\hat{\theta}_{n})p_{\hat{\theta}_{n}}(x)\lambda(\mathrm{d}x)$
$\overset{p}{\to}\int_{\mathcal{X}}S^{*}(f,x,\theta_{0})S^{*}(f,x,\theta_{0})p_{\theta_{0}}(x)\lambda(\mathrm{d}x)=\sigma(f,f),$ (41)
where the last line follows from the continuous mapping theorem with the convergence of $\hat{\theta}_{n}\to\theta_{0}$ in probability.
Verification of Condition (CLB)
To verified the conditional Lindeberg’s condition, we need to show that for any $f\in\mathcal{H}$ and any $\varepsilon>0$, as $n\to\infty$,
$\mathbb{E}\left[S^{*}(f,\widetilde{X},\hat{\theta}_{n})^{2}\bm{1}\left(S^{*}(f,\widetilde{X},\hat{\theta}_{n})\geq\varepsilon\sqrt{n}\right)\Big|\mathcal{F}_{n}\right]\overset{p}{\to}0.$ (42)
Through the same argument as in Eqn. (41), we have for $n>n_{0}$,
$\mathbb{E}\left[S^{*}(f,\widetilde{X},\hat{\theta}_{n})^{2}\bm{1}\left(S^{*}(f,\widetilde{X},\hat{\theta}_{n})\geq\varepsilon\sqrt{n}\right)\Big|\mathcal{F}_{n}\right]$
$\leq\mathbb{E}\left[S^{*}(f,\widetilde{X},\hat{\theta}_{n})^{2}\bm{1}\left(S^{*}(f,\widetilde{X},\hat{\theta}_{n})\geq\varepsilon\sqrt{n_{0}}\right)\Big|\mathcal{F}_{n}\right]$
$\overset{p}{\to}\mathbb{E}_{\mathbb{P}_{\theta_{0}}}\left[S^{*}(f,X,\theta_{0})^{2}\bm{1}\left(S^{*}(f,X,\theta_{0})\geq\varepsilon\sqrt{n_{0}}\right)\right].$ (43)
By Lemma 3, we know that $\mathbb{E}_{\mathbb{P}_{\theta}}[S^{*}(f,X,\theta)^{2}]$ is uniformly bounded almost surely. For any $\varepsilon,\delta>0$, we can find a $n_{0}$ large enough such that
$\mathbb{E}_{\mathbb{P}_{\theta_{0}}}\left[S^{*}(f,X,\theta_{0})^{2}\bm{1}\left(S^{*}(f,X,\theta_{0})\geq\varepsilon\sqrt{n_{0}}\right)\right]<\frac{\delta}{2}.$
Thus, we have
$\mathbb{P}\left(\mathbb{E}\left[S^{*}(f,\widetilde{X},\hat{\theta}_{n})^{2}\bm{1}\left(S^{*}(f,\widetilde{X},\hat{\theta}_{n})\geq\varepsilon\sqrt{n}\right)\Big|\mathcal{F}_{n}\right]\geq\delta\right)$
$\leq\mathbb{P}\left(\mathbb{E}\left[S^{*}(f,\widetilde{X},\hat{\theta}_{n})^{2}\bm{1}\left(S^{*}(f,\widetilde{X},\hat{\theta}_{n})\geq\varepsilon\sqrt{n_{0}}\right)\Big|\mathcal{F}_{n}\right]\geq\delta\right)$
$\leq\mathbb{P}\bigg(\mathbb{E}\left[S^{*}(f,\widetilde{X},\hat{\theta}_{n})^{2}\bm{1}\left(S^{*}(f,\widetilde{X},\hat{\theta}_{n})\geq\varepsilon\sqrt{n_{0}}\right)\Big|\mathcal{F}_{n}\right]$
$\qquad\qquad\qquad\qquad-\mathbb{E}_{\mathbb{P}_{\theta_{0}}}\left[S^{*}(f,X,\theta_{0})^{2}\bm{1}\left(S^{*}(f,X,\theta_{0})\geq\varepsilon\sqrt{n_{0}}\right)\right]>\delta/2\bigg).$
The last probability goes to 0 by the convergence in probability result in Eqn. (43). Thus, we finally get for any $\varepsilon,\delta>0$, as $n\to\infty$,
$\mathbb{P}\left[\mathbb{E}\left[S^{*}(f,\widetilde{X},\hat{\theta}_{n})^{2}\bm{1}\left(S^{*}(f,\widetilde{X},\hat{\theta}_{n})\geq\varepsilon\sqrt{n}\right)\Big|\mathcal{F}_{n}\right]\geq\delta\right]\to 0,$
which verifies Eqn. (42).
With Eqns. (40) and (42) in hands, we can apply Lemma 2 to get
$\{\widetilde{S}_{n}(f)^{2}\}_{f\in\mathcal{H}}\overset{d}{\to}\{S(f)^{2}\}_{f\in\mathcal{H}}\quad\mathcal{F}-\text{stably},$
which recovers the desired Condition (C1.1).
F.7.2 Proof of Condition (C1.2)
Lemma 3 tells us that
$\sigma(\phi_{i},\phi_{i})=\mathbb{E}_{\mathbb{P}_{\theta_{0}}}[S^{*}(\phi_{i},X,\theta_{0})^{2}]\lesssim\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}[\|\phi_{i}(X)\|^{2}]+\left\|\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}\left[\left[\nabla_{\theta}s_{\theta_{0}}(X)\right]^{\top}\phi_{i}(X)\right]\right\|^{2}.$ (44)
To prove the desired condition, we sum both sides of Eqn. (44) over index $i$ to get
$\sum_{i=1}^{\infty}\sigma(\phi_{i},\phi_{i})\lesssim\underbrace{\sum_{i=1}^{\infty}\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}[\|\phi_{i}(X)\|^{2}]}_{T_{1}}+\underbrace{\sum_{i=1}^{\infty}\left\|\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}\left[\left[\nabla_{\theta}s_{\theta_{0}}(X)\right]^{\top}\phi_{i}(X)\right]\right\|^{2}}_{T_{2}}.$ (45)
We now bound terms $T_{1}$ and $T_{2}$ separately.
Bound of $T_{1}$.
Notice that $\{\phi_{i}\}_{i=1}^{\infty}$ is an orthonormal basis of $\mathcal{H}^{d}$, by the reproduction property of RKHS, we have
$\sum_{i=1}^{\infty}\|\phi_{k}(x)\|^{2}=\left\|\sum_{i=1}^{\infty}\phi_{i}(x)\right\|^{2}=\mathrm{Tr}(K(x,x)I_{d})=d\cdot K(x,x).$
Taking expectation on both sides leads to
$T_{1}=\sum_{i=1}^{\infty}\mathbb{E}_{\theta_{0}}[\|\phi_{i}(X)\|^{2}]=\mathbb{E}_{\theta_{0}}\sum_{i=1}^{\infty}[\|\phi_{i}(X)\|^{2}]\leq d\cdot\sup_{x\in\mathcal{X}}K(x,x)\leq dC_{K}<\infty,$ (46)
where Assumption 5 is applied in the last step.
Bound of $T_{2}$.
Applying Cauchy-Schwarz inequality gives
$\left\|\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}\left[\left[\nabla_{\theta}s_{\theta_{0}}(X)\right]^{\top}\phi_{i}(X)\right]\right\|^{2}$ $\leq\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}\left[\|\nabla_{\theta}s_{\theta_{0}}(X)\|_{\mathrm{F}}^{2}\right]\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}\left[\|\phi_{i}(X)\|^{2}\right]$
$\leq\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}\left[d\|\nabla_{\theta}s_{\theta_{0}}(X)\|_{\mathrm{2}}^{2}\right]\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}\left[\|\phi_{i}(X)\|^{2}\right]$
$\leq dC_{2}\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}\left[\phi_{i}(X)^{2}\right],$
where we apply Assumption 4 in the last inequality. Similar to the bound of $T_{1}$, we take summation over index $i$ to obtain
$T_{2}\leq dC_{2}\sum_{i=1}^{\infty}\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}\left[\|\phi_{i}(X)\|^{2}\right]\leq dC_{2}C_{K}<\infty.$ (47)
Combining Eqns. (46) and (47), we conclude that
$\sum_{i=1}^{\infty}\sigma(\phi_{i},\phi_{i})\lesssim d(C_{2}+1)C_{K}<\infty.$
F.7.3 Proof of Condition (C1.3)
We will mainly use a slightly modified version of [fernandez2024general, Proposition 6 ]. The version of usage is stated as below:
Lemma 11.
Suppose there exists a sequence of random measures $\nu_{n}$ on $\mathcal{X}$, such that for all $n\geq 1$, it holds that
$\mathbb{E}\left[\widetilde{S}_{n}(f)^{2}|\mathcal{F}\right]\leq C\int_{\mathcal{X}}\|f(x)\|^{2}\mathrm{d}\nu_{n}(x)\quad\text{a.s.},$ (48)
for some numerical constant $C$. Moreover suppose there exists a measure $\nu$ such that for every continuous $g\in L_{1}(\mathcal{X},\nu)$, we have
$\int_{\mathcal{X}}g(x)\mathrm{d}\nu_{n}(x)\overset{p}{\to}\int_{\mathcal{X}}g(x)\mathrm{d}\nu(x),$ (49)
Then, Condition (C1.3) holds.
The proof of Lemma 11 is provided in Appendix H.5. We now check Statements (48) and (49) hold in this setting.
Verification of Statement (48).
Choose $\nu_{n}=\mathbb{P}_{\hat{\theta}_{n}}$. The right hand side of Eqn. (48) is
$C\int_{\mathcal{X}}\|f(x)\|^{2}\mathrm{d}\nu_{n}(x)=C\,\mathbb{E}_{\hat{\theta}_{n}}\left[\|f(X)\|^{2}\right]=C\,\|f\|^{2}_{L_{2}(\mathbb{P}_{\hat{\theta}_{n}})}.$
Recall we use $\widetilde{X}$ as sample from $\mathbb{P}_{\hat{\theta}_{n}}$. We begin derivation with the left hand side. By AM-GM inequality, we obtain
$\mathbb{E}\left[\widetilde{S}_{n}(f)^{2}|\mathcal{F}\right]$ $\leq 2\underbrace{\mathbb{E}\left[\left[\mathcal{A}_{\hat{\theta}_{n}}f(\widetilde{X})\right]^{2}\Big|\mathcal{F}\right]}_{T_{3}}+2\underbrace{\mathbb{E}\left[\Big\langle\mathbb{E}_{\hat{\theta}_{n}}\left[\nabla_{\theta}s_{\hat{\theta}_{n}}(\widetilde{X})f(\widetilde{X})\right],I(\widetilde{X},\hat{\theta}_{n})\Big\rangle^{2}\Big|\mathcal{F}\right]}_{T_{4}}.$
We bound $T_{3}$ and $T_{4}$ separately.
Bound of $T_{3}$.
Applying Cauchy-Schwarz inequality implies that
$T_{3}=\mathbb{E}_{\widetilde{X}\sim\mathbb{P}_{\hat{\theta}_{n}}}[\mathcal{A}_{\hat{\theta}_{n}}f(\widetilde{X})]^{2}\leq\sup_{\theta^{\prime}\in\Theta}\|\mathcal{A}_{\theta^{\prime}}\|_{\text{op}}^{2}\mathbb{E}_{\widetilde{X}\sim\hat{\theta}_{n}}[\|f(\widetilde{X})\|^{2}]\leq C_{A}^{2}\|f\|^{2}_{L_{2}(\mathbb{P}_{\hat{\theta}_{n}})},$ (50)
where $\|\mathcal{A}_{\theta}\|_{\text{op}}$ is the operator norm of $\mathcal{A}_{\theta}$ as an linear operator in $\mathcal{H}$ equipped with $L_{2}(\mathbb{P}_{\theta})$ norm, and since $\overline{\Theta}$ is a compact set in $\mathbb{R}^{p}$ and $\mathcal{A}_{\theta}$ is continuous in $\theta$, we conclude that for some universal constant $C_{A}$, i.e., $\sup_{\theta^{\prime}\in\Theta}\|A_{\theta^{\prime}}\|_{\text{op}}\leq C_{A}$.
Write $f=(f_{1},\ldots,f_{d})^{\top}$ as a vector in $\mathcal{H}^{d}$. The existence of the $L_{2}(\mathbb{P}_{\hat{\theta}_{n}})$ norm is guaranteed noticing
$\|f(x)\|^{2}=\sum_{i=1}^{d}\langle f_{i},K_{x}\rangle^{2}\leq\sum_{i=1}^{d}\|f_{i}\|_{\mathcal{H}}^{2}\|K_{x}\|_{\mathcal{H}}^{2}=\sum_{i=1}^{d}\|f_{i}\|_{\mathcal{H}}^{2}K(x,x)\leq\sum_{i=1}^{d}\|f_{i}\|_{\mathcal{H}}^{2}C_{K}<\infty.$ (51)
where we apply Cauchy-Schwarz inequality in the first inequality and Assumption 5 in the last inequality.
Bound of $T_{4}$.
A direct calculation reveals us
$T_{4}$ $=\mathbb{E}_{\hat{\theta}_{n}}\left[\Big\langle\mathbb{E}_{\hat{\theta}_{n}}\left[\left[\nabla_{\theta}s_{\hat{\theta}_{n}}(\widetilde{X})\right]^{\top}f(\widetilde{X})\right],I(\widetilde{X},\hat{\theta}_{n})\Big\rangle^{2}\right]$
$\overset{\mathrm{(a)}}{\leq}\left\|\mathbb{E}_{\hat{\theta}_{n}}\left[\left[\nabla_{\theta}s_{\hat{\theta}_{n}}(\widetilde{X})\right]^{\top}f(\widetilde{X})\right]\right\|^{2}\ E_{\hat{\theta}_{n}}\left[\left\|I(\widetilde{X},\hat{\theta}_{n})\right\|^{2}\right]$
$\overset{\mathrm{(b)}}{\lesssim}\left\|\mathbb{E}_{\hat{\theta}_{n}}\left[\left[\nabla_{\theta}s_{\hat{\theta}_{n}}(\widetilde{X})\right]^{\top}f(\widetilde{X})\right]\right\|^{2}$
$\overset{\mathrm{(c)}}{\leq}\mathbb{E}_{\hat{\theta}_{n}}\left[\left\|\nabla_{\theta}s_{\hat{\theta}_{n}}(\widetilde{X})\right\|_{\mathrm{F}}^{2}\right]\mathbb{E}_{\hat{\theta}_{n}}\left[\|f(\widetilde{X})\|^{2}\right]\leq dC_{2}\|f\|^{2}_{L_{2}(\mathbb{P}_{\hat{\theta}_{n}})},$ (52)
where we apply Cauchy-Schwarz inequality in steps (a) and (c) and Definition 2 in step (b).
Combining the upper bounds in Eqns. (50) and (F.7.3), we have
$\mathbb{E}\left[\widetilde{S}_{n}(f)^{2}|\mathcal{F}\right]\leq 2T_{3}+2T_{4}\leq 2C_{A}^{2}\|f\|_{L_{2}(\mathbb{P}_{\hat{\theta}_{n}})}^{2}+2C_{2}\|f\|_{L_{2}(\mathbb{P}_{\hat{\theta}_{n}})}^{2}\lesssim\|f\|_{L_{2}(\mathbb{P}_{\hat{\theta}_{n}})}^{2},$ (53)
which concludes the Statement (48). ∎
Verification of Statement (49).
Choose $\nu=\mathbb{P}_{\theta_{0}}$ for $\theta_{0}$ specified in Assumption 1. By our Assumption 1, we have $\hat{\theta}_{n}\overset{p}{\to}\theta_{0}$. By Assumption 4, we have the density function $p_{\theta}(x)$ is a continuous function in both $\theta$ and $x$. By continuous mapping theorem, Statement (49) holds. ∎
Appendix G Proofs of results in Section 4
G.1 Proof of Theorem 5
Proof of Theorem 5.
We present the proof for the multiplicative local alternatives. For the additive local alternatives, the proof is similar and we briefly discuss the differences at the end of the section.
We divide the proof into three steps. Recall the definition of $\mu_{i}$ in Theorem 5:
$\mu_{i}$ $\equiv\mathbb{E}_{\theta_{0}}[S^{*}(\phi_{i},X,\theta_{0})h(X)],$ (54)
where $\phi_{i}$ is the $i$-th orthonormal basis of $\mathcal{H}^{d}$. By the expression for the multiplicative local alternatives, we can write the full density function of $\mathcal{L}_{n}$ as
$p_{1,n}(x)=\frac{1}{Z_{n}}p_{\theta_{0}}(x)\left(1+\frac{h(x)}{\sqrt{n}}\right),$
where $Z_{n}\equiv 1+\mathbb{E}_{\theta_{0}}[h(X)]/\sqrt{n}<\infty$. We stick to this notation of density in the proof.
Step 1: Estimation under local alternative.
The first step is to characterize the asymptotic performance of the estimation procedure under the local alternative. It turns out that the estimator shares the asymptotic linear expansion as in the null hypothesis with a non-centered influence function. The result is stated in the following lemma with proof in Appendix H.6.
Lemma 12.
The estimator $\hat{\theta}_{n}$ respects the following asymptotically linear expansion:
$\sqrt{n}(\hat{\theta}_{n}-\theta_{0})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}I(X_{i},\theta_{0})+o_{p}\left(1\right),$ (55)
and it weakly converges to a normal distribution with a non-zero mean $\tau=(\tau_{1},\ldots,\tau_{k})$ and covariance matrix $\Sigma$.
$\sqrt{n}(\hat{\theta}_{n}-\theta_{0})\overset{d}{\rightarrow}\mathcal{N}(\tau,\Sigma),$ (56)
where $\tau_{j}=\mathrm{Cov}_{\theta_{0}}[[I(X,\theta_{0})]_{j},h(X)]$ for $j\in[k]$ and $\Sigma=\mathrm{Cov}_{\theta_{0}}[I(X,\theta_{0})]$.
Lemma 12 is used to characterize the influence of the estimating procedure on the test statistic under the local alternative.
Step 2: Asymptotic expansion.
The second step of the proof is to derive the asymptotic expansion of the test statistic under the local alternative.
Lemma 13.
Define functional $S_{1,n}:\mathcal{H}^{d}\to\mathbb{R}$ as
$S_{1,n}(f)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathcal{A}_{\theta_{0}}f(X_{i})+\Big\langle\mathbb{E}_{X\sim\mathcal{L}_{n}}\left[\left[\nabla_{\theta}s_{\theta_{0}}(X)\right]^{\top}f(X)\right],I(X_{i},\theta_{0})\Big\rangle.$
The asymptotic variational form of the test statistic under local alternatives can be written as
$nT_{n}=\sup_{f\in\mathcal{B}_{\mathcal{H}^{d}}(1)}S_{1,n}(f)^{2}+o_{p}(1).$
The proof of Lemma 13 is similar to the proof of Theorem 3, so we omit the proof here. The variational form of the test statistic provides a convenient way to analyze the asymptotic behavior of the test statistic in the functional space, which we focus on in the remaining proof.
Step 3: Weak convergence in the functional space.
To prove the weak convergence of the test statistic, we modify the conditions in Lemma 5 to fit the local alternative setting.
Lemma 14.
With the same setup in Lemma 5, we introduce the new conditions:
  1. (1)”
    There exists a mean functional $\mu(f)$ and a covariance functional $\sigma(f_{1},f_{2})$ such that for any $f\in\mathcal{H}^{d}$, we have
    $S_{1,n}(f)\overset{d}{\to}S_{1}(f)\sim\mathcal{N}(\mu(f),\sigma(f,f)),$ (C3.1)
  2. (3)”
    For some orthonormal basis $(\phi_{i})_{i=1}^{\infty}$ of RKHS $\mathcal{H}$ and any $\varepsilon>0$, we have
    $\lim_{k\to\infty}\limsup_{n\to\infty}\mathbb{P}_{\mathcal{L}_{n}}\left(\sum_{i=k}^{\infty}S_{1,n}(\phi_{i})^{2}\geq\varepsilon\right)=0,$ (C3.3)
Suppose the Conditions (C3.1), (C1.2) and (C3.3) hold, then under measure $\mathcal{L}_{n}$, we have
$\{S_{1,n}(f)\}_{f\in\mathcal{H}^{d}}\overset{d}{\to}\{S_{1}(f)\}_{f\in\mathcal{H}^{d}}.$
The proof of Lemma 14 is similar to the proof of Lemma 5, so we omit the proof here. To apply Lemma 14, we need to verify the conditions in the lemma. Notice that (C1.2) is not a probabilistic statement, and it is already verified in the proof of Theorem 3, and we can directly use it here. The verification of conditions (C3.1) and (C3.3) can be found in section G.2. After verifying Conditions (C3.1) and (C3.3), we conclude that
$\{S_{1,n}(f)\}_{f\in\mathcal{H}^{d}}\overset{d}{\to}\{S_{1}(f)\}_{f\in\mathcal{H}^{d}}.$
thus by continuous mapping theorem, we have under $\mathcal{L}_{n}$,
$nT_{n}\overset{d}{\to}\|S_{1}(f)\|_{\mathcal{H}^{d}\to\mathbb{R}}\equiv\sup_{\|f\|_{\mathcal{H}^{d}}\leq 1}S_{1}(f).$
To get the explicit expression for $\|S_{1}(f)\|_{\mathcal{H}^{d}\to\mathbb{R}}$, we notice that
$\sup_{\|f\|_{\mathcal{H}^{d}}\leq 1}S_{1}(f)^{2}$ $=\left(\sup_{\|f\|_{\mathcal{H}^{d}}\leq 1}\Big\langle\sum_{i=1}^{\infty}\sqrt{\lambda_{i}}(Z_{i}+\mu_{i})\phi_{i},f\Big\rangle\right)^{2}$
$=\left\|\sum_{i=1}^{\infty}\sqrt{\lambda_{i}}(Z_{i}+\mu_{i})\phi_{i}\right\|^{2}_{\mathcal{H}^{d}}$
$=\sum_{i=1}^{\infty}\lambda_{i}(Z_{i}+\mu_{i})^{2}=W_{1}$
which is a non-central chi-squared distribution with infinite degrees of freedom. ∎
G.2 Verification of conditions in Lemma 14
G.2.1 Proof of Condition (C3.1)
In our setting, the limit functional $S_{1}(f)$ can be explicitly constructed as
$S_{1}(f)=\sum_{i=1}^{\infty}\langle\phi_{i},f\rangle_{\mathcal{H}^{d}}\sqrt{\lambda_{i}}(Z_{i}+\mu_{i})\quad\text{and}\quad\mu(f)=\sum_{i=1}^{\infty}\langle\phi_{i},f\rangle_{\mathcal{H}^{d}}\,\mu_{i},$
where $\mu_{i}=\mathbb{E}_{\theta_{0}}[S^{*}(\phi_{i},X,\theta_{0})h(X)]$ and other notations are defined the same as in Condition (C1.1). Recall in Eqn. (27), we define the functional $S_{n}(f)$. We prove this result by showing the weak convergence of the functional $S_{n}(f)$ under $\mathcal{L}_{n}$, then consider the difference between $S_{n}(f)$ and $S_{1,n}(f)$.
Weak convergence of $S_{n}(f)$ under $\mathcal{L}_{n}$.
With the same proof as in Lemma 10, we show that under $\mathbb{P}_{\theta_{0}}$ the functional $S_{n}(f)$ satisfies
$S_{n}(f)\overset{d}{\to}S(f)\sim\mathcal{N}(0,\sigma(f,f)).$
Furthermore, recall that in Eqn. (86), we prove that
$\log\left(\frac{\mathrm{d}\mathcal{L}_{n}}{\mathrm{d}P_{\theta_{0}}}(X)\right)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(h(X_{i})-\mathbb{E}_{\theta_{0}}h(X)\right)-\frac{1}{2}\mathrm{Var}_{\theta_{0}}[h(X)]+o_{p}(1).$
Combining the expression of $S_{n}(f)$ and the above expansion, we have under measure $\mathbb{P}_{\theta_{0}}$,
$\begin{pmatrix}S_{n}(f)\\ \log\left(\frac{\mathrm{d}\mathcal{L}_{n}}{\mathrm{d}P_{\theta_{0}}}(X)\right)\end{pmatrix}$ $\overset{d}{\rightarrow}\mathcal{N}\left(\begin{pmatrix}0\\ -\frac{1}{2}M_{2}^{h}\end{pmatrix},\begin{pmatrix}\sigma(f,f)&\mu(f)\\ \mu(f)&M_{2}^{h}\end{pmatrix}\right)$.
By Le Cam’s third lemma [van2000asymptotic, Theorem 6.6], we conclude that under local alternative $\mathcal{L}_{n}$,
$S_{n}(f)\overset{d}{\to}S_{1}(f)\sim\mathcal{N}(\mu(f),\sigma(f,f)).$ (57)
Difference between $S_{n}(f)$ and $S_{1,n}(f)$.
Comparing the expression of $S_{n}(f)$ and $S_{1,n}(f)$, we have
$S_{1,n}(f)-S_{n}(f)$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\Big\langle\mathbb{E}_{X\sim\mathcal{L}_{n}}\left[\left[\nabla_{\theta}s_{\theta_{0}}(X)\right]^{\top}f(X)\right]-\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}\left[\left[\nabla_{\theta}s_{\theta_{0}}(X)\right]^{\top}f(X)\right],I(X_{i},\theta_{0})\Big\rangle$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\Big\langle\int_{\mathcal{X}}\nabla_{\theta}s_{\theta_{0}}(x)f(x)p_{\theta_{0}}(x)\left(\frac{1+\frac{1}{\sqrt{n}}h(x)}{Z_{n}}-1\right)\lambda(\mathrm{d}x),I(X_{i},\theta_{0})\Big\rangle$
$\leq\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\sum_{j=1}^{k}\left(\mathbb{E}_{\theta_{0}}\left[\left[\left[\nabla_{\theta}s_{\theta_{0}}(X)\right]^{\top}f(X)\right]_{j}^{2}\right]\right)^{1/2}\cdot\left[I(X_{i},\theta_{0})\right]_{j}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\cdot\mathbb{E}_{\theta_{0}}\left[\left(1+\frac{1}{\sqrt{n}}h(x)-Z_{n}\right)^{2}\right]^{1/2},$
where we use Cauchy-Schwarz inequality in the inequality above.
To get the final bound the above expression, we analysis terms separately in the above inequality. By Cauchy-Schwarz inequality, we reach
$\sum_{j=1}^{k}\left(\mathbb{E}_{\theta_{0}}\left[\left[\left[\nabla_{\theta}s_{\theta_{0}}(X)\right]^{\top}f(X)\right]_{j}^{2}\right]\right)^{1/2}\cdot\left[I(X_{i},\theta_{0})\right]_{j}$
$\leq\sum_{j=1}^{k}\mathbb{E}_{\theta_{0}}\left[\left[\left[\nabla_{\theta}s_{\theta_{0}}(X)\right]^{\top}f(X)\right]_{j}^{2}\right]\cdot\sum_{j=1}^{k}\left[I(X_{i},\theta_{0})\right]_{j}^{2}$
$\leq\sum_{j=1}^{k}\mathbb{E}_{\theta_{0}}\left[\left[\left[\nabla_{\theta}s_{\theta_{0}}(X)\right]^{\top}f(X)\right]_{j}^{2}\right]\cdot\|I(X_{i},\theta_{0})\|^{2}.$ (58)
By the boundness of $f\in\mathcal{H}^{d}$ (See Eqn. (51)), we have, for any $j\in[k]$,
$\mathbb{E}_{\theta_{0}}\left[\left[\left[\nabla_{\theta}s_{\theta_{0}}(X)\right]^{\top}f(X)\right]_{j}^{2}\right]$ $\leq\mathbb{E}_{\theta_{0}}\left[\left\|\nabla_{\theta}s_{\theta_{0}}(X)_{j,\cdot}\right\|^{2}\right]\sup_{x\in\mathcal{X}}\|f(x)\|^{2}$
$\leq\mathbb{E}_{\theta_{0}}\left[\left\|\nabla_{\theta}s_{\theta_{0}}(X)_{j,\cdot}\right\|^{2}\right]C_{K}^{2}\|f\|^{2}_{\mathcal{H}^{d}}.$ (59)
Plugging Eqn. (59) into Eqn. (58) gives
$\sum_{j=1}^{k}\left(\mathbb{E}_{\theta_{0}}\left[\left[\left[\nabla_{\theta}s_{\theta_{0}}(X)\right]^{\top}f(X)\right]_{j}^{2}\right]\right)^{1/2}\cdot\left[I(X_{i},\theta_{0})\right]_{j}$
$\leq\mathbb{E}_{\theta_{0}}\left[\|\nabla_{\theta}s_{\theta_{0}}(X)\|^{2}_{\mathrm{F}}\right]C_{K}^{2}\|f\|_{\mathcal{H}^{d}}\cdot\|I(X_{i},\theta_{0})\|^{2}$
$\leq dC_{2}C_{K}^{2}\|f\|_{\mathcal{H}^{d}}^{2}\cdot\|I(X_{i},\theta_{0})\|^{2},$
where we apply Assumption 4 in the last step. Recall Assumption 1 implies a uniform bound on $\mathbb{E}[\|I(X_{i},\theta_{0})\|^{2}]$, we have
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\sum_{j=1}^{k}\left(\mathbb{E}_{\theta_{0}}\left[\left[\left[\nabla_{\theta}s_{\theta_{0}}(X)\right]^{\top}f(X)\right]_{j}^{2}\right]\right)^{1/2}\cdot\left[I(X_{i},\theta_{0})\right]_{j}=O_{p}(1)$ (60)
Also, by Assumption 2, we have
$\mathbb{E}_{\theta_{0}}\left[\left(1+\frac{1}{\sqrt{n}}h(x)-Z_{n}\right)^{2}\right]=O_{p}(n^{-1/2}).$ (61)
Combining Eqns. (60) and (61) with Assumption 1, we have, for any fixed $f\in\mathcal{H}^{d}$,
$S_{1,n}(f)-S_{n}(f)=O_{p}(1)\cdot O_{p}(n^{-1/2})=O_{p}(n^{-1/2}).$
Plugging it into Eqn. (57) implies that under measure $\mathcal{L}_{n}$,
$S_{1,n}(f)\overset{d}{\to}S_{1}(f)\sim\mathcal{N}(\mu(f),\sigma(f,f)),$
which verifies Condition (C3.1).
G.2.2 Proof of Condition (C3.3)
By Markov ineuqlaity, we have
$\limsup_{n\to\infty}\mathbb{P}_{\mathcal{L}_{n}}\left(\sum_{i=k}^{\infty}S_{1,n}(\phi_{i})^{2}\geq\varepsilon\right)$ $\leq\limsup_{n\to\infty}\frac{1}{\varepsilon}\sum_{i=k}^{\infty}\mathbb{E}_{\mathcal{L}_{n}}[S_{1,n}(\phi_{i})^{2}]$
$=\frac{1}{\varepsilon}\sum_{i=k}^{\infty}\left(\mu_{i}^{2}+\sigma(\phi_{i},\phi_{i})\right).$
It is proven in Condition (C1.2) that $\sum_{i=1}^{\infty}\sigma(\phi_{i},\phi_{i})<\infty$. Noticing that
$\sum_{i=1}^{\infty}\mu_{i}^{2}\leq\sum_{i=1}^{\infty}\mathbb{E}_{\theta_{0}}\left[S^{*}(\phi_{i},X,\theta_{0})^{2}\right]\mathbb{E}_{\theta_{0}}\left[h(X)^{2}\right]\leq\sum_{i=1}^{\infty}\sigma(\phi_{i},\phi_{i})\mathbb{E}_{\theta_{0}}\left[h(X)^{2}\right]<\infty,$
we can conclude that
$\lim_{k\to\infty}\limsup_{n\to\infty}\mathbb{P}_{\mathcal{L}_{n}}\left(\sum_{i=k}^{\infty}S_{1,n}(\phi_{i})^{2}\geq\varepsilon\right)\leq\lim_{k\to\infty}\frac{1}{\varepsilon}\sum_{i=k}^{\infty}\left(\mu_{i}^{2}+\sigma(\phi_{i},\phi_{i})\right)=0,$
which proves Condition (C3.3).
G.3 Proof of the additive case in Theorem 5
Proof of the additive case in Theorem 5.
We now prove Theorem 5 for the additive case with
$\mu_{i}$ $=\gamma\mathbb{E}_{\mathcal{L}_{g}}[S^{*}(\phi_{i},X,\theta_{0})],$
where $\phi_{i}$ is the $i$-th orthonormal basis of $\mathcal{H}$. The density function of $\mathcal{L}_{n}$ in this case can be written as
$p_{1,n}(x)$ $=\left(1-\frac{\gamma}{\sqrt{n}}\right)p_{\theta_{0}}(x)+\frac{\gamma}{\sqrt{n}}g(x)$
$=p_{\theta_{0}}(x)\left(1+\frac{\gamma}{\sqrt{n}}\left(\frac{g(x)}{p_{\theta_{0}}(x)}-1\right)\right),$
which can be fitted into the multiplicative local alternatives with $h(x)=\gamma(g(x)/p_{\theta_{0}}(x)-1)$. It can be directly checked that Assumption 2 hold. Thus, by Theorem 5, we immediately have Theorem 5 holds with
$\mu_{i}$ $=\mathbb{E}_{\theta_{0}}\left[S^{*}(\phi_{i},X,\theta_{0})\cdot\gamma\left(\frac{g(x)}{p_{\theta_{0}}(x)}-1\right)\right]$
$=\gamma\mathbb{E}_{\mathcal{L}_{g}}[S^{*}(\phi_{i},X,\theta_{0})]-\gamma\mathbb{E}_{\theta_{0}}[S^{*}(\phi_{i},X,\theta_{0})]$
$=\gamma\mathbb{E}_{\mathcal{L}_{g}}[S^{*}(\phi_{i},X,\theta_{0})],$
where we apply Lemma 3 in the last line. ∎
Appendix H Proofs of main lemmas
H.1 Proof of Lemma 3
Proof of Lemma 3(1).
Notice that $X\sim\mathbb{P}_{\theta}$, the property of kernelized Stein class in Eqn. (6), we have
$\mathbb{E}_{X\sim\mathbb{P}_{\theta}}\left[\mathcal{A}_{\theta}f(X)\right]=0.$ (62)
Additionally, Assumption 1 ensures that $\mathbb{E}_{X\sim\mathbb{P}_{\theta}}[I(X,\theta)]=0$, which leads to
$\mathbb{E}_{X\sim\mathbb{P}_{\theta}}\left[\Big\langle\mathbb{E}_{X^{\prime}\sim P_{\theta}}\left[[\nabla_{\theta}s_{\theta}(X^{\prime})]^{\top}f(X^{\prime})\right],I(X,\theta)\Big\rangle\right]$
$=\Big\langle\mathbb{E}_{X^{\prime}\sim P_{\theta}}\left[[\nabla_{\theta}s_{\theta}(X^{\prime})]^{\top}f(X^{\prime})\right],\mathbb{E}_{X\sim\mathbb{P}_{\theta}}\left[I(X,\theta)\right]\Big\rangle=0.$ (63)
Combining Eqn. (62) and (63) gives $\mathbb{E}_{X\sim\mathbb{P}_{\theta}}[S^{*}(f,X,\theta)]=0$. ∎
Proof of Lemma 3(2).
The second moment of $S^{*}(f,X,\theta)$ can be bounded as
$\mathbb{E}_{X\sim\mathbb{P}_{\theta}}[S^{*}(f,X,\theta)^{2}]$
$\leq 2\mathbb{E}_{X\sim\mathbb{P}_{\theta}}[|\mathcal{A}_{\theta}f(X)|^{2}]+2\mathbb{E}_{X\sim\mathbb{P}_{\theta}}\left[\Big\langle\mathbb{E}_{X^{\prime}\sim P_{\theta}}\left[[\nabla_{\theta}s_{\theta}(X^{\prime})]^{\top}f(X^{\prime})\right],I(X,\theta)\Big\rangle^{2}\right]$
$\leq 2\mathbb{E}_{X\sim\mathbb{P}_{\theta}}[|\mathcal{A}_{\theta}f(X)|^{2}]+2\mathbb{E}_{X^{\prime}\sim\mathbb{P}_{\theta}}\left[\left\|[\nabla_{\theta}s_{\theta}(X^{\prime})]^{\top}f(X^{\prime})\right\|^{2}\right]\mathbb{E}_{X\sim\mathbb{P}_{\theta}}[\|I(X,\theta)\|^{2}],$ (64)
where AM-GM inequality is applied in the first line and Cauchy-Schwarz inequality is applied in the second line. With the same argument as in Eqn. (50), we obtain for some constant $C_{A}$ independent of $f$ and $\theta$,
$\mathbb{E}_{X\sim\mathbb{P}_{\theta}}[|\mathcal{A}_{\theta}f(X)|^{2}]$ $\leq C_{A}\|f\|^{2}_{L^{2}(\mathbb{P}_{\theta})}.$ (65)
Moreover, by Assumption 1, we have, for some constant $C_{I}$ independent of $\theta$,
$\mathbb{E}_{X\sim\mathbb{P}_{\theta}}[\|I(X,\theta)\|^{2}]\leq C_{I}.$ (66)
Combining Eqns. (64), (65) and (66) gives
$\mathbb{E}_{X\sim\mathbb{P}_{\theta}}[S^{*}(f,X,\theta)^{2}]$ $\lesssim\mathbb{E}_{X\sim\mathbb{P}_{\theta}}\left[\|f(X)\|^{2}\right]+\mathbb{E}_{X^{\prime}\sim\mathbb{P}_{\theta}}\left[\left\|[\nabla_{\theta}s_{\theta}(X^{\prime})]^{\top}f(X^{\prime})\right\|^{2}\right]$
which proves the bound Lemma 3(2). Also, the right-hand side of the above inequality can be uniformly bounded in $\theta$ by the boundness of $f$ in $\mathcal{H}^{d}$ (i.e., definition of RKHS) and Assumption 1. ∎
H.2 Proof of Lemma 4
Proof of Lemma 4.
We split the derivation the variational form of the bootstrap statistic into two steps.
Step 1: Derive the intermediary form.
Recall the KSD kernel $h_{p}(\cdot,\cdot)$ defined in Eqn. (7). We denote the kernel under measure $\mathbb{P}_{\theta}$ as $h_{\theta}(\cdot,\cdot)\coloneq h_{\mathbb{P}_{\theta}}(\cdot,\cdot)$. Define $F_{\theta}(x,x^{\prime})=\nabla_{\theta}h_{\theta}(x,x^{\prime})$ and $G(x,\theta)=\mathbb{E}_{X\sim\mathbb{P}_{\theta}}F_{\theta}(x,X)$. Let $\widetilde{X}=\{\widetilde{X}_{1},\ldots,\widetilde{X}_{n}\}$ be the bootstrap sample.
The bootstrap statistic admits the following intermediary decomposition:
$n\widetilde{T}_{n}$ $=nT_{n}(\widetilde{X},\hat{\theta}_{n})+\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}G(\widetilde{X}_{j},\hat{\theta}_{n})^{\top}I(\widetilde{X}_{i},\hat{\theta}_{n})$
$\qquad\qquad\qquad+\frac{1}{2n}\sum_{i=1}^{n}\sum_{j=1}^{n}I(\widetilde{X}_{i},\hat{\theta}_{n})^{\top}H_{n}I(\widetilde{X}_{i},\hat{\theta}_{n})+o_{p}(1),$ (67)
where $H_{n}=\nabla_{\theta}^{2}T_{n}(\widetilde{X},\hat{\theta}_{n})$.
We prove Statement (67) in the following part of this step. We first use Taylor expansion around $\hat{\theta}_{n}$, which leads to
$n\widetilde{T}_{n}$ $=nT_{n}(\widetilde{X},\hat{\theta}_{n})+\big(n\widetilde{T}_{n}-nT_{n}(\widetilde{X},\hat{\theta}_{n})\big)$
$=nT_{n}(\widetilde{X},\hat{\theta}_{n})+n\nabla_{\theta}T_{n}(\widetilde{X},\hat{\theta}_{n})^{\top}(\widetilde{\theta}_{n}-\hat{\theta}_{n})$
$\qquad\qquad+\frac{n}{2}(\widetilde{\theta}_{n}-\hat{\theta}_{n})^{\top}\nabla_{\theta}^{2}T_{n}(\widetilde{X},\hat{\theta}_{\gamma})(\widetilde{\theta}_{n}-\hat{\theta}_{n})+o_{p}(n\|\widetilde{\theta}_{n}-\hat{\theta}_{n}\|^{2})$
$=nT_{n}(\widetilde{X},\hat{\theta}_{n})+n\nabla_{\theta}T_{n}(\widetilde{X},\hat{\theta}_{n})^{\top}(\widetilde{\theta}_{n}-\hat{\theta}_{n})$
$\qquad\qquad\qquad\qquad+\frac{n}{2}(\widetilde{\theta}_{n}-\hat{\theta}_{n})^{\top}H_{n}(\widetilde{\theta}_{n}-\hat{\theta}_{n})+o_{p}(n\|\widetilde{\theta}_{n}-\hat{\theta}_{n}\|^{2}).$ (68)
where the last inequality uses Lemma 15. To further address the last term, we first use Assumption 1 to get that $\hat{\theta}_{n}$ is a asymptotically linear estimator of $\theta_{0}$. Thus, we have $\hat{\theta}_{n}\overset{p}{\rightarrow}\theta_{0}$, which implies
$\lim_{n\to\infty}\mathbb{P}\left(\widehat{\theta}_{n}\notin B_{\theta_{0}}(\delta)\right)=0.$ (69)
In the next step, we apply the definition of uniform asymptotically linear estimator (Def. 2) on the estimator $\hat{\theta}_{n}$. Working on the event $\left\{\widehat{\theta}_{n}\in B_{\theta_{0}}(\delta)\right\}$, we recall Assumption 1 to get For any $\varepsilon>0$, we have
$\mathbb{P}\left(\left|\sqrt{n}(\widetilde{\theta}_{n}-\hat{\theta}_{n})-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}I(\widetilde{X}_{i},\hat{\theta}_{n})\right|\geq\varepsilon\right)$
$\leq\mathbb{P}\left(\left|\sqrt{n}(\widetilde{\theta}_{n}-\hat{\theta}_{n})-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}I(\widetilde{X}_{i},\hat{\theta}_{n})\right|\geq\varepsilon\bigg|\left\{\widehat{\theta}_{n}\in B_{\theta_{0}}(\delta)\right\}\right)+\mathbb{P}\left(\widehat{\theta}_{n}\notin B_{\theta_{0}}(\delta)\right).$
Then first term is a vanishing term by the definition of uniform asymptotically linear estimator. Combining this result with equation (69), we have
$\lim_{n\to\infty}\mathbb{P}\left(\left|\sqrt{n}(\widetilde{\theta}_{n}-\hat{\theta}_{n})-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}I(\widetilde{X}_{i},\hat{\theta}_{n})\right|\geq\varepsilon\right)=0.$
Writing into a compact form:
$\sqrt{n}(\widetilde{\theta}_{n}-\hat{\theta}_{n})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}I(\widetilde{X}_{i},\hat{\theta}_{n})+o_{p}(1),$ (70)
which in turn implies that
$n\|\widetilde{\theta}_{n}-\hat{\theta}_{n}\|^{2}$ $=O_{p}\left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}I(\widetilde{X}_{i},\hat{\theta}_{n})\right)=O_{p}(1).$ (71)
Plugging Eqns. (70) and (71) into the previous Eqn. (68), we have
$n\widetilde{T}_{n}$ $=nT_{n}(\widetilde{X},\hat{\theta}_{n})+\big[\sqrt{n}\nabla_{\theta}T_{n}(\widetilde{X},\hat{\theta}_{n})^{\top}\big]\left[\sqrt{n}(\widetilde{\theta}_{n}-\hat{\theta}_{n})\right]$
$\qquad\qquad\qquad\qquad+\frac{1}{2}\left[\sqrt{n}(\widetilde{\theta}_{n}-\hat{\theta}_{n})^{\top}\right]H_{n}\left[\sqrt{n}(\widetilde{\theta}_{n}-\hat{\theta}_{n})\right]+o_{p}(1)$
$=nT_{n}(\widetilde{X},\hat{\theta}_{n})+\bigg[\frac{1}{\sqrt{n}}\sum_{i=1}^{n}G(\widetilde{X}_{i},\hat{\theta}_{n})\bigg]\left[\frac{1}{\sqrt{n}}\sum_{i=1}^{n}I(\widetilde{X}_{i},\hat{\theta}_{n})\right]$
$\qquad\qquad\qquad\qquad+\frac{1}{2}\bigg[\frac{1}{\sqrt{n}}\sum_{i=1}^{n}I(\widetilde{X}_{i},\hat{\theta}_{n})\bigg]H_{n}\bigg[\frac{1}{\sqrt{n}}\sum_{i=1}^{n}I(\widetilde{X}_{i},\hat{\theta}_{n})\bigg]+o_{p}(1)$
$=nT_{n}(\widetilde{X},\hat{\theta}_{n})+\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}G(\widetilde{X}_{j},\hat{\theta}_{n})^{\top}I(\widetilde{X}_{i},\hat{\theta}_{n})$
$\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{2n}\sum_{i=1}^{n}\sum_{j=1}^{n}I(\widetilde{X}_{i},\hat{\theta}_{n})^{\top}H_{n}I(\widetilde{X}_{i},\hat{\theta}_{n})+o_{p}(1),x$ (72)
where in the second equality we use Lemma 16.
Step 2: Equivalence of intermediary form and variational form.
To prove the equivalence between the intermediary form and the variational form, it suffices to show that
$\sup_{f\in\mathcal{B}_{\mathcal{H}^{d}}(1)}\widetilde{S}_{n}(f)^{2}$ $=nT_{n}(\widetilde{X},\hat{\theta}_{n})+\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}G(\widetilde{X}_{j},\hat{\theta}_{n})^{\top}I(\widetilde{X}_{i},\hat{\theta}_{n})$
$\qquad\qquad\qquad+\frac{1}{2n}\sum_{i=1}^{n}\sum_{j=1}^{n}I(\widetilde{X}_{i},\hat{\theta}_{n})^{\top}H_{n}I(\widetilde{X}_{i},\hat{\theta}_{n})+o_{p}(1).$ (73)
We prove Statement (73) in the following part of this step. To simplify the notation in the proof, we define auxiliary functionals
$\widetilde{A}_{n}(f)$ $\equiv\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathcal{A}_{\hat{\theta}_{n}}f(\widetilde{X}_{i})$
$\widetilde{B}_{n}(f)$ $\equiv\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\Big\langle\mathbb{E}_{X\sim P_{\hat{\theta}_{n}}}\left[[\nabla_{\theta}s_{\hat{\theta}_{n}}(X)]^{\top}f(X)\right],I(\widetilde{X}_{i},\hat{\theta}_{n})\Big\rangle$
Since we work on bivariate kernel function $K(\cdot,\cdot)$, we use the superscript $1,2$ to denote the order of arguments the functional applied on. For example, for functional $\widetilde{A}_{n}:\mathcal{H}^{d}\to\mathbb{R}$, notation $\widetilde{A}_{n}^{1}K(x,y)$ means taking $y$ as a constant and applying $\widetilde{A}_{n}$ onto $K(\cdot,y)$.
We begin by explicitly calculate the left hand side of Equation (73). It is straightforward to check that $\widetilde{S}_{n}(f)$ is a linear functional on $\mathcal{H}^{d}$ By Riesz representation theorem [yosida2012functional, Page 81 ], there exists a unique $f_{n}^{\star}\in\mathcal{H}^{d}$ such that $\widetilde{S}_{n}(f)=\langle f_{n}^{\star},f\rangle_{\mathcal{H}^{d}}$. Then we obtain
$\sup_{f\in\mathcal{B}_{\mathcal{H}^{d}}(1)}\widetilde{S}_{n}(f)^{2}=\sup_{\|f\|_{\mathcal{H}^{d}}\leq 1}\langle f_{n}^{\star},f\rangle_{\mathcal{H}^{d}}^{2}=\|f_{n}^{\star}\|_{\mathcal{H}^{d}}^{2},$
where we apply Cauchy-Schwarz inequality. To get the expression of $f_{n}^{\star}$, we observe
$\widetilde{S}_{n}(f_{n}^{\star})=\langle f_{n}^{\star},f_{n}^{\star}\rangle_{\mathcal{H}^{d}}=\|f_{n}^{\star}\|_{\mathcal{H}^{d}}^{2},$ (74)
and by reproducing property, we have for any $j\in[d]$,
$[f_{n}^{\star}(x)]_{j}=\langle[f_{n}^{\star}]_{j},K(\cdot,x)\rangle_{\mathcal{H}}=\langle[f_{n}^{\star}]_{j},K_{x}\rangle_{\mathcal{H}}.$ (75)
Directly from equation (74) and (75), we notice
$S_{n}(K_{x})=\langle f_{n}^{\star},K_{x}\bm{1}_{d}\rangle_{\mathcal{H}^{d}}=\sum_{j=1}^{d}[f_{n}^{\star}(x)]_{j},$
where $\bm{1}_{d}$ is the vector in $\mathbb{R}^{d}$ with all entries one. Finally, we achieve
$\|f_{n}^{\star}\|_{\mathcal{H}^{d}}^{2}=S_{n}(f_{n}^{\star})=S_{n}^{1}S_{n}^{2}[\mathbf{K}(x,x^{\prime})],$
where we define $\mathbf{K}(x,x^{\prime})\equiv K(x,x^{\prime})I_{d}$, and the superscript indicates the argument operator is applied on. We follow this notation below.
$\sup_{f\in\mathcal{B}_{\mathcal{H}^{d}}(1)}\widetilde{S}_{n}(f)^{2}$ $=(\widetilde{A}_{n}^{1}+\widetilde{B}_{n}^{1})(\widetilde{A}_{n}^{2}+\widetilde{B}_{n}^{2})[\mathbf{K}(x,x^{\prime})]$
$=\underbrace{\widetilde{A}_{n}^{1}\widetilde{A}_{n}^{2}[\mathbf{K}(x,x^{\prime})]}_{\widetilde{T}_{1}}+2\underbrace{\widetilde{A}_{n}^{1}\widetilde{B}_{n}^{2}[\mathbf{K}(x,x^{\prime})]}_{\widetilde{T}_{2}}+\underbrace{\widetilde{B}_{n}^{1}\widetilde{B}_{n}^{2}[\mathbf{K}(x,x^{\prime})]}_{\widetilde{T}_{3}}.$
We then turn to derive the explicit formula for $\widetilde{T}_{1},\widetilde{T}_{2}$ and $\widetilde{T}_{3}$ separately.
Calculation of $\widetilde{T}_{1}$.
We notice that
$\widetilde{T}_{1}$ $=\widetilde{A}_{n}^{1}\widetilde{A}_{n}^{2}[\mathbf{K}(x,x^{\prime})]$
$=\widetilde{A}_{n}^{1}\left[\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\mathcal{A}_{\hat{\theta}_{n}}^{2}\mathbf{K}(x,\widetilde{X}_{j})\right]$
$=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}\mathcal{A}_{\hat{\theta}_{n}}^{1}\mathcal{A}_{\hat{\theta}_{n}}^{2}\mathbf{K}(\widetilde{X}_{i},\widetilde{X}_{j})=nT_{n}(\widetilde{X},\hat{\theta}_{n}).$ (76)
Calculation of $\widetilde{T}_{2}$.
We again have
$\widetilde{T}_{2}$ $=\widetilde{A}_{n}^{1}\widetilde{B}_{n}^{2}[\mathbf{K}(x,x^{\prime})]$
$=\widetilde{A}_{n}^{1}\bigg[\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\Big\langle\mathbb{E}_{\widetilde{X}\sim\mathbb{P}_{\hat{\theta}_{n}}}[\nabla_{\theta}\mathcal{A}_{\theta}^{2}\mathbf{K}(x,\widetilde{X})],I(\widetilde{X}_{j},\hat{\theta}_{n})\Big\rangle\bigg]\Big|_{\theta=\hat{\theta}_{n}}$
$=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}\Big\langle\mathbb{E}_{\widetilde{X}\sim\mathbb{P}_{\hat{\theta}_{n}}}[\nabla_{\theta}\mathcal{A}_{\hat{\theta}_{n}}^{1}\mathcal{A}_{\theta}^{2}\mathbf{K}(\widetilde{X}_{i},\widetilde{X})],I(\widetilde{X}_{j},\hat{\theta}_{n})\Big\rangle\Big|_{\theta=\hat{\theta}_{n}}$
$=\frac{1}{2n}\sum_{i=1}^{n}\sum_{j=1}^{n}\Big\langle G(\widetilde{X}_{i},\hat{\theta}_{n}),I(\widetilde{X}_{j},\hat{\theta}_{n})\Big\rangle$ (77)
$=\frac{1}{2n}\sum_{i=1}^{n}\sum_{j=1}^{n}G(\widetilde{X}_{i},\hat{\theta}_{n})^{\top}I(\widetilde{X}_{j},\hat{\theta}_{n}),$ (78)
where Eqn. (77) is due to direct calculation on the derivative.
Calculation of $\widetilde{T}_{3}$.
we can similarly compute that
$\widetilde{T}_{3}$ $=\widetilde{B}_{n}^{1}\widetilde{B}_{n}^{2}[K(x,x^{\prime})]$
$=\widetilde{B}_{n}^{1}\bigg[\frac{1}{\sqrt{n}}\sum_{j=1}^{n}\Big\langle\mathbb{E}_{\widetilde{X}\sim\mathbb{P}_{\hat{\theta}_{n}}}\left[\nabla_{\theta}\mathcal{A}_{\theta_{1}}^{2}\mathbf{K}(x,X)\right],I(\widetilde{X}_{j},\hat{\theta}_{n})\Big\rangle\bigg]\bigg|_{\theta_{1}=\hat{\theta}_{n}}$
$=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}\bigg\langle\mathbb{E}_{\widetilde{X}^{\prime}\sim\mathbb{P}_{\hat{\theta}_{n}}}\bigg[\nabla_{\theta_{2}}\mathcal{A}_{\theta_{2}}^{1}\Big\langle\mathbb{E}_{\widetilde{X}\sim\mathbb{P}_{\hat{\theta}_{n}}}[\nabla_{\theta_{1}}\mathcal{A}_{\theta_{1}}^{2}\mathbf{K}(\widetilde{X}^{\prime},\widetilde{X})],I(\widetilde{X}_{j},\hat{\theta}_{n})\Big\rangle\bigg],I(\widetilde{X}_{j},\hat{\theta}_{n})\bigg\rangle\bigg|_{\theta_{1}=\theta_{2}=\hat{\theta}_{n}}$
$=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}\bigg\langle\Big\langle\mathbb{E}_{\widetilde{X},\widetilde{X}^{\prime}\sim\mathbb{P}_{\hat{\theta}_{n}}}\left[\nabla_{\theta}s_{\hat{\theta}_{n}}(X)\nabla_{\theta}s_{\hat{\theta}_{n}}(X^{\prime})\mathbf{K}(\widetilde{X}^{\prime},\widetilde{X})\right],I(\widetilde{X}_{j},\hat{\theta}_{n})\Big\rangle,I(\widetilde{X}_{i},\hat{\theta}_{n})\bigg\rangle$
$=\frac{1}{2n}\sum_{i=1}^{n}\sum_{j=1}^{n}\bigg\langle\Big\langle\nabla_{\theta}^{2}\mathrm{KSD}(\mathbb{P}_{\theta},\mathbb{P}_{\hat{\theta}_{n}})|_{\theta=\hat{\theta}_{n}},I(\widetilde{X}_{j},\hat{\theta}_{n})\Big\rangle,I(\widetilde{X}_{i},\hat{\theta}_{n})\bigg\rangle$ (79)
$=\frac{1}{2n}\sum_{i=1}^{n}\sum_{j=1}^{n}\bigg\langle\Big\langle H_{n},I(\widetilde{X}_{j},\hat{\theta}_{n})\Big\rangle,I(\widetilde{X}_{i},\hat{\theta}_{n})\bigg\rangle+o_{p}(1)$ (80)
$=\frac{1}{2n}\sum_{i=1}^{n}\sum_{j=1}^{n}I(\widetilde{X}_{j},\hat{\theta}_{n})^{\top}H_{n}I(\widetilde{X}_{i},\hat{\theta}_{n})+o_{p}(1),$ (81)
where Eqn. (79) is due to direct calculation on the derivative and in Eqn. (80) we use U-statistics law of large numbers [van2000asymptotic, Theorem 12.3]). Combining Eqn.(76), (78) and (81) directly yields that
$\sup_{f\in\mathcal{B}_{\mathcal{H}}(1)}\widetilde{S}_{n}(f)^{2}$ $=nT_{n}(\widetilde{X},\hat{\theta}_{n})+\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}G(\widetilde{X}_{j},\hat{\theta}_{n})^{\top}I(\widetilde{X}_{i},\hat{\theta}_{n})$
$\qquad\qquad\qquad+\frac{1}{2n}\sum_{i=1}^{n}\sum_{j=1}^{n}I(\widetilde{X}_{j},\hat{\theta}_{n})^{\top}H_{n}I(\widetilde{X}_{i},\hat{\theta}_{n})+o_{p}(1)$
$=nT_{n}(\widetilde{X},\widetilde{\theta}_{n})+o_{p}(1).$
H.3 Proof of Lemma 5
Proof of Lemma 5.
In [fernandez2024general, Theorem 1 ], the similar result is shown under standard weak convergence instead of stable convergence. We aim to translate all the conditions and result into stable convergence setting.
In Lemma 1, it is proved that stable convergence of random elements is equivalent to weak convergence holds for all measurable events $F$ with $\mathbb{P}(F)>0$, where $\mathbb{P}$ is the data generating measure. Thus, to get the stable convergence result, we only need Conditions (C1.1), (C1.2) and (C1.3) in Lemma 5 to hold under any measurable event $F$ with $\mathbb{P}(F)>0$.
Using Lemma 1 again, we can see Condition (C1.1) is equivalent to the weak convergence of the sequence $\widetilde{S}_{n}(f)$ under any $F$ with $\mathbb{P}(F)>0$. Condition (C1.2) is not a probabilistic statement, so no modification is needed here. Also, Condition (C1.3) is modified according to $F$. With all these modifications, we can directly apply [fernandez2024general, Theorem 1 ] to get the stable convergence result. ∎
H.4 Proof of Lemma 6
Proof of Lemma 6.
Rewrite the given condition as
$\lim_{n\to\infty}\mathbb{E}\bigl[f(X_{n})\,(X_{n}-C)\bigr]=0\quad\text{for all bounded continuous }f.$
We now show that for any $t>C$, $\lim_{n\to\infty}\mathbb{P}(X_{n}\geq t)=0$. Fix $\varepsilon=t-C>0$. Define a bounded continuous function $f_{\varepsilon}\colon\mathbb{R}\to\mathbb{R}$ by
$f_{\varepsilon}(x)\;=\;\begin{cases}-1,&x\leq C-\varepsilon,\\ \text{linear interpolation},&x\in[C-\varepsilon,\,C+\varepsilon],\\ +1,&x\geq C+\varepsilon.\end{cases}$
We apply contradiction method here. Suppose for some $t>C$, we have
$\limsup_{n\to\infty}\mathbb{P}(X_{n}\geq t)=\lim_{k\to\infty}\mathbb{P}(X_{n_{k}}\geq t)=\delta>0,$ (82)
where $n_{k}$ is a subsequence of $\{1,2,\cdots\}$. Then we obtain
$\mathbb{E}\bigl[f_{\varepsilon}(X_{n_{k}})\,(X_{n_{k}}-C)\bigr]$
$=\mathbb{E}\bigl[f_{\varepsilon}(X_{n_{k}})\,(X_{n_{k}}-C)\bm{1}(X_{n_{k}}>C)\bigr]+\mathbb{E}\bigl[f_{\varepsilon}(X_{n_{k}})\,(X_{n_{k}}-C)\bm{1}(X_{n_{k}}<C)\bigr]$
$\geq\mathbb{E}\bigl[f_{\varepsilon}(X_{n_{k}})\,(X_{n_{k}}-C)\bm{1}(X_{n_{k}}\geq t)\bigr]$
$\geq\varepsilon\cdot\mathbb{P}(X_{n_{k}}\geq t)\rightarrow\varepsilon\delta>0\quad\text{ as }k\rightarrow\infty,$
where in the first inequality, we use the fact that $f(x)\leq 0$ for $x\leq C$ and $f(x)\geq 0$ for all $x\geq C$. This directly contradicts the assumption in Lemma 6. Thus, we conclude that $\lim_{n\rightarrow\infty}\mathbb{P}(X_{n}\leq t)=0$.
By symmetric analysis, we also get that $\lim_{n\rightarrow\infty}\mathbb{P}(X_{n}\geq t)=1$ for all $t<C$. Thus, we conclude that $X_{n}\overset{d}{\to}C$, which is equivalent to $X_{n}\overset{p}{\to}C$.
H.5 Proof of Lemma 11
Proof of Lemma 11.
The difference between the statement of Lemma 11 and Proposition 6 in [fernandez2024general] is we only require convergence in probability in Eqn. (49) instead of almost sure convergence. However, it can be directly checked that the whole proof of Proposition 6 in [fernandez2024general] still goes through with this weaker condition.
In fact, we need an additional argument that if a sequence of almost surely bounded random variable $Y_{n}\overset{p}{\to}C_{4}$ for some constant $C_{4}$, then we have
$\limsup_{n\to\infty}\mathbb{E}[Y_{n}|F]\leq C_{4}\quad\forall F\subseteq\mathcal{F},\kern 5.0pt\mathbb{P}(F)>0.$ (83)
Proof of Statement (83).
For any $\varepsilon>0$, define event $\mathcal{E}_{n}^{\prime}=\{Y_{n}\leq C_{4}+\varepsilon\}$. Since $Y_{n}\overset{p}{\to}C_{4}$, we have $\mathbb{P}(\mathcal{E}_{n}^{\prime})\to 1$. Suppose $Y_{n}\leq B<\infty$ almost surely. Thus, we have
$\mathbb{E}[Y_{n}|F]$ $=\mathbb{E}[Y_{n}\mathbbm{1}(\mathcal{E}_{n}^{\prime})|F]+\mathbb{E}[Y_{n}\mathbbm{1}(\mathcal{E}_{n}^{c})|F]$
$\leq C_{4}+\varepsilon+(\sup Y_{n})\cdot\mathbb{P}(\mathcal{E}_{n}^{\prime}|F)$
$\leq C_{4}+\varepsilon+B\cdot\mathbb{P}(\mathcal{E}_{n}^{\prime})/\mathbb{P}(F)$
$\to C_{4}+\varepsilon.$
Since $\varepsilon$ is arbitrary, we have for any $F\subseteq\mathcal{F}$ with $\mathbb{P}(F)>0$,
$\limsup_{n\to\infty}\mathbb{E}[Y_{n}|F]\leq C_{4}.$
In the original proof of Proposition 6 in [fernandez2024general], the Statement (83) is derived used under almost sure convergence assumption. Therefore, the original proof is still valid under convergence in probability.
In order to apply Proposition 6 in [fernandez2024general] to prove Lemma 11, we pick $\mathcal{G}=\mathcal{H}$, $Q=\mathrm{id}$ and $L=K$ in the original statement. The whole proof goes through, which straightly implies Condition (C1.3) holds. ∎
H.6 Proof of Lemma 12
Proof of Lemma 12.
We apply Le Cam’s first lemma [van2000asymptotic, Lemma 6.4] to prove this result. By Assumption 2, we have
$\mathbb{E}_{\theta_{0}}\left[\frac{\mathrm{d}\mathcal{L}_{n}}{\mathrm{d}P_{\theta_{0}}}\right]=\frac{1}{Z_{n}}\mathbb{E}_{\theta_{0}}\left[1+\frac{h(X)}{\sqrt{n}}\right]=\frac{1+\mathbb{E}_{\theta_{0}}[h(X)]/\sqrt{n}}{Z_{n}}=1,$
which implies $\mathcal{L}_{n}\lhd\mathbb{P}_{\theta_{0}}$. Conversely, under measure $X\sim\mathcal{L}_{n}$, we conclude
$\frac{\mathrm{d}\mathcal{L}_{n}}{\mathrm{d}P_{\theta_{0}}}=\frac{1}{Z_{n}}\left(1+\frac{h(X)}{\sqrt{n}}\right)\overset{d}{\to}1$
by noticing that $Z_{n}\to 1$ and $h(X)$ is a finite random variable, so it also holds that $\mathcal{L}_{n}\rhd\mathbb{P}_{\theta_{0}}$. Thus measures $P_{\theta_{0}}$ and $\mathcal{L}_{n}$ are mutually contiguous. Since under measure $\mathbb{P}_{\theta_{0}}$, Assumption 1 tells us
$\left|\sqrt{n}(\hat{\theta}_{n}-\theta_{0})-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}I(X_{i},\theta_{0})\right|\overset{p}{\to}0,$
Le Cam’s first lemma confirms that it also holds under measure $\mathcal{L}_{n}$, which is exactly Eqn. (55).
To further derive the asymptotic distribution of the estimator $\hat{\theta}_{n}$, we need to apply the Le Cam’s third lemma [van2000asymptotic, Theorem 6.6]. For simplicity, we denote
$\frac{\mathrm{d}\mathcal{L}_{n}}{\mathrm{d}P_{\theta_{0}}}(X)\equiv\prod_{i=1}^{n}\left[\frac{\mathrm{d}\mathcal{L}_{n}}{\mathrm{d}P_{\theta_{0}}}(X_{i})\right].$
It holds that under measure $\mathbb{P}_{\theta_{0}}$,
$\log\left(\frac{\mathrm{d}\mathcal{L}_{n}}{\mathrm{d}P_{\theta_{0}}}(X)\right)$ $=\sum_{i=1}^{n}\log\left(1+\frac{h(X_{i})}{\sqrt{n}}\right)-n\log Z_{n}$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(h(X_{i})-\mathbb{E}_{\theta_{0}}h(X)\right)-\frac{1}{2n}\sum_{i=1}^{n}h(X_{i})^{2}+\frac{1}{2}\left(\mathbb{E}_{\theta_{0}}[h(X)]\right)^{2}$
$\qquad\qquad\qquad\qquad\qquad+O_{p}\left(\frac{1}{n^{3/2}}\sum_{i=1}^{n}h(X_{i})^{3}+\frac{1}{n^{1/2}}\left(\mathbb{E}_{\theta_{0}}[h(X)]\right)^{3}\right),$
where we use the Taylor expansion $\log(1+x)=x-x^{2}/2+O(x^{3})$ in the second line. To address remaining term, we first notice that by strong law of large numbers [durrett2019probability, Theorem 2.4.1], we have
$\lim_{n\to\infty}\frac{1}{2n}\sum_{i=1}^{n}h(X_{i})^{2}=\frac{1}{2}\mathbb{E}_{\theta_{0}}[h(X)^{2}]\quad\text{almost surely}.$ (84)
Then by Lemma 21 with choice $p=2/3$ in the lemma, we obtain
$\frac{1}{n^{3/2}}\sum_{i=1}^{n}h(X_{i})^{3}\to 0\quad\text{almost surely}.$ (85)
Combining Eqns. (84), (85) and the fact that $\mathbb{E}_{\theta_{0}}[h(X)]<\infty$, we conclude
$\log\left(\frac{\mathrm{d}\mathcal{L}_{n}}{\mathrm{d}P_{\theta_{0}}}(X)\right)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(h(X_{i})-\mathbb{E}_{\theta_{0}}h(X)\right)-\frac{1}{2}\mathrm{Var}_{\theta_{0}}[h(X)]+o_{p}(1),$ (86)
Under measure $\mathbb{P}_{\theta_{0}}$, we have the following jointly weak convergence
$\begin{pmatrix}\sqrt{n}(\hat{\theta}_{n}-\theta_{0})\\ \log\left(\frac{\mathrm{d}\mathcal{L}_{n}}{\mathrm{d}P_{\theta_{0}}}(X)\right)\end{pmatrix}$ $=\begin{pmatrix}0\\ -\frac{1}{2}\mathrm{Var}_{\theta_{0}}[h(X)]\end{pmatrix}+\begin{pmatrix}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}I(X_{i},\theta_{0})\\ \frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(h(X_{i})-\mathbb{E}_{\theta_{0}}h(X)\right)\end{pmatrix}+o_{p}(1)$
$\overset{d}{\rightarrow}\mathcal{N}\left(\begin{pmatrix}0\\ -\frac{1}{2}M_{2}^{h}\end{pmatrix},\begin{pmatrix}\Sigma&\tau\\ \tau^{\top}&M_{2}^{h}\end{pmatrix}\right)$,
where $M_{2}^{h}=\mathrm{Var}_{\theta_{0}}[h(X)]$ and $\tau=\mathrm{Cov}_{\theta_{0}}[I(X,\theta_{0}),h(X)]$. Eqn. (86) is applied in the first line, and the standard CLT [durrett2019probability, Theorem 3.4.1] is used in the second line. By Le Cam’s third lemma, when under measure $\mathcal{L}_{n}$,
$\sqrt{n}(\hat{\theta}_{n}-\theta_{0})\overset{d}{\rightarrow}\mathcal{N}(\tau,\Sigma),$ (87)
which concludes Eqn. (56).
H.7 Proof of Lemma 13
Proof of Lemma 13.
The proof here is similar to the proof of Lemma 4, where we derive the intermediary form of expansion. Thus, we state the intermediary results and omit the detailed proof.
We introduce the following notation. Recall that $F_{\theta}(x,x^{\prime})=\nabla_{\theta}h_{\theta}(x,x^{\prime})$, and we define $G_{n}(x,\theta)\equiv G_{\mathcal{L}_{n}}(x,\theta)=\mathbb{E}_{X\sim\mathcal{L}_{n}}F_{\theta}(x,X)$. Under local alternative hypothesis $H_{1,n}$, we apply the similar calculation in the proof of Statement (67) to obtain
$nT_{n}=nT_{n}(X,\theta_{0})$ $+\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}G_{n}(X_{j},\theta_{0})^{\top}I(X_{i},\theta_{0})$
$\qquad\qquad+\frac{1}{2n}\sum_{i=1}^{n}\sum_{j=1}^{n}I(X_{i},\theta_{0})^{\top}H_{0}I(X_{i},\theta_{0})+o_{p}(1),$ (88)
where $H_{0}=\nabla_{\theta}^{2}T_{n}(X,\theta)|_{\theta=\theta_{0}}$. Furthermore, the calculation in the proof of Statement (73) gives that
$\sup_{f\in\mathcal{B}_{\mathcal{H}}(1)}S_{1,n}(f)^{2}=nT_{n}(X,\theta_{0})$ $+\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}G_{n}(X_{j},\theta_{0})^{\top}I(X_{i},\theta_{0})$
$\qquad+\frac{1}{2n}\sum_{i=1}^{n}\sum_{j=1}^{n}I(X_{i},\theta_{0})^{\top}H_{0}I(X_{i},\theta_{0})+o_{p}(1),$ (89)
Combining Eqn. (88) and Eqn. (89) proves Lemma 13. ∎
Appendix I Auxiliary lemmas
Lemma 15.
Under Assumption 1, for any $0\leq\gamma\leq 1$, let $\theta_{\gamma}=\gamma\theta_{0}+(1-\gamma)\hat{\theta}_{n}$ and $\hat{\theta}_{\gamma}=\gamma\hat{\theta}_{n}+(1-\gamma)\widetilde{\theta}_{n}$, it holds that
$\|\nabla_{\theta}^{2}T_{n}(X,\theta_{0})-\nabla_{\theta}^{2}T_{n}(X,\theta_{\gamma})\|_{2}=o_{p}(1),$ (90)
Moreover, for the bootstrap statistics, we have
$\|\nabla_{\theta}^{2}T_{n}(\widetilde{X},\hat{\theta}_{n})-\nabla_{\theta}^{2}T_{n}(\tilde{X},\hat{\theta}_{\gamma})\|_{2}=o_{p}(1).$ (91)
Proof of Lemma 15.
Recall the KSD kernel $h_{p}(\cdot,\cdot)$ defined in Eqn. (7). We write $h_{\mathbb{P}_{\theta}}$ as $h_{\theta}$ for simplicity in the proof. By Newton-Leibniz formula, we have
$\nabla_{\theta}^{2}T_{n}(X,\theta_{0})-\nabla_{\theta}^{2}T_{n}(X,\theta_{\gamma})$ $=\int_{0}^{1-\gamma}\nabla_{\theta}^{3}T_{n}(X,\theta_{0}+t(\hat{\theta}_{n}-\theta_{0}))\cdot(\theta_{0}-\theta_{\gamma})\mathrm{d}t,$
which leads to
$\left\|\nabla_{\theta}^{2}T_{n}(X,\theta_{0})-\nabla_{\theta}^{2}T_{n}(X,\theta_{\gamma})\right\|_{2}$ $\leq\int_{0}^{1-\gamma}\left\|\nabla_{\theta}^{3}T_{n}(X,\theta_{0}+t(\hat{\theta}_{n}-\theta_{0}))\right\|_{2}\cdot\|\theta_{0}-\theta_{\gamma}\|_{2}\mathrm{d}t$
$\leq\int_{0}^{1}\left\|\nabla_{\theta}^{3}T_{n}(X,\theta_{0}+t(\hat{\theta}_{n}-\theta_{0}))\right\|_{2}\cdot\|\hat{\theta}_{n}-\theta_{0}\|_{2}\mathrm{d}t.$
Using the definition of $h_{p}$ and Assumption 4, we obtain the expression
$\|\nabla_{\theta}^{3}T_{n}(X,\theta)\|_{2}$ $\leq\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\|\nabla_{\theta}^{3}h_{\theta}(X_{i},X_{j})\|_{2},$
which implies the bound on the operator norm of of the term above can be derive from the moment bound on $\nabla^{3}h_{\theta}(X_{i},X_{j})$. For any $\theta\in\Theta$, we have
$\mathbb{E}\left[\left\|\nabla_{\theta}^{3}h_{\theta}(X_{i},X_{j})\right\|_{2}\right]\lesssim(C_{1}+C_{2}+1)C_{3}C_{K}<\infty,$
where Hölder inequality is applied, and some numerical constants are omitted. The moment bound above combined with Markov inequality reveals that $\|\nabla_{\theta}^{3}T_{n}(X,\theta)\|_{2}=o_{p}(n^{1/2})$ uniformly in $\theta\in\Theta$. Thus, we achieve
$\left\|\nabla_{\theta}^{2}T_{n}(X,\theta_{0})-\nabla_{\theta}^{2}T_{n}(X,\theta_{\gamma})\right\|_{2}$ $\leq\left(\int_{0}^{1}\left\|\nabla_{\theta}^{3}T_{n}(X,\theta_{0}+t(\hat{\theta}_{n}-\theta_{0}))\right\|_{2}\mathrm{d}t\right)\cdot\|\hat{\theta}_{n}-\theta_{0}\|_{2}$
$=o_{p}(n^{1/2})\cdot O_{p}(n^{-1/2})=o_{p}(1),$
where the last line is due to Assumption 1 that $\|\hat{\theta}_{n}-\theta_{0}\|_{2}=o_{p}(1)$. Thus, we conclude Eqn. (90) holds.
For the second part, we can apply the similar argument as above. The only distinction is that we need $\|\hat{\theta}_{n}-\widetilde{\theta}_{n}\|=o_{p}(1)$, which is proved in Appendix H.2, Step 1. With this additional result, we can directly apply Lemma 17 to obtain Eqn. (91). ∎
Lemma 16.
Suppose we have $X_{i}\overset{i.i.d.}{\sim}\mathcal{L}$ and denote $G_{\mathcal{L}}(x,\theta)=\mathbb{E}_{X\sim\mathcal{L}}[F_{\theta}(x,X)]$. Then we have
$\sqrt{n}\nabla_{\theta}T_{n}(X,\theta)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}G_{\mathcal{L}}(X_{i},\theta)+o_{p}\left(1\right).$ (92)
Specifically, conditioning on a fixed sequence of data $\mathcal{D}_{n}$, we have
$\sqrt{n}\nabla_{\theta}T_{n}(\widetilde{X},\hat{\theta}_{n})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}G_{\mathbb{P}_{\hat{\theta}_{n}}}(\widetilde{X}_{i})+o_{p}\left(1\right),$ (93)
and it also holds unconditionally.
Proof of Lemma 16.
By definition, we have
$\sqrt{n}\nabla_{\theta}T_{n}(X,\theta)$ $=\sqrt{n}\left(\frac{1}{n^{2}}\sum_{i=1}^{n}\nabla_{\theta}h_{\theta}(X_{i},X_{i})+\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j\neq i}\nabla_{\theta}h_{\theta}(X_{i},X_{j})\right)$
$=\frac{1}{n^{3/2}}\sum_{i=1}^{n}\nabla_{\theta}h_{\theta}(X_{i},X_{i})+\frac{1}{n^{3/2}}\sum_{i=1}^{n}\sum_{j\neq i}\nabla_{\theta}h_{\theta}(X_{i},X_{j})$
$=\frac{1}{\sqrt{n}(n-1)}\sum_{i=1}^{n}\sum_{j\neq i}\nabla_{\theta}h_{\theta}(X_{i},X_{j})+o_{p}(1),$ (94)
where the last equality is due to the fact that $\mathbb{E}[\nabla_{\theta}h_{\theta}(X_{i},X_{i})]<\infty$, which can be directly obtained from Assumption 4.
By the central limit theorem of U-statistic [van2000asymptotic, Theorem 12.3], we have
$\frac{1}{\sqrt{n}(n-1)}\sum_{i=1}^{n}\sum_{j\neq i}\nabla_{\theta}h_{\theta}(X_{i},X_{j})$ $=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\mathbb{E}_{X\sim\mathcal{L}}[\nabla_{\theta}h_{\theta}(X_{i},X)]+o_{p}(1)$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}G_{\mathcal{L}}(X_{i},\theta)+o_{p}\left(1\right).$ (95)
Combining Eqn. (94) and Eqn. (95), we obtain the desired result in Eqn. (92). The Eqn. (93) can be directly obtained by applying Lemma 17 and repeat the proof above. ∎
To be consistent with the notation in [niu2022reconciling], we use the notation $W_{n}|\mathcal{F}_{n}\overset{d,p}{\to}W$ to indicate that $W_{n}$ converges in distribution to a random variable $W$ conditionally on $\mathcal{F}_{n}$, i.e., for all $t\in\mathbb{R}$, at which the CDF of $W$ is continuous, we have
$\mathbb{P}[W_{n}\leq t|\mathcal{F}_{n}]\overset{p}{\to}\mathbb{P}[W\leq t].$ (96)
Lemma 17 (Conditional convergence implies unconditional convergence).
For a sequence of random variables $W_{n}$ and a fixed continuous random variable $W$,
$W_{n}|\mathcal{F}_{n}\overset{d,p}{\to}W\quad\text{implies}\quad W_{n}\overset{d}{\to}W.$
Proof of Lemma 17.
By iterated expectation formula, we have
$\mathbb{E}[\mathbb{P}[W_{n}\leq t|\mathcal{F}_{n}]]=\mathbb{E}[\mathbbm{1}(W_{n}\leq t)]=\mathbb{P}[W_{n}\leq t].$ (97)
Notice that $\mathbb{P}[W_{n}\leq t|\mathcal{F}_{n}]\leq 1$. We apply the bounded convergence theorem [durrett2019probability, Theorem 1.5.3] to obtain
$\lim_{n\to\infty}\mathbb{E}[\mathbb{P}[W_{n}\leq t|\mathcal{F}_{n}]]=\mathbb{E}[\mathbb{P}[W\leq t]]=\mathbb{P}[W\leq t].$ (98)
Combining Eqn. (97) and Eqn. (98), we have
$\lim_{n\to\infty}\mathbb{P}[W_{n}\leq t]=\mathbb{P}[W\leq t],\quad\forall t\in\mathbb{R},$
which implies the unconditional convergence in distribution.
Lemma 18 (Lemma 5, [niu2022reconciling]).
Let $W_{n}$ be a sequence of random variables and $W$ be a continuous random variable. If $W_{n}|\mathcal{F}_{n}\overset{d,p}{\to}W$, we have
$\limsup_{n\rightarrow\infty}\mathbb{P}\left[\sup_{t\in\mathbb{R}}\Big|\mathbb{P}[W_{n}\leq t|\mathcal{F}_{n}]-\mathbb{P}[W\leq t]\Big|>\delta\right]=0.$
Lemma 19 (Lemma 1, [niu2022reconciling]).
Let $W_{n}$ be a sequence of random variables and $\alpha\in(0,1)$. If $W_{n}|\mathcal{F}_{n}\overset{d,p}{\to}W$, and $W$ is a continuous random variable with strictly increasing CDF, then we have
$\mathbb{Q}_{\alpha}(W_{n})\overset{p}{\to}\mathbb{Q}_{\alpha}(W).$
Lemma 20 (Theorem 6, [niu2022reconciling]).
Let $W_{n}$ be a sequence of random variables satisfies $W_{n}|\mathcal{F}_{n}\overset{d,p}{\to}W$. For any random variable $\epsilon_{n}$ satisfies $\epsilon_{n}\overset{p}{\to}0$, we have
$W_{n}+\epsilon_{n}|\mathcal{F}_{n}\overset{d,p}{\to}W.$
Lemma 21 (Marcinkiewicz-Zygmund strong law of large numbers, Theorem 2.5.12 and Theorem 2.5.13, [durrett2019probability]).
Let $\{X_{i}\}_{i=1}^{n}$ be i.i.d. random variables with $\mathbb{E}[|X_{i}|^{p}]<\infty$ for some $0<p<2$, then
$n^{-1/p}\sum_{i=1}^{n}(X_{i}-\mathbb{E}[X_{i}])\to 0\quad\text{almost surely if }\kern 5.0pt1<p<2.$
$n^{-1/p}\sum_{i=1}^{n}X_{i}\overset{a.s.}{\to}0\quad\text{almost surely if }\kern 5.0pt0<p<1.$
Appendix J Additional results on experiments
J.1 Testing normality
We provide detailed experimental setups on the normality test in Section 5.1 here. We consider the following four different distributions to generate data:
  1. 1.
    Gaussian distribution $\mathcal{N}(\mu,1)$, which is under $H_{0}$ for any $\mu\in\mathbb{R}$.
  2. 2.
    Non-centered Student-$t$ distribution with degree of freedom $\nu$ and mean shift $\mu=10/(\nu+1)$. The density function is given by $p(x)\propto(1+(x-\mu)^{2}/\nu)^{-\nu+1/2}$, $x\in\mathbb{R}$.
  3. 3.
    Mixture of two Gaussians, i.e., $w\mathcal{N}(0,1)+(1-w)\mathcal{N}(\delta,(1+\delta)^{2})$.
  4. 4.
    Non-centered generalized $\chi^{2}$ distribution with power parameter $\alpha$ and a mean shift, where the density function is given by $p(x)\propto x^{\alpha-1}\exp(-(x-1)^{2}/2)$, $x\in\mathbb{R}$.
J.2 Order detection for the kernel exponential family
We provide additional details on the order detection for the kernel exponential family in Section 5.2. The and null models and data generating processes in the two experiments are set as follows:
  1. 1.
    In the first setting, we set $p=2$ and $\theta_{1}=10$, $\theta_{2}$ is varied from 4-4 to $4$. The null model is the kernel exponential family with $p=1$ and $\theta_{1}\in\mathbb{R}$. Thus, when $\theta_{2}=0$, the data generating distribution is from the null model, otherwise it is from the alternative model.
  2. 2.
    In the second setting, we set $p=5$ to generate data. The parameters are varied for different instances and $\theta_{2},\ldots,\theta_{5}$ are randomly generated from $\mathrm{Unif}[-10,10]$. We aim to evaluate the power of the SKSD test when the size of null model increases.
To sample from the kernel exponential family, we apply metroplis-adjusted Langevin algorithm (MALA) with standard Gaussian as the proposal distribution. The burn-in period is set as $10^{4}$ iterations, and we collect samples every $20$ iterations afterwards.
For our choice of Gaussian kernel in the kernel exponential density family, the basis functions $\{\phi_{\ell}\}_{\ell=1}^{\infty}$ can be explicitly expressed as
$\phi_{\ell}(x)=\frac{x^{\ell}}{\sqrt{\ell!}}\exp\left(-\frac{x^{2}}{2\sigma^{2}}\right),\quad\ell=1,2,\ldots,$
which allows us to compute the score function $\nabla_{x}\log p_{\theta}(x)$ in closed form.
J.3 Conditional Gaussian graphical model
In this section, we provide additional details on the conditional Gaussian graphical model experiment in Section 5.3. The data generating process (DGP) is set as follows: for different $d$, we set $\gamma^{(1)}_{l}=2$ and $\gamma^{(2)}_{k}=-0.5$. For the interaction terms, we consider the 1D grid structure with reach $r=2$:
$\Sigma_{ij}$ $=\begin{cases}-w_{i},&i-j=1\quad\mathrm{mod}\ d,\\ -w_{j},&i-j=d-1\quad\mathrm{mod}\ d,\\ -10^{-2}\varepsilon,&|i-j|=2\quad\mathrm{mod}\ d,\\ 0,&\text{otherwise},\end{cases}$
where $w=(w_{1},\ldots,w_{d})$ indicates the interaction magnitude between adjacent nodes in the graph. We consider two different settings for $w$:
  1. 1.
    For the first experiment, we fix $d=8$ and independently sample $w_{i}\sim\mathrm{Unif}([2,2+d/2])$ for $i\in[d]$. We vary $\varepsilon$ from 0 to $1.5$ to generate data from both null and alternative models, where the null model corresponds to $\varepsilon=0$.
  2. 2.
    For the second experiment, we set $w_{i}=2$ for $i\in[d]$, $\varepsilon=0.5$, and vary $d\in\mathbb{Z}^{+}$ to generate data from different distributions. Here, all the data generating distributions are from alternative models.
Throughout both experiments, we take $\gamma^{(1)}$ and $\gamma^{(2)}$ as fixed, known parameters. Gibbs sampling is used to sample from the model for both data generation and bootstrap procedure. We set the burn-in period as $10^{4}$ iterations and collect samples every $20$ iterations afterwards.
For completeness of the properties of this model, we establish the conditional distributions for the conditional Gaussian graphical model. For $i=1,\ldots,d$, we have
$p_{\theta}(x^{(i)},x^{(-i)})=\frac{1}{Z}\exp\left(\sum_{1\leq i\neq j\leq d}\Sigma_{ij}(x^{(i)})^{2}(x^{(j)})^{2}+\sum_{k=1}^{d}\gamma_{k}^{(2)}(x^{(k)})^{2}+\sum_{l=1}^{d}\gamma_{\ell}^{(1)}x^{(\ell)}\right)$
$=\frac{1}{Z}\exp\Bigg(\left(\sum_{j\neq i}2\Sigma_{ij}(x^{(j)})^{2}+\gamma_{i}^{(2)}\right)(x^{(i)})^{2}+\gamma_{i}^{(1)}x^{(i)}$
$\qquad\qquad\qquad+\sum_{\begin{subarray}{c}1\leq i^{\prime}\neq j^{\prime}\leq d,\\ i^{\prime}\neq i,j^{\prime}\neq i\end{subarray}}\Sigma_{ij}(x^{(i^{\prime})})^{2}(x^{(j^{\prime})})^{2}+\sum_{\begin{subarray}{c}1\leq k\leq d,\\ k\neq i\end{subarray}}\gamma_{k}^{(2)}(x^{(k)})^{2}+\sum_{\begin{subarray}{c}1\leq\ell\leq d,\\ \ell\neq i\end{subarray}}\gamma_{\ell}^{(1)}x^{(\ell)}\Bigg),$
for some normalizing constant $Z$. By Bayes’ rule, we have
$p_{\theta}(x^{(i)}|x^{(-i)})\propto\exp\left(\left(\sum_{j\neq i}2\Sigma_{ij}(x^{(j)})^{2}+\gamma_{i}^{(2)}\right)(x^{(i)})^{2}+\gamma_{i}^{(1)}x^{(i)}\right)$
$\Leftrightarrow X^{(i)}|X^{(-i)}\sim\mathcal{N}\left(-\frac{\gamma_{i}^{(1)}}{2\left(\sum_{j\neq i}2\Sigma_{ij}(x^{(j)})^{2}+\gamma_{i}^{(2)}\right)},-\frac{1}{2\left(\sum_{j\neq i}2\Sigma_{ij}(x^{(j)})^{2}+\gamma_{i}^{(2)}\right)}\right),$
which gives the conditional distribution of $X^{(i)}$ given $X^{(-i)}$.
When $\Sigma_{ij}<0$ and $\gamma_{i}^{(2)}<0$, we can check that
$Z=\int_{\mathbb{R}^{d}}\exp\left(\sum_{1\leq i\neq j\leq d}\Sigma_{ij}(x^{(i)})^{2}(x^{(j)})^{2}+\sum_{k=1}^{d}\gamma_{k}^{(2)}(x^{(k)})^{2}+\sum_{l=1}^{d}\gamma_{\ell}^{(1)}x^{(\ell)}\right)\mathrm{d}x$
$<\int_{\mathbb{R}^{d}}\exp\left(\sum_{k=1}^{d}\gamma_{k}^{(2)}(x^{(k)})^{2}+\sum_{l=1}^{d}\gamma_{\ell}^{(1)}x^{(\ell)}\right)\mathrm{d}x$
$=\prod_{k=1}^{d}\int_{\mathbb{R}}\exp\left(\gamma_{k}^{(2)}(x^{(k)})^{2}+\gamma_{k}^{(1)}x^{(k)}\right)\mathrm{d}x^{(k)}$
$=\prod_{k=1}^{d}\sqrt{\frac{\pi}{-\gamma_{k}^{(2)}}}\exp\left(-\frac{(\gamma_{k}^{(1)})^{2}}{4\gamma_{k}^{(2)}}\right)<\infty,$
which confirms that the density $p_{\theta}$ is well-defined. When $\Sigma_{ij}>0$ for some $j\neq i$, the density $p_{\theta}(x^{(i)}\mid x^{(-i)})$ is not integrable when $x^{(j)}\to\infty$, thus $p_{\theta}$ is not a valid density function. Similar argument applies to the case when $\gamma_{i}^{(2)}>0$, where the density $p_{\theta}(x^{(i)}\mid x^{(-i)})$ is not integrable when $x^{(i)}\to 0$. Thus, when $\Sigma_{ij}>0$ for any $i\neq j$ or $\gamma_{i}^{(2)}>0$ for some $i$, the density $p_{\theta}$ is not well-defined.
In this setting, both the minimum-KSD and the implicit score matching estimators admit closed-form solutions, which are given by the following: Denote
$t(x)$ $=\mathrm{vec}\left((x^{(i)}x^{(j)})^{2}\right)_{1\leq i\leq j\leq d}\in\mathbb{R}^{d^{2}},$
$b(x)$ $=\sum_{k=1}^{d}\gamma_{k}^{(2)}(x^{(k)})^{2}+\sum_{l=1}^{d}\gamma_{\ell}^{(1)}x^{(\ell)}\in\mathbb{R},$
where $\mathrm{vec}(\cdot)$ is the vectorization operator that stacks the columns of a matrix into a vector, and $b(x)=\sum_{k=1}^{d}\gamma_{k}^{(2)}x_{k}^{2}+\sum_{l=1}^{d}\gamma_{l}^{(1)}x_{i}\in\mathbb{R}$. Then, the estimators are given by
$\mathrm{vec}(\hat{\Sigma}_{\mathrm{KSD}})$ $=-\left(\sum_{i,j=1}^{n}K(x_{i},x_{j})\nabla t(x_{i})\nabla t(x_{j})^{\top}\right)^{-1}$
$\qquad\qquad\left(\sum_{i,j=1}^{n}K(x_{i},x_{j})\nabla t(x_{i})\nabla b(x_{j})^{\top}+\nabla t(x_{i})\nabla_{2}K(x_{i},x_{j})^{\top}\right),$
$\mathrm{vec}(\hat{\Sigma}_{\mathrm{SM}})$ $=-\left(\sum_{i=1}^{n}\nabla t(x_{i})\nabla t(x_{i})^{\top}\right)^{-1}\left(\sum_{i=1}^{n}\Delta t(x_{i})+\nabla t(x_{i})\nabla b(x_{i})^{\top}\right),$
where $\nabla t(x)$, $\Delta t(x)$ are the gradient and Laplacian of $t(x)$, respectively, and $\nabla_{2}K(x,x^{\prime})$ is the gradient of $K(x,x^{\prime})$ with respect to the second argument $x^{\prime}$.
Since the model is a little more complex, we validate the assumptions for this model. We first choose some convex and open set $\Theta\subseteq\mathbb{R}^{d^{2}+2d}$ with compact closure such that it is large enough and contains the true parameter $\Sigma$. To ensure Assumption 3, we project our estimators onto the set $\Theta$ if the estimators fall outside of $\Theta$. In numerical experiments, the estimators almost surely fall inside of $\Theta$ when $n$ is large enough, so the projection step does not change the estimators in practice. The score function for this model is given by
$s_{\theta}(x)$ $=\nabla_{x}\log p_{\theta}(x)$
$=\begin{pmatrix}\sum_{j\neq 1}4\Sigma_{1j}x^{(1)}(x^{(j)})^{2}+2\theta_{1}^{(2)}x^{(1)}+\theta_{1}^{(1)}\\ \sum_{j\neq 2}4\Sigma_{2j}x^{(2)}(x^{(j)})^{2}+2\theta_{2}^{(2)}x^{(2)}+\theta_{2}^{(1)}\\ \vdots\\ \sum_{j\neq d}4\Sigma_{dj}x^{(d)}(x^{(j)})^{2}+2\theta_{d}^{(2)}x^{(d)}+\theta_{d}^{(1)}\end{pmatrix}$.
We then verify Assumption 4 by checking the moment conditions on $s_{\theta}(x)$ and its derivatives.
$\mathbb{E}_{X\sim p_{\theta}}\left[\|s_{\theta}(X)\|^{2}\right]$
$=\int_{\mathbb{R}^{d}}\|s_{\theta}(x)\|^{2}p_{\theta}(x)\,\mathrm{d}x$
$=\int_{\mathbb{R}^{d}}\left(\sum_{i=1}^{d}\left(\sum_{j\neq i}4\Sigma_{ij}x^{(i)}(x^{(j)})^{2}+2\theta_{i}^{(2)}x^{(i)}+\theta_{i}^{(1)}\right)^{2}\right)p_{\theta}(x)\,\mathrm{d}x$
$=\int_{\mathbb{R}^{d}}\left(\sum_{i=1}^{d}\left(\sum_{j\neq i}4\Sigma_{ij}x^{(i)}(x^{(j)})^{2}+2\theta_{i}^{(2)}x^{(i)}+\theta_{i}^{(1)}\right)^{2}\right)$
$\qquad\qquad\cdot\frac{1}{Z}\exp\left(\sum_{1\leq i\neq j\leq d}\Sigma_{ij}(x^{(i)})^{2}(x^{(j)})^{2}+\sum_{k=1}^{d}\theta_{k}^{(2)}(x^{(k)})^{2}+\sum_{l=1}^{d}\theta_{\ell}^{(1)}x^{(\ell)}\right)\,\mathrm{d}x$
For fixed $x\in\mathbb{R}^{d}$, the integrand function is continuous in $\Sigma_{ij}$ for all $1\leq i\neq j\leq d$. Moreover, the partial derivatives of the integrand with respect to $\Sigma_{ij}$ exist and its $L_{1}$ norm is uniformly bounded, which can be verified in a similar way as the proof of well-definedness of $p_{\theta}$ above. Thus, by Fubini’s theorem, $\mathbb{E}_{X\sim p_{\theta}}\left[\|s_{\theta}(X)\|^{2}\right]$ is continuous in $\theta$. Since the closure of the parameter space $\Theta$ is compact, $\mathbb{E}_{X\sim p_{\theta}}\left[\|s_{\theta}(X)\|^{2}\right]$ is uniformly bounded in the parameter space $\Theta$. Similar argument applies to the other moment conditions in Assumption 4, thus we conclude that Assumption 4 holds. The Assumption 5 holds since we choose Gaussian kernel in this task, so we omit the details.
Appendix K Local power validation
In Section 4, we established the theory for asymptotic distribution of the SKSD tests under both multiplicative and additive local alternative hypotheses. In this section, we conduct simulations to evaluate the finite-sample performance of our tests under local alternatives.
Experiment setup.
In this experiment, the null model $H_{0}$ is the Gaussian model $\mathcal{N}(\mu,\sigma^{2})$, where $\mu\in\mathbb{R}$ and $\sigma^{2}\in\mathbb{R}_{++}$ are the mean and variance of the Gaussian distribution, respectively. Two families of data generation processes (DGPs) are considered under the alternative hypothesis $H_{1}$:
  1. 1.
    For the multiplicative local alternatives, we consider the generalized $\chi^{2}$ distribution with power parameter $\alpha$, where the density function is given by $p(x)\propto x^{\alpha-1}\exp(-x^{2}/2)$, $x\in\mathbb{R}$.
  2. 2.
    For the additive local alternatives, we consider the mixture of Gaussian distribution $w\cdot\mathcal{N}(2,3)+(1-w)\cdot\mathcal{N}(0,1)$, where $w\in(0,1)$ is the mixing weight.
The local alternatives degenerate to the null model when $\alpha\to 1$ for the multiplicative case and $w\to 0$ or $1$ for the additive case. In the experiment, we start with the degenerated case mentioned above, and then vary the value of $\alpha$ and $w$ to evaluate the local power of the tests.
We perform the tests with level to be $0.05$, and calibrated the $95\%$-cutoff for all four test statistics using the parametric bootstrap procedure with $B=300$ time resamples and choose the resample size to be $n$ in order to balance the computation cost and power. The testing Type-I error and power are estimated based on $500$ replications. We use this bootstrap test setting by default for the following experiments.
Testing Methods.
The estimation and testing procedures used in this experiment:
  1. 1.
    Estimation: minimum-KSD estimator and MLE are used to estimate the parameters of the null model $\mathcal{N}(\mu,\sigma^{2})$.
  2. 2.
    Testing: We consider the SKSD test with $95\%$-parametric bootstrap cutoff.
Figure 4: Power curves of SKSD goodness-of-fit tests under local alternatives with different estimation procedures. The null model is the Gaussian model $\mathcal{N}(\mu,\sigma^{2})$ Left panel: multiplicative local alternatives; Right panel: additive local alternatives. The red dash line indicates the test level.
Results.
In Figure 4, we present the power curves of SKSD tests under both multiplicative and additive local alternatives with different estimation procedures. When the observed data is generated from the null model, all tests control the Type-I error at the nominal level $0.05$. In the left panel of Figure 4, we consider the multiplicative local alternatives with data generated from generalized $\chi^{2}$ distributions. When $\alpha$ increases from $1$, the observed data become more heavy tailed, thus more distinguishable from the null model, and the power of both tests increases to nearly 1, which is consistent with the theoretical justification. In the right panel of Figure 4, we consider the additive local alternatives with data generated from the mixture of Gaussian distributions. When $w\in(0,1/2)$ increases, the observed data become more apparently multi-modal, thus the power of both tests increases, which again validates our theoretical results. In addition, we can see that the SKSD test with MLE has a slightly larger power than the one with minimum-KSD estimator in the multiplicative case, which confirms the influence of the estimation procedure on the power of the tests.
Appendix L Power comparison across score-based tests
In Section 2, we discussed the connection between score-based tests and distance-based tests, especially the IPM-based tests. We focus on SKSD tests since it can be viewed as a special case of IPM-based tests with computation efficiency under certain circumstances. However, the finite-sample power comparison between SKSD tests and other distance-based tests is still under-explored. In this section, we compare the power of our proposed goodness-of-fit tests with different distance-based matrices for composite null hypothesis.
Experiment setup.
In this experiment, the null model $H_{0}$ is the Gaussian model $\mathcal{N}(\mu,\sigma^{2})$. Two families of data generation processes are considered under the alternative hypothesis $H_{1}$:
  1. 1.
    The Student-$t$ distribution with degree of freedom $\nu$, where the density function is given by $p(x)\propto(1+x^{2}/\nu)^{-\nu+1/2}$, $x\in\mathbb{R}$.
  2. 2.
    The generalized $\chi^{2}$ distribution with power parameter $\alpha$, where the density function is given by $p(x)\propto x^{\alpha-1}\exp(-x^{2}/2)$, $x\in\mathbb{R}$.
Both Student-$t$ and generalized $\chi^{2}$ distributions are generalizations of the Gaussian family, thus under alternative hypotheses. They degenerate to the null model when $\nu\to\infty$ ($1/\nu\to 0$) and $\alpha\to 1$. In the experiment, we start with the degenerated case mentioned above, and then vary the value of $\nu$ and $\alpha$ to evaluate the power of the goodness-of-fit tests under the alternative hypotheses. For each DGP, we generate $n=100$ samples from the distribution.
Testing Methods.
We apply our goodness-of-fit test framework for this experiment setup. Our semiparametric testing consists of two steps, estimation and testing:
  1. 1.
    Estimation: For both DGPs, we estimate the parameters of the null model $\mathcal{N}(\mu,\sigma^{2})$ using the maximum likelihood estimator (MLE).
  2. 2.
    Testing: We consider the following distance-based matrices for the goodness-of-fit tests: (1) Kolmogorov-Smirnov distance, (2) Wasserstein-$1$ distance, (3) Maximum Mean Discrepancy (MMD), and (4) Kernel Stein discrepancy (KSD).
Figure 5: Power curves of SKSD Gaussian goodness-of-fit tests with different distance-based matrices. The left panel uses the Student-$t$ distribution with degree of freedom $\nu$, while the right panel uses the generalized $\chi^{2}$ distribution with power parameter $\alpha$. The red dash line indicates the test level.
Results.
In the first case with data generated from Student-$t$ distribution, when $1/\nu$ goes larger, the observed data become more heavy tailed, thus more distinguishable from a Gaussian. Therefore, we expect the tests to have larger power against larger $1/\nu$. In Figure 5, we can see for all four matrices, the Type-I error is controlled at the nominal level $0.05$ under the null hypothesis. When $1/\nu$ increases, the power of all the tests increases to nearly 1, so all the tests are consistent against this alternative hypothesis. Among the four distance-based matrices, the SKSD test has the largest power, followed by $W_{1}$ distance, and K-S distance, while the MMD-based test shows a significant gap in power curves.
Next, we consider the case when observed data is generalized $\chi^{2}$ distribution with power parameter $\alpha$. Similarly to the previous case, we expect the power of the tests to increase as $\alpha$ goes larger. Although the Type-I error is controlled at level $0.05$ for all four matrices, the MMD test does not show a consistent power curve as $\alpha$ increases. The other three tests with KSD, $W_{1}$, and K-S distance, show a similar consistent power curve, while the $W_{1}$ distance test has the largest power, followed by K-S distance and SKSD tests.
In summary, we can see that the KSD, $W_{1}$ and K-S distance-based tests show comparable power against the both two alternative hypotheses, while the MMD-based test is not consistently powerful in our settings. Taking the computation cost and intractability of other tests in higher dimensionality into account, the SKSD test is the most competitive candidate among the four distance-based matrices.